The Ferrari Affair

2

Contents

1 Linda Logel 7 2 The Fire 29 3 Ken and Jeff 39 4 The Investigation 47 5 Paul Selinger 65 6 The Ferrari 75 7 October 11 81 8 Nate Haslett 87 9 Looking for Nate 101 10 Owen Sound 109 11 The Money Envelope 117 4 CONTENTS 12 Jackie and Steve 121 13 Sunday, October 14 129 14 Nate in Hiding 141 15 Paul Returns to Mount Dunham 149 16 Pursuit of Thieves 159 17 Home to Toronto 167 18 Search for Nate 177 19 Nate Relaxes 191 20 Paul Sacked 197 21 Nate Returns 205 22 The Night Heats Up 213 23 Nate Makes Headlines 219 24 Snow Day 231 25 Simon Greenwood 239 26 Linda’s Next Case 243 27 New Year 253 CONTENTS 5 28 Wilhelmina 265 6 CONTENTS

Chapter 1

Linda Logel

Linda Logel drove her new, used, red VW Golf west on highway 12 to Mount Dunham on her way to her first job. Linda recently graduated from Barrie Collegiate as a fire scene investigator. She was anxious and excited about leaving the safety of her parents home in Barrie, and beginning in a fresh new place, and on her own. At the same time, she was glad to be leaving some dark memories of her past behind. Up until 5 years ago she had planned to become a normal housewife raising a few children, near her parents home. One night, however, was all it took to change her life completely. This new job was a chance to start over.

At 5 feet, 5 inches tall, and 24 years old, Linda had a lot to look forward to in life. With short cut dark hair, and brown eyes, she was pleasing to the male eye. She was very fit, not overweight, with an outgoing attitude towards other people, but at the same time she always seemed to be on her guard. She noticed other people, what they were doing, and she seemed to study them. She was taking an impression of everyone she met, trying to decide whether to trust them or not.

As she drove to Mount Dunham, she noticed that houses became farms that were farther apart than houses as she distanced herself from Barrie. Trees became fewer, and the land grew flatter. In early September some fields had been plowed and some re-seeded in wheat. In others the corn was beginning to dry and whistle in the wind. The landscape made her feel isolated and alone. She began to worry about being able to find things in Mount Dunham, and how far she would have to go to buy new clothes or food. She wondered if she would be able to find somewhere to live around Mount Dunham. She was not a night or party type person, but she was hoping that it would not be too lonely.

She finally came into Mount Dunham and hit Main Street, which runs north and south. She turned left and drove slowly south down Main Street. She saw the fire station, the police station, a jewelry store, a hardware store, a grocery store, a liquor store, a library, a convenience store, a CIBC bank, a real estate agent, and an insurance dealer. As she was leaving Mount Dunham there was the Black Forest Restaurant and Motel. She had counted only 3 cross streets, including the one she had driven in on. She turned around and went north on Main Street and made mental notes of the shops again. On the north side of highway 12 were two more cross streets. Off of one she could see the public school, a mix of old and new buildings, and its athletics facilities. A PetroCan gas station was on the east side of Main Street. She pulled into the gas station and filled up her car.

“Wow. This is a small place.” she told herself, “And I thought Barrie was small. There wasn’t one clothing store here.” Linda decided to check in at the Black Forest Motel that she had seen. Then she would go to the real estate agent and see about renting something. She did not know how long she would have this job.

Linda drove back to the Black Forest Motel. As she walked into the office, there was no one there. The room was very warm due to the sun shining directly through the big window into the office. She rang the buzzer on the counter. She heard someone stirring in the back rooms behind the desk.

“In a minute!” a woman yelled.

Linda noticed a bulletin board with various meetings posted, but most of the events had already occurred. There was a notice for a house to rent with strips to tear off having the contact phone number on it. She tore one off and put it in her pocket. There was a notice about an auction for a farm with house and barns in the next month. Just then a big woman came through wearing a short light green shirt and apron. She was taking off a pair of yellow rubber gloves that were stuck on her

9

hands. An odour of Pine-Sol drifted in with her.

10 CHAPTER 1. LINDA LOGEL “Sorry, I was cleaning back there. What would you like deary?” she asked as she brushed back some loose hair hanging over her eyes, and wiped her brow with her forearm.

“I’d like a room, please.”

“One night?”

“I don’t know. It depends how quickly I can find my own place to stay.” Linda said.

“You’re moving to Mount Dunham?”

“Yes, I’m going to be a new fire scene investigator here.” Linda smiled.

“Tell you what. You can stay as long as you want, we’re not that busy here. It’s 60 dollars a night or 400 dollars for a week.”

“That would be helpful. I have a pet cat, is that a problem? He’s house broken.”

“A cat should be alright, but you’re responsible for any scratched furniture.”

“Oh yes, of course.”

Linda paid for a week using her MasterCard, and was given the key for Room 7.

“What can you tell me about this house for rent?” Linda asked while pointing to the notice on the bulletin board.

“That house is on the other side of Main and one road over. It’s a nice place. Big too. The owners live in Owen Sound. The Mitchells. They used to operate the liquor store in town, but they retired. The Mrs became ill, and it was easier for them to move to Owen Sound where their son lives, to be closer to medical facilities, than to travel back and forth.”

“Thanks, I’ll check it out then.” Linda said. With that she left the office and moved her car down to the end where Room 7 was located. She only brought her suitcase into the motel room, and the travel case with her cat. As she entered the room, there was a musty odour, covered by Pine-Sol, to the room. The furniture in the room must have dated back about 50 years. A television, the size of a three slice toaster was hanging from the ceiling in a corner. Linda thought she might need binoculars to see anything on the TV. There was a sink with a mirror above it, and to the left of that was the bathroom with a tub and toilet, and barely enough room to turn around in. Linda thought this room would hasten her to find her own place, as soon as possible. She went back to the car and got the litter box, litter, and some kibble for the cat. She let the cat loose in the room and it went about exploring every nook and cranny.

Linda pulled the little slip of paper with the contact phone number on it and sat by the phone. She dialed the number and waited for it to ring. Her watch indicated 12:30. After three rings a male voice answered.

“Hello, Mitchell’s residence.”

Linda was expecting the voice of an older gentleman or woman, but this voice was from a younger, stronger man. She gave her details to the person on the phone and said she was phoning about the house for rent in Mount Dunham.

“Oh yes, the rent is $2500 a month plus utilities and phone. You’re expected to take care of the yard and shovel your own snow in the winter.”

“Can I take a look at the place first?” Linda asked.

“Sure, see Colin Freeman at the real estate agency in Mount Dunham. He has the keys and can set up your lease.”

“Okay, thank you.” Linda hung up and looked at Ruffles, her cat.

“That one was pretty expensive Ruff. But I have to look at it to get an idea about things around here. Finding a good place here might be difficult.” Ruffles had already made himself at home on the bed. Linda flipped through the phone book, which included numbers for several small towns in that area. She found the Yellow Pages and looked for the real estate agent’s number. She called the number.

“Coldwell’s, Colin speaking.”

Linda repeated everything she had said earlier.

“I could meet you there.” Colin said, “Are you familiar with the area?”

“No, I’m not. I’m at the Black Forest Motel and can be at your office in five minutes.”

“Look, I’ll pick you up in my car, and we’ll go there from the Motel.”

“Thank you. I’ll be waiting in the front office.” She hung up and went over to the mirror and checked out her appearance. She ran a comb through her Dorothy Hamill style hairgut. She contemplated changing clothes, but didn’t think she had time to change. She put kibble and water down for the cat, grabbed the room key and her purse, then left the room and locked it.

Within a few minutes, a dark blue Nissan Altima pulled up to the office where Linda was waiting. A middle aged man in a sport coat and tie jumped out and introduced himself as Colin. He had a large head with reddish hair, and his arms seemed too short for his body. His face had faint remnants of freckles. Linda was reminded of a used car dealer, rather than a real estate salesperson. After they shook hands and introduced themselves, Colin said,

“There are two other houses for rent that I could show you, if you want to look around a bit.”

“Yes, I was thinking that $2500 a month was a little high for me.” Linda said.

“Okay then, hop in, and lets go.” Colin ushered Linda around the front of the car and opened the front door for her. Once Colin was seated and had the car running he turned and headed north, then took the first street and turned left. He went two streets further west then pulled into the driveway of a fairly large lot with a big old three story house, that looked like a mansion.

“This one?” Linda asked in amazement. She twisted her head to look up at the top of it from inside the car.

“Yep, it is probably too much for one person. Sorry, I assumed you were single, is that correct?” Colin asked.

“Just me and my cat.” Linda said. “This is definitely too big. Could we look at some of the others you have? Do you have any bungalows?” she asked.

“Yes, there is one bungalow. Only $500 a month.” he said.

“Let’s look at that one.”

Colin put the car in reverse and backed out. Then continued west out of town. Two roads further, about two kilometers, Colin turned south and pulled in to the first drive on the right. A small bungalow with a shed and one-car garage, all in goldfinch yellow siding and a black-gray roof. There were two large trees, one in back and one in front of the house. Along the south and at the back of the property were bushes that were seven feet high, planted specifically to be a windbreak in winter. The house itself sat on top of a little mound so that rain water drained away from all sides of the house.

“This is more like it.” said Linda, and she climbed out of the car and headed for the front doorway.

“The front door doesn’t open. You have to go in the side entrance.” Colin said as he pointed to the side of the house. “The owners kept the front door closed and had furniture sitting in front of the door. It is actually screwed shut.”

Linda nodded and followed him to the side door. The side entrance went directly into the kitchen area. There was a table and single chair in the kitchen. Linda went to the sink and tried to turn on the water, but Colin stopped her and said,

“The hydro is not turned on right now. You have to apply to the company in Owen Sound before they will turn it on.”

She then went to the refrigerator and opened the freezer on top. It was empty. The stove was electric and had four burners and an oven. She looked in the oven and noticed it needed cleaning. She looked up and opened a couple of cupboards. They were bare and a little dusty. It seemed the house had been sitting empty for several months.

Colin began walking past the living room and Linda followed,

“There’s a master bedroom and a smaller bedroom, and one bathroom down this way. This room back here has the furnace and water heater. The furnace works on oil that is delivered according to your usage.” Colin explained.

“Do you know how much the hydro and oil cost, say per month?” she asked.

“Both of those depend on the time of the year. Per year you are looking at about another $2000 in total.”

Linda looked through the rest of the house. The living room had a large picture window that reached from the floor to the ceiling. Linda thought how Ruffles would love that. The front door was indeed screwed shut tightly.

Linda did not fall in love with the house, but it was functional for her purposes, and the price was right. She was surprised that such a place was available. She turned to Colin and said,

“This should work out just fine. I’ll take it. What do I need to do?”

“We can go back to my office, and you can fill in the forms there. Do you have a bank account here yet?”

“No, I just arrived today.”

“And you said you’ll be working in Mount Dunham? Where was that again?”

“I’ll be a fire scene investigator.”

“We can’t sign the lease until you have your bank account, and until I can confirm your employment with the fire station.” Colin said.

“No problem, I’ll try to get that done as soon as I can. Tomorrow hopefully.”

Linda walked through the house once more. There were no beds and no other furniture. She would have to ‘rough it’ for the first month or two until she could find some furniture. She had brought camping gear with her, an air mattress, and tin cookware. She could manage without furniture for a little while. Her previous apartments had always been furnished. She would have to scrounge around for reject furniture.

Linda went outside and looked around in the separate garage. It was empty, and large enough for her car. She would have to buy a lawnmower, or hire someone to cut the grass. Soon it would be winter and that would bring snow. Again, she might have to hire someone to clear her driveway. She walked around the perimeter of the property. The house sat on a one acre parcel of land that obviously belonged to the farmstead that surrounded the house to the west and south. The roads were gravel. Roads were set up north-south and east-west and about 1 kilometer apart. That meant she could run 4 kilometers in the morning by just running around a square. Yes, she thought to herself, this place should work out very well.

* * * * *

The next day Linda walked to the fire station from the motel, on the east side of Main Street. After greeting the fire chief and some of the fire fighters, she was informed she had come to the wrong place. The fire scene investigator’s office was not with the fire station, but was associated with the police station, which was across the street on the west side of Main Street. Linda apologized and felt embarrassed. She thanked the men and crossed the street.

Inside the police station, an officer at the front desk greeted her in an official manner.

“How may I help you?” asked a tall, well-built, uniformed officer.

“My name is Linda Logel, and I’m supposed to begin work here today.” Linda said.

“Do you have your Offer of Employment letter and your driver’s license, please?” asked the officer in a deadly serious tone. A tone that strongly indicated that humour would not be appreciated or tolerated. He studied the offer letter and her driver’s licence. He typed in Linda’s information into the computer, and the fact she was there to begin a new job. After hitting enter, the computer searched several databases for the name Linda Logel. Linda was becoming a little nervous. No warnings came up on the computer screen, meaning that there were no outstanding warrants against Linda in the system. Then he spoke into the microphone attached to an interoffice loudspeaker system and said,

“Captain Muller, you have a visitor at the front desk.” You could hear the announcement echo throughout the police station. Then he smiled at Linda and said,

“The captain will be with you shortly, and welcome to Mount Dunham.” He handed back her letter and driver’s licence.

“Oh, why thank you, er.” Linda felt a little relieved by his smile.

“Justin Jankowisz” the officer said.

“Nice to meet you.” she smiled, and looked around a little. Captain Muller came out from one of the offices at the back and walked to where Linda was standing. He was not as tall as Officer Justin. He was thin with greying brown hair, and wore glasses. He was not in uniform, but wore a suit with tie. He had the look of someone who was always busy, but also like someone in charge.

“Hello, I’m Don Muller. You must be Linda Logel. I’ll be your supervisor. You’ve met Justin here. There are three other police officers at this station that you’ll meet over time. There is always one officer in the station at all times, and one in a patrol car. Over there is Tracy Morgan who essentially keeps everything flowing around here. She files reports and gathers information about anything from anywhere at the speed of light, almost. Without her this station would not be able to function properly. If you need to know anything, ask Tracy first. She either knows already or she can find out very quickly. She has worked here for over ten years, and she knows everyone in this town. Whatever you do, don’t upset her.”

“Not a problem. Are there other fire investigators?” Linda asked.

“Yes, myself and Gord Walker, but he has elected to be part-time, and offered to help when we need it. Otherwise, he prefers to raise cattle west of town. You are taking his place essentially. You will be the main fire scene investigator here, after me. I’ll try to break you in slowly so you have a chance to learn procedures around here. Eventually I hope that you will take over most of the investigating duties. The fire trucks are called out about once a week. Most of those are small fires that don’t require any investigation by our department. We only investigate fires where an entire house is burned, or if people have been injured or killed. Sometimes that means investigating car accidents in which cars have gone up in flames. When there is little for you to do, you may be asked to help with police investigations, but these are usually handled by the Ontario Provincial Police. You never know when life will become complicated around this little town.”

“We covered all types of fires in college. I’m ready to go, sir.” Linda said.

“Excellent. You’re probably wondering why the police and fire investigators are in the same office and why I oversee both areas?” he asked.

“I assume it is to save money for the town and to be efficient at the same time. After all Mount Dunham is not very large.” Linda said

“Not large, but the area we cover is as large as some of the bigger cities, like London, Woodstock, or Kitchener. Just not as many people.” he said. “Where is it you come from again?”

“Barrie, sir.” Linda said.

“Just call me Don. We are not so formal here.” Captain Muller said. “I’ll show you to your office, if you’ll follow me this way, please.” Don showed Linda through the building. Besides offices there was an interrogation room with a supply and storage room across the hall. Next was a lunch room with vending machines. At the end of the hall was an exit to the rear of the building, which was the way that all of them came in and out of the building. There was a parking area at the rear and a garage for the police cars and SUVs. Back inside and marching past the lunch room, supply room, and interrogation room again, they stopped at an office which was 14 feet by 11 feet with small windows on one wall above the 6 foot level. Lots of indirect sunlight came through. The ceilings were ten feet high, and all the walls were white. There was a desk and two tables in her office. A computer was sitting on the desk, with a phone, and the two tables were empty. A set of three wall mounted shelves were bare and a little dusty.

“This used to be Gord’s office, but he moved out about 3 months ago. Gave us notice that he only wanted to be part-time. Then we had to get permission to hire a new person, you. Bureaucracy exists even in Mount Dunham, and it moves as slowly as anywhere else in this province. This will be your office now.” said Don.

Linda went to the computer and noticed it was not plugged in or connected to anything. Captain Muller noticed too, and said,

“Tracy will get you started on that. She’ll have to get you registered for the police and fire systems, so you can review other crimes and so forth.”

20 CHAPTER 1. LINDA LOGEL “Is there anything in particular you would like me to start working on today?”

“Its pretty quiet right now, so just get yourself settled in and talk with Tracy. She’ll have forms for you to sign for payroll and insurance. And she’ll get you a badge and ID so you’ll be official. For now, your work hours will be from 8 am to 5 pm, but if we happen to get a big fire, then you may be called out at any time. You’ll have to get uniforms to wear to work. Tracy can tell you where to get those too. You should keep a spare uniform here in the office. Sometimes they get dirty or wet. Well, I’ll leave you with Tracy, and welcome aboard Linda.” He shook her hand and noticed the softness of it, not like a man’s. He passed Linda off to Tracy, making the introductions, and then excused himself and went back to his office.

Tracy was around forty years old, and oozed with the confidence that she was very good and efficient at her job. She looked at Linda with some skepticism, but spoke kindly to her. Tracy never had another female employee at the police station. Tracy had already reviewed Linda’s application for the job and had performed a background check on her, so that she knew everything about Linda, except her personality, before Linda was offered the job. Within an hour Linda was officially signed in and registered as a new employee. In the next hour, Tracy had Linda’s computer doing searches of the police databases. At the lunch break, Tracy took Linda to a clothing store in Owen Sound to get her fitted for uniforms. The uniforms would be available the following week.

During the shopping trip, Linda learned that Tracy was married with two daughters, ages 6 and 8. Her husband, Pete, worked for the telephone and cable companies of Orangeville, servicing phones and satellite dishes in several counties. Tracy was fairly laidback and humourous. She had funny stories about her girls and the things they had gotten into. She was glad that her two girls seemed to be close to each other and spent a lot of time together. Linda knew, however, that at some time sisters split up over something and never talk to each other again, as she had done with her younger sister, Jenny. About the time Linda discovered boys, Jenny almost ceased to exist to Linda. There was no fight between them, just a parting of paths. Jenny was too young and not interested in keeping up with Linda’s quest for boy attention. Tracy’s kids could be entirely different than she and Jenny had been.

While Tracy was talking about her girls, Linda’s mind drifted off thinking about her early obsession with boys. She had met Tim Harrington, who was 3 years her senior, in high school. Tim was handsome and friendly and made Linda feel special, so that after only a week Linda agreed to a date with Tim. After dinner and a movie, Tim was driving her home, but she didn’t recognize where they were going. Tim drove down a deserted road some distance from town. She had become uncomfortable with him during the date, and now she was really getting nervous. When Tim stopped the car and turned it off. It was dark and quiet. Linda said,

“Take me home please. Now.”

Tim shook his head negatively and pulled out a gun from under the seat.

“Take off your clothes.” he said.

“No, please, just take me home.” Linda begged.

“Take’m off.” he said more forcefully, and he pointed the gun at her head.

Linda started crying and took off her clothes. When she was naked and shaking, he said,

“Now get in the back seat, and quit crying. You’ll enjoy it, believe me, they always do.”

Linda climbed over the front seat into the back. Tim got out of the car and started to take off his clothes. Linda took the opportunity to jump out the other back door and run off the road. There was a woods and she ran for her life in the dark. She heard Tim swearing, then she heard two gun shots and bullets whizzing past her. She kept running and put lots of trees between her and the car. Her bare feet were hurting from the sticks and stones that she had been tramping on in her escape. Her body was scratched as well from branches she ran through. She stopped and leaned against a tree in the dark rubbing her feet which were actually bleeding. Suddenly she heard someone else tramping into the woods.

“Where are you, you stupid bitch?” Tim yelled.

She could hear him stumbling in the dark. He was going in the wrong direction to the right of her and she kept quiet. She began to move away from him whenever he would move. However, she broke a stick and it attracted his attention. Another gunshot and a bullet whizzed through the trees near her. Then she heard him moving in her direction. She grabbed the end of a thick branch and bent it back as much as she could, then leaned tightly against a tree trunk. Tim came closer, but she could not see him. He was starting to walk away from her, so she shifted her feet to make a noise. Tim stopped and listened,

“You’re close now, aren’t you bitch.” he said, and he crept a step towards her, in her direction. She caught a glimpse of him and waited,

“Come on, one more step forward please.” she said to herself. Tim moved one more step.

“There you ... ” is all Tim said, then Linda let loose the branch and “thwack” the branch hit Tim full force in the face and knocked him to the ground unconscious. Linda grabbed the gun from his hand and started back towards the car. She hid the gun under a fallen tree trunk on her way out. Once she was at the car she dressed herself, and got her things. To her surprise, the car keys were still in the ignition, so Linda took the car. She drove straight to a police station and reported the assault. Tim was arrested and sentenced to five years in jail. His nose had been broken by the tree branch in his face, and this tended to detract from his otherwise good looks. On the day he was sentenced he stared hard at Linda across the court room.

Linda would never trust the male gender again. She took martial arts training in order to protect herself in the future, and she became very proficient. All of these memories raced through Linda’s mind in a second or two. She became aware that Tracy had stopped talking, and she came back to the now.

After Tracy and Linda returned from lunch, Linda stayed in her office and retrieved files covering the last two fires handled by Captain Muller and Gord Walker. She particularly noticed the types of documentation that were required and the amount of detail in the reports. One fire resulted in a casualty and the other resulted in significant damage to two houses within Mount Dunham, the result of an accident due to negligence 24 CHAPTER 1. LINDA LOGEL

in an outdoor burning of old wood. She found the actual case files and browsed through them as well.

About 3 pm, Justin called Linda to the front desk. He explained about logging in phone calls and walk-ins into the computer system for every event. Linda was nervous about taking over on her first day. Almost as soon as Justin was out the door, a call came in and Linda answered.

“Police station, how may I help you?” she said. “There’s a truck parked across the end of my driveway.” an elderly

woman said.

“Have you talked to the driver about moving it?” Linda asked. “Driver’s not there.” the woman said abruptly. “Are you in need of moving your car out? To go somewhere?” “I don’t have a car, and I don’t need to go anywhere.” “Then what is the problem?” “The problem is that the truck is illegally parked, and if I get any

visitors they will not be able to park in my driveway.” the woman said.

“Okay. Can you give me your name and address please?” Linda grabbed a pen and piece of paper.

“This is Wilhelmina Schrandt.” she said. “And your address is?” Linda waited. “You people know where I live. Put my name into your computer,

it knows who I am.”

Linda was bewildered and looked at Tracy and whispered, “Wilhelmina Schrandt?” Tracy nodded, so Linda said,

“We’ll send someone over shortly. Be patient please.” “Do you want to know the colour and make of the truck, and the

licence plate number?” Wilhelmina asked.

“That won’t be necessary, the police officer will get that information when he stops by your house.” Linda replied.

“Thank you young lady.” she said, and hung up.

Linda looked at Tracy questioningly. Tracy explained,

“That woman phones here at least once a week. Most of the time it is for small things, like kids cutting across her lawn. She sits at home all day by herself. She must be in her 80’s. I think she calls us just to talk to someone. What did she want this time?”

“There’s a truck parked across the end of her driveway blocking entrance to any potential visitors.”

“That’s a new one. Her address is in the computer. She just lives around the block from here on Coyote street.” Linda began keying in the information about the call.

“You try to get their name and other information from their driver’s licence before they start in with their complaints or questions.” Tracy said.

Linda called an officer on patrol and relayed the information. She had to explain who she was, too. Linda keyed in the name of the officer who was going to respond to the call, and the time.

About ten minutes later, a tall man wearing a baseball cap, button up shirt hanging out over his well-worn blue jeans, with a moustache and goatee came into the station. He was solidly built with a slight paunch. He was not someone Linda would take home to meet her parents.

“How may I help you sir?” Linda asked, and she noticed a faint alcoholic odour and cigarette smoke emanating from the man as he leaned on the desk. She guessed he must have been around thirty years old. His eyebrows were bushy and made his brown eyes look darker. Linda was repulsed by the man, part sasquatch she thought to herself.

“There’s an empty farmhouse east on highway 12, about 15 to 20 kilometers out. I was wondering if it was for sale.”

Tracy heard the question and answered,

“That one’s abandoned. The current owner is the bank. It’s scheduled for auction in a month or two. You can go to the bank for more details about that.”

“Thank you, ma’am.” and the man turned and left before Linda could ask to look at his driver’s licence.

“How am I supposed to log that one?” she asked. “Did you know that person?”

“New to me. Just enter the time and what he asked about in the log book, and maybe a description of him.” Linda tried to put in something, but the computer seemed to want certain fields to be entered. She made up a name, John Doe, for the name field, guessed he was about 32 years old, maybe 220 pounds. Tracy came over and showed Linda the surveillance camera and how to retrieve his picture and add it to the report.

“Wow,” said Linda, “that was cool. Do we need to do that each time?”

“Doesn’t hurt.” Tracy replied.

Finally Justin came back from his break. She told him about the phone call and the male visitor. She explained about the man and how he left before she could ask for his ID.

“You always ask for their name and address first, using their driver’s licence if possible, before you listen to their stories.” Justin scolded her gently.

“Sorry,” said Linda.

“You never know about people. Once there was a guy who came in and asked about a woman that lives around here. The computer came up with a court order against the guy to keep away from his wife. We gave him a warning and asked him to leave town. Next thing you know, the woman was taken to hospital beaten up rather badly.” he said. “He was arrested and is now in jail. We usually call the person they are looking for, to see if that person wants to see the visitor.”

“Thanks, I’ll remember that.” Linda said. At the end of the day Linda was not very happy with her performance on her first day. Although she was not trained as a police officer, she thought she should have known more about office protocol. She wondered if she would last in this job. On the bright side, she had not been fired.

Linda stopped at the bank on the way back to the motel. She set up a savings account and a chequing account. She deposited most of the money she had brought with her to Mount Dunham. Next she stopped at the Coldwell Realty office and gave her information to Colin Freeman in order to complete the lease agreement. Colin gave her two sets of keys to the house, and explained about the monthly payments. He also gave her the name and telephone number of the owners of the house. If she had any problems with the house, repairs and the like, she was to phone the owners directly. From there she went back to the motel and collected her clothes and her cat and drove out to the house, even though there was no hydro. Linda drove back into town and ate at the Black Forest Restaurant for the second night in a row.

The new house echoed without furniture in it. The floors were parquet wood flooring, so every footstep she made was heard. Every sound, inside or out, seemed magnified in the darkness that night. She thought about getting surveillance cameras for the outside of the house, or at least lights that would come on whenever movement was sensed. She shivered as she thought of the night of her date with Tim Harrington. She vowed that that would never happen to her again. Ruffles snuggled next to her legs on the air mattress, and the cat purring made her feel better.

The next morning she awoke at 5:30 am, got dressed into her sweat clothes and laced on her running shoes. Out the door Linda went for a run. She headed south from the house, turned right at the next road heading west, then right again heading north, and right again heading east, until she was back to her road. On the run she counted only two other houses, farms really, within that square. By the time she got back in the house, the sun’s rays were coming through the living room window. “Day 2” she thought. The cat meowed.

A couple strokes of a brush to her hair and she was ready to go. She would have to go to work without a shower today. Linda put on some 28 CHAPTER 1. LINDA LOGEL clean clothes and left the house for her second day of work.

The Fire

On a cool October night, out in the country east of Mount Dunham, one could see billions of stars, and an owl hooted somewhere in the distance warning creatures that it was on the hunt. At the abandoned farmhouse of Ruth and Morris Blanchard, a tall man in jeans, boots, t-shirt and denim jacket leaned against the fender of his pick-up truck drinking a can of beer and smoking a cigarette. His eyes sparkled with excitement. Beer foam clung to his moustache and goatee, which he wiped off on his sleeve. After emptying the beer can he squashed it against his leg and tossed it on the ground. He walked towards the porch of the abandoned farm house. He flicked his cigarette onto the porch, and swoosh flames leaped up as the cigarette bounced and landed in a pool of gasoline.

The flames leaped as if to grab at him, and he jumped back quickly, knocking his baseball cap to the ground as he covered his face.

“Whoa!” he laughed. The flames always seemed to leap at him. He watched with a smile on his face. The fire spread quickly and ran into the two storey house. Within seconds the entire first floor was engulfed in flames. He picked up his cap then backed towards his pickup. He climbed into his truck and admired the fire. He liked way the flames licked upwards randomly, and the way that burning items reacted to the heat were fascinating and hypnotizing to watch.

Flames had now reached the second storey, as well as to the attached garage. He picked up his cell phone from the front seat of the truck and was about to make a call, when he noticed a farmer and a dog running across a field from the nearby farmhouse, about a kilometer to the west. The dog would be there soon.

“Shit, stupid dog.” he said to himself. “Sorry, I can’t stay.” He could hear the farmer calling for the dog to stop. He started the truck, then backed up quickly, turned and headed for the road. He turned right and headed west squealing his tires and throwing gravel as he left.

The farmer stumbled on the dirt clods in the plowed field getting dirt into his boots. He stopped and watched as the dark coloured pickup truck drove out the laneway and onto the road. The dog’s outline was visible against the bright fire. The farmer was breathing heavily by the time he reached the front of the house. The dog came to the farmer and he leashed the panting dog. Together they watched the house burn. The grass, weeds, and two trees around the house were on fire too.

Suddenly, a small explosion in the garage caused flames to leap into the sky for a few seconds. Parts of the garage roof soared into the sky and headed towards the farmer. The farmer and the dog flinched at the explosion, then turned and ran away from the house. The roof debris landed about six inches short of hitting the farmer’s foot. The two kept running, not looking back.

When they stopped, there was a loud crack of wood splitting and they watched as the roof over the main house caved in on top of the second floor. The heat increased and the farmer and dog had to move further away from the house. The farmer noticed someone walking up the lane.

“Mrs Palmer, did you bring any marshmallows?” he asked. Mrs Palmer lived in the house to the east by herself.

“No thank you, I’m diabetic, you know.” she said. “Hello Mickey” she addressed the dog and patted his head, as he wagged his tail in greeting.

“I was just kidding, Mrs Palmer.” he said.

“I know that Fred. Isn’t this just dreadful? Did you see the man in the truck?” she asked.

“No, I just saw the truck leaving. Did you see anything?” Fred said.

“Not really. I saw someone poking around over here, but didn’t pay much attention to it. Until I saw the flames. Then I phoned 911.” she said.

“There’s not going to be much to save by the time the fire department gets here. Luckily the barns are not on fire.” he said.

“Why do people have to destroy things that don’t belong to them?” Mrs Palmer asked.

“It’s a general lack of respect in the younger generation, I think.” said Fred. “Their parents haven’t given them the basic discipline they need. My father would have stripped my hide if I even looked like I was going to be disrespectful. These days kids get away with everything.”

“That wasn’t a kid I saw.” said Mrs Palmer.

“I was just talking generally. I don’t know if that was a kid or not.”

“Good thing this house was abandoned. At least no one has been hurt. Say, do you know what happened to the Blanchards anyway?” Mrs Palmer asked.

“I don’t know where they went or why, but I expect they ran out of money and just left. One day they were here and the next day gone. Not a word from them since. The bank sold all of their equipment and furniture.”

“I heard there was going to be an auction to sell the property sometime soon.” she said. “How much do you expect this place is worth?”

“The house wasn’t that great anyway, but the barns are useful and the land would be the main value. Maybe six hundred thousand dollars? Maybe more?” Fred said.

Flashing lights and sirens of fire trucks from Mount Dunham could be seen and heard rolling down the road. The land in this part of Ontario was flat and almost treeless. One could see for many kilometers in every direction and sound traveled unimpeded. Thus, even though the fire trucks could be seen and heard, they were still some distance away.

Fred held on tightly to Mickey’s leash as two trucks came up the laneway and stopped. Firefighters jumped out of the trucks and attached hoses to the trucks with great efficiency. The fire chief came up the laneway next in an SUV with flashing lights. The lights of all vehicles were aimed at the house, as well as the spotlights on the trucks. Soon water was being sprayed onto the fire.

The chief ordered his men saying, “Keep the fire from spreading to the barns.” Turning to the two people watching the fire, the fire chief said, “Thank you Mrs Palmer for phoning us. Fred, what are you doing here?”

“Oh, Mickey took off on me when I let him out tonight. I had to chase him over here.” Fred said. “We saw a pickup truck here, but didn’t get any make or licence plate numbers, not even a colour, except that it was dark coloured. Mrs Palmer thinks there was just one man in the truck. You might have passed the truck on the way here. It was headed for town.”

“No, we didn’t pass any vehicles, so he must have turned off and hid until we passed. So the fire was deliberately set. This house has been abandoned, hasn’t it?” asked the fire chief.

“That’s right, about three years now. Mrs Palmer and I were just talking about that.” said Fred.

“What was their name?” the chief asked.

“Blanchards. Ruth and Morris.” said Fred.

“We’ll just have to control the fire for now, and send an investigator in the morning. Thanks for your help. You may go home now, we’ll take care of things here.” said the chief.

As they walked down the laneway, Mrs Palmer asked, “Fred, could you have your boy, Ken, come over and cut my grass one more time before it snows please?”

“Sure thing Mrs Palmer.” said Fred. Fred’s boys were eager to earn extra money and Mrs Palmer besides paying with cash, usually had chocolate chip cookies by the dozen, even if she was diabetic. Fred decided to take the road to his home rather than tramping through the field again. Mrs Palmer also walked on the road. At the end of the laneway the two turned and headed in opposite directions. The farther they got away from the fire the colder it got. October nights could go below 0 C. Fred realized he should have had a coat. Mickey, a black lab, pulled on the leash a little. Fred looked back at the fire. He was surprised that the entire house had burned to the ground so quickly. He could smell smoke on his clothes.

The fire fighters stayed with the fire until it appeared to be entirely finished. Yellow caution tape was put up in a wide circle around the house. Two men were left with the SUV to keep watch until the next morning when the investigator would arrive. This was to ensure that the fire did not flare up again or spread to the barns.

* * * * *

Nate Haslett, sitting in his pick-up truck in the dark on a sideroad could see that his fire was out, and watched the firetrucks return to Mount Dunham. That stupid dog had prevented Nate from finishing the job and setting fire to the barns. He started his truck and drove towards the fire site. He cursed to himself when he saw the SUV parked there and drove past. He turned right at the next sideroad and decided to go home. He lit another cigarette and popped open another can of beer. Perhaps the abandoned farmhouse was too close to the neighboring houses. He would have to be more careful next time in selecting a place.

When he got home almost an hour later, the lights in the house were all out. His mother and sister were already in bed. He could hear his dog barking inside. Brutus could tell from the truck’ engine noise that his owner, Nate, was home. Nate opened the door to the house and Brutus stuck his nose through to greet Nate. Brutus’ tail was wagging and the dog had stopped barking.

“Shhh. Quiet.” Nate whispered, “You’ll wake everyone up.” Although that command was much too late to make any difference. Brutus was a dark, almost black, German shepherd, ex-guard dog for a security company. He was retired from service after he was severely injured from a break-in on the premises to which he was assigned. The intruder shot Brutus in the rear which crippled his right rear leg. Nate took the dog and nursed it back to health. Brutus still had a little stiffness in his rear leg, but could get around. Brutus became attached to Nate, his mother and sister, and became a good watchdog that would bark and snarl at any strangers on the property. Brutus had scared away several unwanted solicitations just by the sound of his bark. Once a Jehovah’s witness came to the door, and ignored the warning sounds of the dog. When Nate’s mother answered the door, Brutus pushed past her and jumped on the unlucky solicitor and bit him in the face and on the arms before Nate’s sister could pull the dog away. After that experience, a “Beware of dog” sign was displayed permanently by the doorbell. Nate had to pay for the medical expenses for that victim.

Nate went up the stairs to his room, and Brutus followed him. Nate had a small desk by the window with a computer on it. He used his computer for two important tasks. Because Nate did farm work on a casual basis, he had to keep accurate employment records for his income taxes. In the last year Nate had worked for five different employers helping them to plant, and weed, and harvest their crops. He worked where he was needed, and this was all over southwestern Ontario. Most of his income was spent on beer and on his truck, so that he saved little money. He gave a small amount to his mother, from time to time. Thus, his pay cheques were not consistent in amount or timing, and during the winter months they almost stopped completely.

The second task for which Nate needed a computer was to search the internet for abandoned properties. His obsession with fire seemed to be getting stronger over time. His need to relieve the pressure of this obsession had also become stronger. He chose abandoned properties because many of them were old and starting to fall apart, or they were not useful in their present condition. Nate figured that it was therefore better to burn these poor buildings and build new ones. In essence he was helping society and saving time and money by burning them down. Also, with abandoned buildings there was no chance of accidentally killing someone.

Nate went to bed, and as sleep came, the nightmare of his first fire would return. When he was just twelve years old, Nate had ridden with his father to the grainery mill with a load of wheat. Nate was outside at the back of the grainery mill playing with his magnifying glass. He learned that the magnifying glass could be used to burn paper and rags and insects. The rags he set on fire that day were saturated in oil, and they suddenly got out of control. Nate stood and watched, frozen, as the rags burned and soon the mill wall had caught fire. Nate ran and hid across the road from the mill. The mill with all of its dust and wooden beams went up in flames quickly.

When it was over, Nate’s father and another worker at the mill had died in the fire. The fire was ruled an accident, blamed on the oily rags, and Nate never told anyone that he had started the fire, or that he had been the cause of his father’s death. Nate’s punishment was that he had to re-live it over and over for the rest of his life. However, he was still entranced by fire and flames. After the mill fire, Nate became more of a loner than he already was. He did not have any friends, male or female, and was not very gregarious.

During high school Nate would often set fires inside garbage barrels around the school. These fires usually just caused a big stink and never harmed anyone. Around his home he had a place where he enjoyed burning branches and leaves and paper and wood garbage from the house. Once a fire was going, Nate would keep adding things to it for several hours, then watch it slowly burn itself out. He frequently got embers landing on him and making little holes in his clothes, which led to little burns on his arms and hands.

On one windy day, sheets of burning newspaper got picked up by the wind and made their way to the roof of the garage. Nate had to quickly hook up a hose and spray the roof. He also sprayed the fire to put it out, to prevent the wind from spreading the fire more.

When Nate had learned to drive, he went on camp outs in nearby woods, and would have fires during the night. By this time, Nate understood fire better and he knew how to control it, or at least that was his belief. He was still getting little burns, so he did not really control fire as much as he thought he did. One might have said that Nate did not show proper respect to the power of fire, and that perhaps one day Nate might be taught a fatal lesson. For now he was young and invincible.

After high school, Nate started working for other farmers who needed his help. Working on a tractor in a field by yourself, was ideal for Nate. He would eat his lunch with the other workers, and on occasion would accompany them to a bar to have a few beers after hot, dusty days in the field. Otherwise Nate kept to himself, and did not really get to know any of the other workers, other than their names. He started smoking at the bars too, although he was genetically inclined to smoking from his mother.

During this time Nate got to see many parts of rural Ontario, and he noticed all of the abandoned farms and wondered about them. There were barns overgrown with vines, and holes everywhere. Other barns had all of the side boards removed, so that only the frame and roof were left standing. Houses were boarded up, or not, and windows all broken, and the houses needing a paint job. One could almost see the families and children that had lived there in the past. Nate wondered what happened to them, what events had forced them to leave.

Over time, Nate became more of a drinker, heavy drinker. When drunk, he was more likely to become violent in nature and disagreeable with anyone else. In a sense his behaviour, while drunk, was unpredictable. One night while drinking he convinced himself that he had been cheated out of a day’s worth of pay. He became loud and abusive to other patrons, and so he was asked to leave the bar. A policeman had been called in to remove him, but Nate had obligingly left on his own before the police arrived. On his way home, he stopped at an abandoned farm and started a fire, his first of many to come. Because he was usually drunk when he set the fires, he seldom remembered setting them, much less where they were.

When Brutus came along, Nate slowed down on his drinking and spent a lot of time with Brutus, nursing his crippled leg. Brutus was a new focus for his life. Nate even set fewer fires after he had Brutus. Now Brutus was turning 6 years of age, for a breed that averages about 7 years of life. His mother and sister both found Brutus to increase their feeling of security at home. Brutus had only injured that one Jehovah witness, such that none ever ventured on their property again, as well as those who had heard about the incident. When Nate was away from home, working, he tended to drink more than usual, and Nate’s ties with his dog became weaker. But Brutus always remembered Nate’s tender care and was forever loyal to Nate.

* * * * *

Being October, Nate had few opportunities for work on farms, and so Nate started looking for anything that would pay. He signed up for running snow ploughs for the winter, but that depended on the amount of snow that would be coming. He thought about learning how to drive the big rigs, but his mother told him they wouldn’t hire a drunk. Nate thought about becoming a mechanic, but he really had trouble remembering details, and could not tell the difference between a what-cha-ma-callit and a do-hingy. That morning he got a phone call from a farmer near Leamington, and he left in his truck right after he had breakfast. There would be enough work for a couple weeks.

38 CHAPTER 2. THE FIRE

Ken and Jeff

“You two should have seen the fire last night.” Fred said as his two boys came downstairs and sat at the kitchen table. They looked at him somewhat drowsily, but the word fire shook them awake.

“Where?” Ken, the older of the two boys, asked.

“The Blanchard’s place.”

Both boys jumped from the table to run outside, but before they took a step their father yelled,

“Whoa, whoa, whoa there. Sit down and finish your breakfasts.” He paused, then added, “They don’t need you two snooping through things. You could get burned or cut. Besides it is none of our business. If you leave any evidence over there, they might think you two started the fire. You can wait a couple of days until the official investigation is over.”

Discouraged, the boys sat down and started their breakfasts, but then they began asking questions about the fire, number of fire trucks, and so on. Finally, their mother said, “Enough about the fire. The school bus will be here in ten minutes.”

“Hey Ken, Mrs Palmer asked me to tell you she would like her grass cut one more time before it snows. So you should probably do that this coming weekend, please.” said Fred.

“Sure thing dad.” said Ken.

Ken ran upstairs as soon as he was excused from the table. He brushed his teeth and combed his hair one more time, then grabbed his backpack and jacket and headed out the door. Roger followed him a couple minutes later. They stretched their necks towards the scene of the fire. Smoke still rose slowly from where the house used to be. There was an SUV with two people sitting in it, in front of the house.

“That must have been cool to see last night.” said Roger, “Are you going to tell Jeff about it?”

“Yeah, when I get to school.” said Ken.

The school bus pulled up to their lane and put out its stop sign. Roger climbed on first and headed towards the rear. Ken usually did not sit with Roger because he was embarrassed to be seen with his younger brother or any of the other smaller kids. Today, however, Ken had to sit with Roger because the other options were less appealing. There was a girl, that no one at school ever talked with, and there were some older boys who made it clear by their postures for Ken to keep moving. And then there were younger kids. Roger was the least objectionable option.

Ken bounced along with the movement of the bus down the road. His eyes were straight ahead, not really seeing anything in particular. He was thinking about the fire and his friend Jeff Blanchard. When Jeff lived at the farm house, they used to play together during the summers, and also helped with the farm work. They enjoyed playing computer games and continued to compete over the internet. At the age of 15, Ken did not understand the legal aspects of the Blanchards’ sudden departure from their farm, just that the authorities were after Jeff’s parents.

At school in Mount Dunham, Ken went directly to the school library, and logged on to one of the computers. He wrote a quick message to Jeff, it said

        Jeff, your old house burned to the
        ground last night.  Someone set it on fire.
        There was nothing left this morning.
        Thought you should know.

        Ken

Ken logged off and went to his homeroom. During the day, Ken and Roger told all their friends about the fire. Everyone had their opinion about who started it. Suspects ranged from older brothers to deranged uncles, to rumours of escaped convicts, and one suggestion was that it was an alien. In truth, no one really knew, but it was fun to speculate. By lunch time, the fire was old news and quickly forgotten. In Ken’s English class that afternoon, they were assigned to write a paragraph about an event, as if they were writing a story for a newspaper. Naturally, Ken wrote about the fire and made up a gang of female cheerleaders as those who started the fire, as well as a family of victims who died in the fire including a goldfish that boiled to death.

* * * * *

When Ken arrived home after school, he went to his room and logged on to his computer. He was eager to see if Jeff had answered his message. His eyes widened when he saw Jeff’s name in the senders list. The message said,

        Call me.
        705-871-2386

Ken was lucky to have his own cell phone which was for calling his parents if he stayed after school for sports or whatever. He called the number, but no one answered, so Ken decided to try again after supper. Ken knew Jeff was active in football and was probably at practice.

After supper he went to his room and closed his door. This time his call was answered.

“Hello?”

“Hi, Jeff? This is Ken.”

“Did our house really burn down or are you bullshitting me?”

“No shit Jeff, your house is toast.” said Ken.

“What about the barns?” asked Jeff.

“The barns are OK, just the house is gone.”

“Listen, I need you to do something for me. I hid an envelope with some money in it, in our old hiding spot. About 50 dollars. You need to get it for me.” said Jeff.

“Why didn’t you take it with you when you left?” Ken said.

“We left pretty quickly one night. I was barely awake. I didn’t know we were leaving permanently. I thought I’d be able to get it later when we returned.” said Jeff. Ken knew that it was also true that his parents would have taken his money, if they had known about it.

“My dad told us to stay away from the place until the investigators were finished with it.” said Ken.

“Its my money, Ken, I earned it, and I want it. Maybe you could go over there tonight? Remember our hiding place? The envelope is attached to the underside of the trap door.”

“Man, going over at night would be worse. Don’t worry I’ll get over there sometime and find it for you. I’ll let you know when I have it.” said Ken.

“Thanks Ken. Don’t tell anyone, please?” said Jeff.

“Don’t worry, I won’t tell anyone. Bye for now.” said Ken and he disconnected.

Ken gazed at the cellphone and thought about Jeff’s request. Ken had to obey his father, but he also wanted to help his friend. He would wait until Saturday. He could go over while his parents were in town shopping for groceries. Technically, he didn’t have to go near the house, so he wouldn’t be in any danger. His parents would never know, if he was quick about it.

Ken’s thoughts drifted to the hiding spot in the barn. The trap door was heavy and lifted up from the floor and attached to the wall behind it to hold it open. The hiding spot was basically a box with some straw in it below the floor level. Jeff’s father had not kept anything in it. Jeff and Ken could fit into the box easily at the same time, at least two years ago they could. When closed the box became stuffy and hot. Roger would use Mickey to help him find the two boys. Mickey could track them anywhere. Once they had tried to take Mickey into the box with them, but Mickey barked until he was let out.

In truth, they had only hid in the box about four times. If he or Jeff found any neat rocks or feathers or anything of unique significance, they kept an old coffee can with a plastic lid in the hiding place. Ken did not know if those treasures were still in the hiding place. He was looking forward to visiting the old hiding place.

Ken changed into his work clothes and went outside to help his father. Ken could drive the tractors and even back up a tractor with a four wheel wagon latched onto the back. Ken was often used for transporting wagons from the fields to the farm, and taking the emptied wagons back out to the fields. He was not allowed to plant seed or disc or combine the crops. These required his Dad who could look after breakages of equipment and come up with quick fixes. Some of the machinery required more strength than Ken had. Ken loved living on the farm and being in the country.

Ken’s father had nothing specific for Ken to do that evening, so Ken went into the barn to shoot baskets on their indoor half-court. The court was an obstacle course with beams that supported the walls in an upright position, one on either side of the basket. At the top of the key was a square hole in the floor which led to the lower level of the barn. So one could either fall through the hole, or trip on a support beam. The basket itself was flush against the wall so that lay-ups were abruptly finished by crashing into the wall.

If the physical hardships were not enough, after ten minutes of bouncing the ball, there would be a dust cloud in the barn that caused any light coming from the windows to be filtered. The dust penetrated nasal passages and eyes and led to sneezing. Regardless, this was Ken’s favourite place to be. He and Jeff had played on this court many times. He also played basketball with Roger, but it was better to play against Jeff who was the same age and size as Ken. He wondered momentarily if he or Jeff were taller now.

Ken knew he would have to get the envelope from the hiding place for Jeff because Jeff would have done the same for him, if things were reversed. He took a last shot, then turned out the lights and went to the house to clean up for supper.

Roger was helping to set the table when Ken came in the house. The smells of supper started to make Ken hungry. Roger was talking about his homework with his mother. Mickey was sitting near his empty food dish trying to get someone to feed him. Roger tried to do everything that Ken did, and when they were not around their school friends, the two of them actually worked together and talked about their friends. Ken had once or twice taken the blame for things that Roger had broken or lost. Being younger, Roger inherited all of Ken’s older clothes that didn’t fit any longer. Roger rarely got any new clothes, except shoes, for himself, but he didn’t mind because they were taught not to waste things.

When supper was ready and on the table, Roger filled Mickey’s dish with kibble, then took his place at the table. Together the whole family said grace before picking up their knives and forks. The talk around the table that night was again about the fire. Ken and Roger took turns telling their parents about the possible suspects that their friends came up with at school.

“I’m sure these ideas will be of immense help to the investigator who was there today.” said their father, “But what I’m most concerned about now is what is for dessert?” and they all laughed. To any outsider, this family showed true love and care towards one another.

That night Ken went to bed thinking about the fire, Jeff’s envelope, and how he was going to get the envelope to Jeff once he had it. He did not know where Jeff currently lived, but because neither of them knew how to drive a car yet, the opportunity for them to get together was not apparent to Ken. Ken did not know how this was going to happen without telling his father or mother, and once he did that his parents would have to tell the bank. He put it out of his mind, and decided to not worry about it any longer until he actually had the envelope. With any luck it would be gone, and he would not have to worry about it at

45

all.

46 CHAPTER 3. KEN AND JEFF

The Investigation

October 10, 2012

During her first month in Mount Dunham, Linda simply adjusted to life in a small town. Not much exciting had happened at work yet. She got her uniforms, and managed to scrounge some used furniture for her house. Her cat Ruffles almost became friends with the local skunk. With coyotes or coywolves running around, she kept her cat indoors most of the time. She became good friends with Tracy and Justin, her two main co-workers.

There was a fire in a garage at a house on the north end of town. The fire had been put out in time and no one was injured. Linda, however, in her investigation discovered that the fire had been deliberately set. Even more impressive was the fact that she found a pair of gloves in the notch of a tree near to the garage, that traced to a young boy who lived nearby. When the boy was questioned by Captain Muller, he quickly admitted starting the fire, which then got out of control, and the boy fled the scene quickly. Otherwise the fire may have been attributed to accidental causes. The parents were left to deal with their boy, and to pay for the damages.

Linda also handled the first grade introduction to Fire Safety at the public school. She demonstrated her fire inspection equipment and let the kids play with a roll of Caution Tape, which they used to wrap their teacher to look like a yellow mummy. Linda showed a short film on how a fire begins and spreads. The teacher told Linda that her presentation was much better than last year’s, when Gord Walker did it. Gord apparently brought in a dog that had been burned to death in a recent fire, and several children had nightmares that night.

On October 10, the morning after the fire at the Blanchard farm, Linda got to the Police Station by 7:45 am. As she came in and walked past the lunch room, supply room, and interrogation room, Tracy walked past going to the lunch room.

“Good morning Linda. Captain wants to see you.” Tracy said.

“Good morning. How’re you doing?” Linda said warmly. Tracy had helped Linda find dishes and pans for Linda’s kitchen, and Tracy’s husband helped move some furniture with his pickup truck on weekends. Linda had met Tracy’s daughters and talked with them one night, and read them a story. She enjoyed that night with the kids.

Linda dropped her lunch and jacket on her desk and checked her hair and face in the mirror behind her office door. She noticed a cat hair on the sleeve of her uniform and picked it off.

Don Muller saw her enter his office and he leaned back in his chair, and said,

“Good morning, Linda. Beautiful day isn’t it?”

“Sure is, sir. I heard there was a fire last night. Are we going to investigate it today?” Linda asked.

“I want you to investigate it by yourself” Captain Muller said.

Linda was taken aback. She wondered to herself if she was ready to do this alone. Sort of like a pilot doing their first solo landing of an airplane, without doing any practice runs with a legal pilot. Captain Muller added,

“I was told it was a deliberate fire. A truck was seen leaving the place in a hurry, but no description of the truck or the person or persons in it. The farm was abandoned so no one was killed. You can take the SUV with the gear in it. Sign it out before you leave. When you get back replace any items that you use so that it is ready to go for the next fire. If you encounter any difficulties you can always phone me, and I’ll come out there.”

He handed her a file containing the fire chief’s report from the previous night.

“Sure captain.” Linda said, “Sounds straight-forward.”

“You need to take photos and collect evidence that might lead us to the person who started it, and of course, how they did it.” Muller said. Then he realized Linda should know what needs to be done.

“Good luck with it.” he said.

Linda was on the road within 30 minutes with directions from Justin. Sunglasses shielded her eyes from the bright sun light on this crisp fall day. The sun was in her eyes as she traveled east on highway 12. After 20 minutes she approached a west bound school bus that was stopping to pick up two boys. Linda had to stop while the boys climbed into the bus. Linda looked at the home where the boys lived. The name on the mailbox was Lawrence. The house was set back from the road about half a kilometer, farther back from the highway than most places. The house was surrounded by four mature maple trees, and lilac bushes on either side of the front door. The trees and bushes were bare of leaves already. Several barns were set behind the house. The laneway went in and curved around to the back. There was a maroon van and a red pickup truck parked in front of a shed or garage behind the house. Linda thought how cold it would be in that house during the winter with little to stop the winds. She thought about the isolation and the distances between farms out here. She wondered how cold her house would be this winter.

50 CHAPTER 4. THE INVESTIGATION Linda suddenly became aware that the bus was moving, and that she was now permitted to drive on too. There were no cars behind her so she did not need to rush. As she passed the bus, she waved to the female driver. Beyond the bus, she could see the smouldering house about a kilometer ahead and another SUV in the laneway near where the house used to be. That must be the place, she thought. Linda turned left into the lane when she reached the premises. She drove on and pulled up alongside the other vehicle and rolled down the window.

“Good morning guys. How’re you doing?” she said cheerfully. Linda noticed they both had stubble on their faces, not having shaved yet today, and maybe not yesterday either. She knew their names, but little else about them. She was glad she did not have to follow the daily ritual of shaving her face. She couldn’t imagine the irritation of shaving every day. But for a man, that was their due punishment for abusing females.

“We’re hungry and cold. Do you need us to help you?” said the fire fighter closest to her.

“I should be okay. Go get a breakfast and some sleep.” she said.

They nodded and the driver signaled goodbye with a short salute. The SUV roared to life and they backed up in a U on the grass and drove forward down the lane.

Linda parked her SUV and got out. There was a warmth still encompassing the area where the house used to be. She walked around the house outside the caution tape that encircled the house as a preliminary survey. The frame of a burnt out car was visible in the garage area. The garage was gone except for its concrete base.

“That’s odd. I thought the bank had sold everything from the house. Why is there a car in the garage?” Linda thought. Also visible were frames of a bicycle and a lawnmower in the garage.

“I guess they didn’t sell off everything, just what they could.” she reasoned.

The house was completely gone except for the chimney along the west wall, and the basement into which nearly everything collapsed. Metal plumbing pipes were bent in odd shapes. Water and electricity had been turned off years ago. Most of the debris was in the basement below the living and dining room parts of the house. The kitchen had been an addition to the house along with the garage, both without a basement below them.

When she got back to the SUV, she put on her fireman’s gear, boots, pants, coat and rubber gloves, and helmet, in case some of the debris was still hot. She began another slow walk carrying a camera, closer to the house and garage trying to determine where the fire might have started, but the burn was so complete that there were few clues about where it had been started. She snapped pictures of the car in the garage, the bike and lawnmower. She looked for articles on the ground that may have been used to start a fire. Near the front of the house was a crushed, empty, Sleeman’s beer can. Linda photographed and bagged the can.

A black cap to a plastic gasoline container, melted by the flames, also lay on the ground. She took photos of it and bagged it, as well. Linda returned to the front of the garage. On the ground were metal latch pieces of both doors still joined together by a padlock, but now not attached to any doors. The screws that held the latch pieces to the wooden garage doors were still in place or close by on the ground. Linda photographed the pieces and bagged them. The car had been locked up in the garage. The person that started the fire did not put the car in the garage, otherwise why bother locking it in, she concluded.

Linda concentrated on the car itself now. It was a sports car. There was room for only a driver and one passenger. On the hood of the car was an emblem of a rearing horse. She photographed it and then used her cellphone to google car emblems. The black prancing horse, as she discovered, was the famous Cavallino Rampante trademark of an Italian Ferrari. She was flabbergasted. A Ferrari typically cost a quarter million dollars or more. How could a car of such value be left in a garage of a deserted farm house in the middle of nowhere? The person who started the fire obviously did not know that there was a Ferrari in the garage, it was far too valuable to burn, she thought. Linda saw the vanity licence plates. She could make out the letters “4PAULS” even though the plates were completely black. She phoned her office and asked Tracy to trace the owner of the Ferrari with those plates. Linda had never seen a Ferrari in real life, even one destroyed by a fire. She also asked Tracy to have a flatbed tow truck sent out for the car, said thank you, then disconnected.

The trunk lid was on the ground a few feet away, still within the garage. The small trunk appeared to have been empty. Linda took many pictures. With digital cameras there was no need to be frugal with film. She worked her way around the car imagining what it would have been like to drive around at top speed when it was new. She imagined a bumblebee yellow colour with black leather interior. She found the necessary identification numbers and jotted them down. There were no obvious human remains in the car to preserve or identify, and all other non-metal objects had been turned to ashes and smoke. Linda’s phone chirped.

“Hey Linda, the car was registered to Paul Selinger of Toronto, and reported stolen on July 4, 2011.” said Tracy.

“That was quick. Good work Tracy. I wonder how and when it got here, locked in the garage, I mean?” Linda said.

“That’s your job to figure out. I’ll leave the details about the car and Paul Selinger’s contact number on your desk. The tow truck is on its way, bye.” Tracy replied. Linda was impressed by another example of Tracy’s efficiency at her job.

Linda needed a ladder to descend to the basement part of the house, so she went to the barns to look for one. She took photos of the barn and surrounding scenery, like the Lawrence’s farm, and the house to the east. Linda walked around the barns and saw no ladders she could use. She opened the double doors on the first larger barn and was scared by a rush of pigeons fluttering out in all directions. After they passed, she peered into the darkness of the barn. Sunlight shone through spaces between the vertically attached boards on the side of the barn and onto the wooden floor. A few pigeons were still flitting about in the rafters. About a dozen bales of straw were stacked in the back corner, but otherwise the barn was bare. A small office to the right of the door contained an empty desk, surrounded by shelves holding jars with different sizes of nails, screws, bolts and nuts. There were also oil cans, and used tubes of lubricants. A first aid kit was attached to the wall near the light switch. The calendar in the office showed April 2009. Outside the office were engine belts of various sizes hanging on the wall, and chains, like bicycle chains. A work bench ran along the wall opposite to the office, with an anvil and vice on one end. The smells in the barn were musty with overtones of oil and gas, and mice. Three old tires leaned against a wall.

Suddenly a big black dog ran into the barn, barked, and came up to Linda with its tail wagging rapidly, surprising Linda yet again. The dog jumped up and put its front paws on her stomach and nearly knocked her over.

“Down” she yelled, and the dog backed off still wagging its tail vigourously, panting, and moving around excitedly watching Linda expectantly.

“Where did you come from?”, Linda said, as she went to the door of the barn. Across the field she saw a woman making her way and yelling “Mickey”.

“So, your name is Mickey?” and Mickey renewed his tail wagging at the mention of his name. Eventually the woman made her way to the barn, and Mickey ran to her. She immediately attached the leash on the dog and continued up to Linda with Mickey pulling her along. She was out of breath, and trying to apologize at the same time. Linda took a picture of the woman and her dog while the woman was getting the dog under control.

“I am so sorry. Mickey wants to play all the time. He saw you and had to come over. My name is Grace Lawrence. We live next door there.” she said as she pointed to the west.

“No problem. He startled me is all.” Linda replied. “My name is Linda Logel. I’m a fire scene investigator.”

“Nice to meet you. You must be new to Mount Dunham?” Grace asked.

“Yes, new. Just a month now. This is my first official case actually.”

After a moment of silence, then Linda asked,

“Were those your two boys I saw getting on the bus?”

“Yes, Ken and Roger.”

“Did anyone in your house see anything unusual over here last night?” Linda asked.

“No, I was taking a bath at the time of the fire. And the boys were already in bed. My husband told us about the fire this morning.” Grace answered.

“Did you know there was a car in the garage?” Linda asked.

“Really? I assumed the Blanchards took everything with them. What they didn’t take, the bank sold off at auction. There shouldn’t have been any car in the garage.” Grace said.

“That’s what I was thinking. Not just any car either. It was an Italian Ferrari.” Linda added.

“Goodness gracious. Those are expensive, aren’t they?” Grace asked.

“Quarter of a million dollars or more.”

“Heavens!” Grace was astonished. Mickey whined a little for attention and Grace petted his head.

“Have you seen any strange cars over here since the Blanchards left?” Linda asked.

“No, we can’t see what goes on in front of the house, as you can see from the angle to our house. As far as I know, we haven’t seen anyone over here. If anyone was here at night, we would be less likely to have seen anything. You might want to talk to Mrs Palmer. She lives over there. She’s a retired professor from the university.”

“Thank you. I don’t think I have any further questions at the moment. Thank you for your time. Goodbye Mickey.” Linda was hinting that Grace could leave now.

“Well yes, we better leave you to your investigating, so bye for now. I’ll keep Mickey at our house for the rest of the day.” Mickey and Grace turned and headed home across the field again.

Linda went back into the barn, and was soon satisfied there was no ladder. Linda went to the second, smaller barn and entered. There were no gaps in the walls of this barn and the smells were different. Stale, dusty, grain-like smells lingered heavily due to a lack of air movement inside the barn. A calendar near the doorway indicated June 2009. There was an old wooden ladder hanging on the wall. Although heavy, Linda took the ladder to the house to complete her investigation.

She eased the ladder over the side down into the basement in a somewhat clear area, and climbed down into the basement carrying a small pick axe. As she moved carefully through the mess, she turned over burnt boards and debris. Puddles of water were on the floor. From the way the burnt flooring was slanted into the basement when it fell, the fire appeared to have been started in the centre of the living room and worked its way towards the outer walls. There was no charring on the basement walls, so that the fire was mainly on the upper levels of the house which caved into the basement. Part of the stairs to the basement were not burned completely. Linda snapped pictures with the camera. Satisfied she had searched everywhere, she climbed up the ladder and pulled it out of the hole.

The kitchen area was totally burned except for the ceramic sink which had cracked. The kitchen led to the garage. Being a wooden structure, and being empty of all household items, the house had burned quickly and completely. Gasoline seemed to have been poured in the centre of the living room with a trail of gasoline to the front steps. A melted gasoline container was inside the house in the kitchen area. She took more pictures.

She carried the ladder back to the second barn and hung it back on the wall inside. Her watch indicated 11:06 am, and Linda returned to the SUV to have a fruit snack and to write down some notes. She made a list of items she had collected. Beer can, gas container cap, Ferrari, and lock pieces from garage door. She relaxed and ate a power bar and orange slices while sitting on a flat rock near the house. She saw the tow truck coming down the road, and packed away her snacks into the SUV, and went to greet the driver.

“Hi, my name’s Al” said the driver.

“Hi Al, I’m Linda Logel.” she said, displaying her badge ID to him.

“There’s a car in the garage we need to take back to the fire station. You can back up over here.” Linda removed a section of the caution tape. Al backed the truck up to the garage entrance. He lowered the back of the truck bed until it hit the ground, so the bed was at a slant. He pulled the wire cable from the winch towards the garage and found a spot on the front of the car to which he could attach the hook. Then he looked more closely at the car.

“Hey, this was a Ferrari!” he nearly fell over. “Not so much now, is it? What’s it doing here?”

“That’s what I have to figure out. Have you seen one before?” she asked.

“Yeah, I did. About a year ago, in Mount Dunham.”

“Mount Dunham, really? What colour was the car?” Linda asked.

“I remember it was grey or brown, not flashy red or yellow like you usually see in ads.”

“Did you see the driver?”

“Not that I can remember. It was parked at the Black Forest Restaurant one evening.”

“When was that? Summer, winter?” Linda queried.

“I think it was summer of 2011.” Al answered. Al started the winch motor which squealed loudly, and the car started sliding through the rubble, effectively stopping further discussion with Linda. About five minutes later the car was on the truck and the bed back to its horizontal position. Al retrieved the trunk lid from the garage and tossed it inside the frame of the car. Then he secured the car to the bed with chains across the top of it. When he finished, he climbed back into the truck. Linda started talking, to herself, or to Al,

“I’ve never seen a Ferrari before in my life, and neither have you until last summer. This Ferrari and the one you saw, most likely, are one and the same. The chances of having two different Ferraris in Mount Dunham within the last year must be very low. I know this Ferrari was stolen on July 4, 2011 from Toronto. Therefore, you must have seen it on or after July 4, and the people driving it must have been the thieves. But they realized that a Ferrari was too visible and people noticed it. So they hid it in the garage of this abandoned farmhouse. They must have planned to return for it later after police gave up looking for it. So where are the thieves now?”

Al did not answer, but indicated he was ready to leave. Linda was disappointed that Al didn’t react to her logic.

“Well, Al, what do you think about what I just said?”

“Uh, sounded correct to me. I wasn’t listening carefully. I did see another Ferrari at a car show in London, also last year. That Ferrari was brilliant red. There is a small chance, you know, that the Ferrari on my truck and the one I saw at the Black Forest are different cars.”

“Thanks Al” she said. Linda waved as he drove down the lane. She realized that she could not assume that the Ferrari in this garage and the one at the Black Forest were the same. She would have to test the theory.

Al had said more than he usually did. Al was a known recluse who did not interact very much with anyone, unless he had to. Customers who needed his towing service did not find him very friendly, and Al generally did not answer their questions or mundane comments. Today, he did not pay any particular attention to Linda, but just kept to his work. The Ferrari had excited him a little. Linda made a note in the file that Al had seen a grey/brown Ferrari in Mount Dunham in the summer of 2011 at the Black Forest restaurant, and a different Ferrari in London.

Her work at the site of the fire seemed to be finished, so she took off her fire gear and packed things up. The sun had made things warmer now, but the sky was still clear of clouds. The caution tape had to be left up because of the dangerous hole made by the house. She decided to leave danger signs around the house. She spaced four signs around the house. After, she finished off her snacks, and she decided to visit Mrs Palmer.

Linda drove the SUV down the lane, turned left onto highway 12, then turned into Mrs Palmer’s laneway. The laneway was wide enough for one car, so she had to park behind Mrs Palmer’s Honda CRV. As she was getting out of the SUV, Mrs Palmer came out of the house with her purse, obviously planning to go somewhere. She was startled to see Linda and her SUV parked behind her car.

“Mrs Palmer, my name is Linda Logel. I’m the fire scene investigator. Do you have time to answer a few questions?” Linda noticed Mrs Palmer’s purse and that it was a very expensive brand. Mrs Palmer was dressed very well, as though she were going to an opera. Linda could smell the perfume.

“I’m afraid not. I have an appointment in Mount Dunham in 30 minutes at the hair dressers. I have to leave now.”

“Could you stop by the Police Station and see me after your appointment then?” Linda asked.

“Yes, I can do that. What was your name again?”

“Linda Logel.”

“Okay, could you please move your vehicle so I may leave?”

Linda got in her SUV and backed out the lane and onto the road. Mrs Palmer followed, then headed for Mount Dunham. Linda watched as the CRV sped away. Linda looked at Mrs Palmer’s house and noticed that a person could easily see the laneway to the Blanchard place and the area in front of the house. Linda took some pictures of the Blanchard place from Mrs Palmer’s driveway. The distance was too great to tell one person from another, but you could see people moving about.

Linda made some more notes in her notebook then drove off towards Mount Dunham. She looked at the scene of the fire as she passed, and then at the home of Fred and Grace Lawrence and their two boys. She could see Mickey curled up against the side of the house in the sun. She thought, “I bet Mickey could tell me if someone had been in Blanchard’s house. If only he could talk.” She turned on the radio and listened to music the rest of the way back to the office.

Back at the Police Station, Linda cleaned the gear she had worn that morning and the pickax she had used, and replaced the plastic bags that she had used for evidence collection. She also stocked the SUV with more caution tape and danger signs from the supply room. Then she took the camera and evidence bags into the station. She asked Tracy to download the pictures and send them to her so she could label them. Linda signed in the SUV.

On her desk was the information about the Ferrari. She checked the vehicle identification number (VIN) with the numbers she had written down at the scene and they matched. She added that confirmation to her notes. She noticed that the colour of the vehicle was brown-grey, which was the same as the car that Al, the tow truck driver, had seen last summer at the Black Forest Restaurant.

Linda started a new page in her notebook and listed things she needed to check. First, she had to contact Paul Selinger about his car. Second, she needed to talk with Mrs Palmer. Third, she needed to go to the Black Forest Restaurant to see if anyone remembered seeing a brown-grey Ferrari in 2011. Lastly, she needed to get the beer can she had collected analyzed for prints and DNA. She sat at her desk looking out the windows at the sky, apparently thinking. A lone cloud floated by the window on a clear blue background.

Suddenly her computer beeped indicating she had received a message. It was the pictures from Tracy. She started through them and labelled them one by one, and stored them in a folder for this case. She sort of graded each picture on its clarity and usefulness to the investigation. She sent the labelled pictures to the printer for her physical files. She got up and went back to the lunch room to get a juice from the vending machine. She sat at the table in the lunch room and drank her juice. Justin came in to get a coffee.

“Justin, do you remember my first day here?” Linda asked.

“Anything in particular?”

“I manned the front desk for you while you went on a break.”

“Yeah, and I had to scold you for not recording things in the computer.” he smiled.

“Besides that, do you remember there was a man that came in to ask about the Blanchard place?”

“Yes, but you didn’t get his name, just a picture, remember?” he said.

“Do you think he could be the one that set the fire last night?” Linda asked.

“Could be, we’ll never know, will we?”

“What about the truck in front of Wilhelmina Schrandt’s house? Did we ever get an ID on the truck or owner? The reason I ask is because a dark pickup truck was seen leaving the scene of the fire.” Linda wondered.

“I think the truck was gone by the time Officer Beechum arrived at Mrs Schrandt’s house, but I’ll check on that.”

“I’d appreciate that, please.” Linda said.

“Sure thing.” he said, and he left the lunch room to go back out front.

The loudspeaker announced that Mrs Palmer was here to see Linda. So she finished her juice and discarded the can in the recycle box, and went out to the front desk.

“Thank you Mrs Palmer for coming in. Follow me this way please.” Linda led her to the interrogation room, after grabbing her notebook and pen from her office.

“Please be seated Mrs Palmer. I have just a couple of questions about the Blanchard place that you might be able to help me with. How long have you been living in your present house? I understand you are retired. Are you married?”

“I retired in January 2011, so I have been living here for a year and 10 months. Before that I lived in Guelph, while I was working. I am not married, my husband passed away in 2007.” she said.

“Why did you pick that house? You are alone and out in the country.” Linda asked.

“Oh, that was the house I was born in. When I retired, it came up for sale, and I thought it would be good to go back home. So I’m used to living in the country. The Lawrences look after me, somewhat. Whenever I need help, they usually oblige.” Mrs Palmer answered.

“I see.” said Linda. “So you moved in after the Blanchards had abandoned their house?”

“That’s correct.”

“In the time you have been living there, have you ever seen other cars or people at the house?” Linda said.

“There was a car there once during my first year. I didn’t know the house was totally abandoned at first, so I thought it might have been the previous owners. The car was only there for a few hours one morning, then it was gone. I saw a man and a woman leave in the car.” she said.

“Did you see what type of car it was?” Linda asked anxiously.

“It was a white sedan type of car, four doors I believe. I don’t really know what brand of car it was.”

Linda wrote white sedan, and two people, a man and a woman in her notebook. “That is excellent, Mrs Palmer. Did you notice anything else about them? Could they have been there overnight or longer?”

“It’s possible they could have been there longer, but I did see them leave that morning and never again.”

“Do you recall the exact dates for that?” Linda asked.

“Dear me, no. Sometime in 2011. They were not wearing any coats, so it must have been in the summer. Since I’ve retired, time goes by rapidly and I don’t pay much attention to it.”

“Don’t worry, you have been a great help. Thank you very much for coming in. I hope I haven’t taken up too much of your time. Your hair looks lovely, by the way. Where did you get it cut?” Linda said. Mrs Palmer gave Linda the name and address of the hairdressers, then Linda took her to the front exit.

“I couldn’t help but notice your purse. Is that a real Chloe style?” Linda asked.

“Yes, it is real. Most people don’t notice. Well, good bye.” said Mrs Palmer and she left the station. Linda did not know many women who would spend $2500 on a purse, and take it to a hairdressers. Linda turned to go back to her office, when Justin stopped Linda and handed her a slip of paper. On the paper was a description of the truck and licence plate number of the truck that was parked in Mrs Schrandt’s driveway.

“This was the info on the truck parked in front of Wilhelmina’s house that day.” he said.

“Oh, thanks Justin.” she said and took it back to her office. She looked at the slip and went to her computer. She logged on the motor vehicle system and typed in licence plate number. Up came a picture of a man called Nate Haslett. He was clean shaven with short hair. The licence plates brought up a Ram truck, black. The address for Nate and the truck were RR 5, in Mill Valley, about 40 km from Mount Dunham. She printed out everything that was there. Linda went out to Tracy’s desk where the printer was. Tracy was watching the pages come through. She picked up the pages and was glancing through them.

“How did you find him?” Tracy wondered.

“What do you mean? Find who?” Linda asked.

Tracy took her pen and drew a moustache and goatee on the face of Nate Haslett. Linda recognized the man now as the one who asked about the Blanchard place.

“He’s the owner of the truck that was parked in front of Wilhelmina’s house on my first day here. But you’re right, he’s the one. Wow, he could be the one who started the fire. And he has a black pickup truck. There was a dark coloured pickup truck seen at the fire. Is this a coincidence, or what?” Linda said.

“All you have to do now is find this guy.” Tracy said.

Linda was smart, but she was also lucky. Now she had a suspect.

64 CHAPTER 4. THE INVESTIGATION

Paul Selinger

Year 2009, Toronto

On a busy weekend night at the Terrozi Bread House in Toronto, Head Chef Paul Selinger was about to collapse after a long night of preparing meals. He had yet to check on his supplies and make the food orders for the next few days. There was usually sufficient food for several days, but some items had to be absolutely fresh and it was good to keep the stores topped up. Also, the ability to get certain types of food, depended on the season of the year and general availability. Bad weather could cause a shortage of some items. His sources had to be informed of his needs a few days ahead of schedule.

At that moment a waiter came into the kitchen, and said

“Chef, there’s a customer that would personally like to compliment you on the meal. Believe me, this will be totally worth your while. This babe is an angel.”

“That’s what you always say.” said Paul.

“This time it’s so true.”

Paul got up and walked to the door to the dining area. The waiter led him out to the table and the lady stood up. Paul gasped when he saw her because he rarely saw such a beautiful face and smile. Her light gray eyes skewered him through his heart. Her eyes were mesmerizing, he couldn’t look away, they were very unusual. Her long blonde hair was graceful and silky. Her body was perfectly proportioned all over. Her perfume was intoxicating to him. She was indeed an angel as the waiter had said. She was obviously younger than he. She started to speak,

“That was an absolutely magnificent meal. The lamb chops melted in my mouth. I just wanted to tell you personally that I enjoyed everything.”

Paul was still looking into those eyes, paralyzed, but knew he had to respond. He didn’t seem capable of forming any words or sounds. She was smiling now with lips that were also unique. He pulled himself together as much as he could, and said,

“Thank you very much, it was my pleasure. Paul Selinger at your service.” as he took her hand gently, bowed and kissed it lightly. He thought this was weird given he had never bowed or taken anyone’s hand before. What was happening here?

“I’m Jackie.” she said. She seemed to be as attracted to Paul as much as he seemed to be smitten with her. She had been expecting an older, more plump person with beard and moustache. She had not prepared herself for a younger man, his height, slimness, and good looks. Instead of a beard he had a ponytail. She was still holding his hand and squeezing back a little.

“I’m new to Toronto, and I have never had such a splendid meal.”

“Well, thank you again.” Paul replied, then a wild thought came into his head and he said hopefully, “I could show you around the city tomorrow, if you like, it’s my day off.” Again, something he had never done before in his life. He was throwing himself at this woman. Take me, please, he was thinking.

“As you see I am here with my friends.” she said as she pointed to two people walking towards the exit.

Paul’s mind had not registered the existence of a man and a woman who had also risen from the table at the same time as Jackie. Jackie took a pen from Paul’s breast pocket and wrote a phone number on the sleeve of his shirt.

“Call me tomorrow and maybe we can plan something.” she said. Paul was still looking at those eyes. But then she turned and was out the door and out of sight. He didn’t remember what he was doing before he met her. He had never fallen so quickly for any girl before. There was a connection between them, from his point of view, the two of them had to be together. It felt right, almost perfect. He had to get to know her better. He could not get her out of his head for the rest of the night.

The next morning he showered and spent more than his usual care in cleaning himself for the day ahead. He was getting anxious to see her again. Maybe he was in a weakened state when he met her last night. Maybe she wasn’t as beasutiful as he originally thought. He ate some breakfast and read the newspaper, then he went to the bathroom and brushed his teeth again. At ten o’clock he gathered his courage to call the number that Jackie had written on his shirt sleeve.

“Hello.” said a male voice on the other end. Paul was surprised, then asked,

“Hi, I’m calling for Jackie. Is she there?”

“Just a minute, I’ll get her.” he answered. At this point, Paul was wondering whether this male person was a boyfriend, a relative, or a friend. Then he remembered her friends at the restaurant, and concluded she must be staying with them. In the background Paul heard a female voice ask who was on the phone, and the male voice replied that the call was for Jackie. In what seemed like an eternity, Jackie finally picked up the phone and said,

“Hello?”

“Jackie? This is Paul Selinger.” No response came immediately, then he said, “The chef from the Terrozi Bread House. We met last night, remember? After your meal.”

“Oh yes, I remember. I gave you my phone number. That was really a tremendous meal. What can I do for you?”

“I wondered if you wanted me to show you around town today.”

“Oh yes” she remembered, then she said, “Give me a minute to talk to my friends, I have to check what they have planned. Just hang on.”

Paul waited while he heard discussion in the background. He began to wonder if he was making a mistake. She came back on the line.

“Hi Paul? My friends were going to take me around today, but they said that you could come along. Would you come with me please?” Her voice was sweet and pulled him in. He would have preferred to be with her alone, but this might work out too. Paul agreed and they decided to meet outside of Paul’s restaurant in an hour.

An hour later Paul was in front of the Bread House and a silver Honda Accord pulled up. Paul spotted Jackie in the back seat and she motioned to him. She was as spectacular looking as she was the night before, only more casually dressed. She was in jeans and a black and white horizontally striped blouse that emphasized her physical attributes. He slid into the car, and away it went. After he put on his seat belt, Jackie put her hand on his, and it felt good, warm and soft. Paul was surprised, but he didn’t object. Jackie introduced her friends. The driver was Steve and the woman was Judy. Steve was brown haired with a moustache, shorter than Paul, dressed in a coat and tie. Judy was dressed very elaborately, with long black hair and heavy makeup. She was about the same height as Steve. Paul thought Judy was pretty without all of the makeup. Steve and Judy were dressed very formally compared to Jackie and him. Judy’s light tan dress was a stunning business type of dress which made her look commanding of any situation.

By comparison, Paul had on jeans and a sweater. Jackie had only a little makeup.

“Steve and Judy are getting engaged today. We’re going to the Eaton Centre so that they can buy an engagement ring.” said Jackie.

“We decided last night after dinner at your restaurant.” said Judy.

“Congratulations!!” Paul exclaimed. “How long have you been together?”

“Almost five years,” Judy answered, “and now I’ve finally got him off the pot.”

Steve pulled into a parking garage near the Eaton Centre and parked. Jackie and Judy made plans to meet later, then Jackie took Paul’s hand again and off they went to look at the stores in the mall.

“I’m usually an outdoors kind of person, not a shopper.” Paul said.

“Me too.” said Jackie, “But Steve and Judy were set on getting some jewelry today. We have about an hour until we meet them again. Where do you want to go?”

“I know a park not too far from here, if you like, we could take a stroll through it.”

Jackie smiled and squeezed his hand. Then she suddenly grabbed him with her other hand and kissed him on the lips. He was dazed, and thought “That was nice.” Paul was ready to follow her anywhere, there was a funny feeling coursing through his body like a drug taking effect. He didn’t remember how, but they were suddenly in the park on a bench with a small ice cream cone each. Paul and Jackie were talking freely about many topics and Paul didn’t seem to remember any of the conversation. He was entranced by Jackie’s light gray eyes, smile, and awesome body. To Paul it seemed like he had known Jackie for a long time. He was comfortable being with her. Time passed quickly. He wanted to stay with her the rest of the day.

“Uh oh.” Jackie said after looking at her watch. “Judy’s expecting us at Brogan’s Jewelry store in the centre. Quick, we’ll have to run.” Together they started jogging down the sidewalk, but the other pedestrians slowed them to a fast walk. They made it to Brogan’s in time. Steve and Judy were inside looking at necklaces and rings with a clerk. Paul was about to go in and ask Steve if they had found anything yet, when Jackie took his hand and pulled him back.

“Let’s pretend that we don’t know them and work our way around the store to them.” she whispered in his ear. Paul agreed, but did not understand why. Jackie drew Paul’s attention to some watches on display on the opposite side of the store from where Steve and Judy were standing. All of a sudden Jackie screeched with joy, jumped and threw her arms around Paul’s neck and gave him a big kiss and held on. Paul was totally surprised and put off balance by Jackie’s lunge at him. He fell backwards against the counter rather hard, jarring it and causing a blaringly loud alarm to go off. The loud ringing noise made Paul cringe as his ears were assaulted. The clerk came running over and Paul was quite embarrassed and held out the palms of his hands in apology to the clerk. Jackie also acted embarrassed. The alarm made it impossible to talk in the store. The clerk motioned for them to leave the store immediately, as she searched for the off switch. As Paul and Jackie turned to go, the clerk noticed that Steve and Judy had disappeared. Then she saw that the necklaces and rings she was showing to them were also gone and she left the alarm ringing. Paul and Jackie hurried away from the store. Security police ran past them towards the store. A crowd gathered outside the jewelry shop.

Jackie grabbed Paul’s hand and they hurried quickly away. Paul realized that he had just been an unknowing accomplice in a jewel robbery that had been planned by Jackie, Steve, and Judy. Jackie had lured him in with her hand holding and kisses. She was still holding his hand and pulling Paul along behind her through the Eaton Centre. They finally came to a bench and Jackie sat down. Paul stared at Jackie, and she said,

“Don’t say anything right now. Wasn’t that exciting?” She was still holding his hand, and it still felt good to him. Her eyes were wide and bright with excitement, she was breathing fast and smiling.

“Do you do this all the...” Paul began, but was stopped by

“Shhh! Later.” she corrected him.

A couple was suddenly standing next to Paul and Jackie, and Paul did not recognize them. There was a man with brown hair and glasses, dressed in slacks and a sweatshirt. The woman had short red hair and sunglasses, in slacks and light purple blouse. The man spoke, “Time to get out of here.” Then Paul recognized the voice and blinked in astonishment at Steve and Judy. He couldn’t believe the transformation in their appearances. The group including Paul, proceeded to walk out of the Centre and to their car in the parking garage.

Once in the car and on the road, Steve said to Paul, “Thanks for your help back there. Wasn’t that fun?”

“But you robbed a jewelry store.” Paul said incredulously.

“No one was hurt.” Judy said.

“But why?” Paul stuttered.

“It gives us a rush.” said Jackie.

“You’ve done this before?” Paul asked, bewildered, and wondering what he had gotten mixed up in.

“We find it challenging and exhilirating. Gets the adrenalin going.” Steve put in.

“And it pays some bills.” added Judy.

“Have you ever been caught?” Paul asked.

“Not so far.” answered Judy, “Do you think we paid for these clothes today?”

Paul looked amazed at Jackie. She smiled, then leaned over and kissed him again on the lips. Paul thought how sweet those lips were and could feel her body against his. He could also feel her hand on his leg moving to his crotch. He couldn’t and didn’t want to stop it. She was so pretty. The robbery and the getaway were actually stimulating.

The car came to a halt in front of an apartment building. Everyone got out and walked in. They climbed stairs to the fourth floor. Halfway down the hall, Steve turned and opened a door. The apartment was not at all what Paul was expecting. The interior gave the impression that it was occupied by very old people. There were frilly curtains, knitted blankets, a tea cozy, and a crucifix on the wall in the living room. Jackie took Paul’s hand again and led him to a bedroom and shut the door. She proceeded to undress him and kiss him at the same time. She was like an animal out of control. Paul did not object. but reacted in kind. The sex was so intense. He forgot about the robbery. He forgot about where he was.

During a break in the sex, Paul could hear Steve and Judy in another bedroom apparently doing the same thing as he and Jackie. This was not the kind of day that Paul had expected when he woke up. He gazed around the bedroom and noticed some pictures of little children. He wondered who they were. Jackie then climbed on top of Paul and started kissing him again. She was unsatiable he thought, but it was great. The night came and eventually everyone in the apartment fell asleep.

The next morning Paul was awakened by a knocking on the bedroom door. He looked next to him on the bed and Jackie wasn’t there. Then he said,

“Come in.” in a singing like voice, assuming that it was Jackie. To his surprise, the door opened and in came two uniformed policemen.

“What’s going on.” Paul asked in bewilderment.

“That’s what we would like to know.” said one officer.

Behind the policemen were an elderly couple totally surprised by Paul’s presence in the bedroom.

“Get dressed sir and come with us.” said the other officer.

Paul dressed and went out to face the police. He learned that the apartment belonged to the elderly couple, and that Jackie, Steve, and Judy were not there. Again, it dawned on him that he had been set up by Jackie. It took a long time to explain his presence in the apartment, and even more difficult to explain the presence of the key to the apartment in his pant’s pocket. Paul felt anger and rage at being so easily betrayed by a woman that he did not really know. It just took a little sex to get him to do what they wanted. Luckily it did not take long for the police to establish that he was indeed Paul Selinger and that he was a chef at the Terrozi Bread House. Paul described his ’friends’ to the police, but did not tell the police anything about the robbery, that would have made explanations more difficult for him. Finally, they let Paul go, and Paul apologized to the elderly couple.

Two weeks passed and Paul did not hear anything from Jackie, if that was her real name. The phone number she had given him at the Bread House was no longer in service. He was still upset that she had disappeared and left him alone in that apartment. He did see a small article in the newspaper about the robbery at Brogan’s with a picture of Steve and Judy in their disguises, in the dark background were he and Jackie, but the faces were obscured in darkness.

He remembered Jackie, her taste and smell, and those bewitching eyes. He longed to see her again. After two weeks, however, Paul was beginning to accept that he would likely not see Jackie again. Paul was at work in his kitchen at the Bread House planning the meals and cooking schedule for that evening. The galley doors opened and someone yelled,

“Paul, visitor out front.”

Paul went through the galley doors to the empty dining area. At the entrance to the restaurant was Jackie. Paul was leery at seeing her, but Jackie went right up to Paul and took his hand and smiled.

“I’m sorry we left you in the apartment. I had no choice. Steve and Judy made me leave. I didn’t know it wasn’t their apartment.” she said.

“I don’t think I can trust you. You’re going to get me into big trouble. Why didn’t you find me sooner?” Paul asked.

“Steve and Judy wanted to lay low until the robbery had been forgotten. We went somewhere out in the country. There was no electricity or phone connections. It was like camping out.”

Paul was doubtful of this explanation. Jackie squeezed his hand then said,

“I had a great time with you after the robbery. I’ve missed being with you.” she licked her lips enticingly. “I need a place to stay. Steve and Judy have gone to Vancouver. I was wondering if I could stay with you?” then she leaned into him and kissed him.

Paul’s knees buckled a little from the kiss. Jackie’s eyes were hypnotizing him again. Paul asked “For how long?”

“At least a year, maybe more.” she replied, giving him a little smile and a wink. Paul weakened more, she was too beautiful to refuse.

“Okay, but you’ll have to come back at closing time. Then I can take you to my place.”

“Don’t worry, I’ll be waiting for you when you get home.” Jackie kissed him again and stroked his crotch with her hand. Paul looked stunned, but was not surprised that she had found his apartment, and that she assumed from the start that he would agree. She knew her power over him, he thought. He was angry with himself for being so weak. Paul knew that Jackie was probably setting him up for another fall. He decided to be happy to have her as long as she wanted to stay. Really, a year?

The Ferrari

A little past midnight Paul left the Terrozi Bread House for home. Jackie was waiting for him at his apartment.

“I brought you something to eat. Lamb chops.”

“I ate something already. I’m not hungry now.”

“That’s alright, I’ll just put them in the fridge.” he said.

“I am hungry for something else, however.” she said letting her eyes look in the direction of the bedroom. She took his hand and pulled him towards the bedroom. The sex seemed to just continue from the day of the robbery, as though they had just taken a short break.

As she promised, Jackie stayed with Paul for a year. It seemed like they were always in bed when he was home, not that he was complaining. Jackie always had something new to try out on him, something to intensify the sex.

Jackie occassionally brought home new clothes, or something they needed around the apartment. Paul did not ask how she got the items, but he knew she did not have a job. Paul concluded she was a thief, and was good at not getting caught. One day they went to a store to return an item. Paul was unknowingly used as a decoy again while Jackie made off with a top of the line food processor. He got angry at her for 76 CHAPTER 6. THE FERRARI

using him as an accomplice. So she stopped involving him in her little robberies. Each robbery gave Jackie enough adrenalin for both of them. He did not mind the sex sessions afterwards.

Jackie’s partners in crime, Steve and Judy, were never mentioned and Paul did not ever see them again. He believed they were in Vancouver as Jackie had said. Until one day he saw them in Toronto walking through the Farmer’s market while he was making orders for the Bread House.

“I saw Steve and Judy at the market today.” Paul told Jackie. “Did you say hello?” she asked. “No, I don’t think they recognized me.” “I’m sure they did, they see us together fairly often.” she said. “What do you mean? I thought you told me they were in Vancou-

ver.”

“Oh yeah, well they came back.” Jackie said. “Have you been pulling robberies with them again?” Paul asked,

knowing the answer was probably affirmative.

“Maybe, it’s better if you don’t know about it Paul.” she said. “You guys are going to get caught one of these days, and I’ll probably

be arrested with you.” Paul said.

“You have alibis for all of the robberies, you were at work.” she said. “Look around this place. Most the things here, you and your gang

have stolen. I could be arrested for having stolen property.”

“Don’t worry Paul.” Jackie said. “I worry about you every day.” Paul said. “I was going to take you out today.” she said. “When?” “How about now?” she said, and she took his had and led him down the street to the bus stop. They went to the subway station and changed from bus to subway.

Paul was surprised when they stopped at a Toyota dealership. From there they went to a Honda dealership, then Volvo, then BMW, then Mercedes. Finally they ended at a small, import auto dealership located in a vacated parking garage. There were Lamborgini, Ferrari, Porsche and other luxurious sport cars on display. Jackie almost began moaning when she saw the cars, and wanted to sit behind the wheel of each one. She had on a fashionable dress and leather boots that came up to her knees, and she wore an expensive royal red coat with a small trim of black mink fur on the collar. Paul didn’t know when or how she acquired this outfit, but it made her look very foxy.

Jackie picked a gray-beige Ferrari, or greige, as she decided to call it. When they test driving it on the 427 headed to the 401, Jackie pulled up her dress to expose her panties to Paul. Paul noticed and squirmed a little in his seat.

“You realize that this car costs almost 300K.” Paul said. “You realize I don’t have anywhere near that amount of money saved up. I could buy a house for that kind of money, albeit a small one.”

“Relax, enjoy the ride Paul.” Jackie smiled. She zoomed around a truck and several cars as though they were parked. She was high on adrenalin again.

“Do I get to drive?” Paul asked.

“Happy Birthday.” Jackie replied.

Paul’s jaw dropped and he wondered how she had determined it was his birthday. Paul hardly ever acknowledged his birthdays to himself. To him, it was just another day in his life. Jackie finally gave Paul a chance to drive the Ferrari, and it was good. People noticed the car, and heads turned at stoplights. After returning to the dealership, Jackie concluded the purchase of the Ferrari, and paid the full amount from her purse, in cash. Some weeks later Jackie got vanity plates that read “4PAULS”. Each day they would take a drive around Toronto, and Jackie would drop Paul off at work. The car made Paul forget about Steve and Judy. He enjoyed the car, and he was in love with Jackie. He cooked fine meals for her. They went places together, movies, plays, concerts, hockey games. As long as Paul ignored her compulsion to steal, he was unconcerned. He most of all enjoyed her endless sex drive.

Then the inevitable day arrived in early July 2011, Jackie and the Ferrari did not come to pick Paul up from work. Paul waited an hour and phoned the apartment several times, with no answer. Finally, he called a taxi. Paul searched the apartment when he got home. Jackie was not there, and her clothes and bathroom items were gone along with a suitcase. Somehow Paul expected this day to come, from the day Jackie moved into his apartment. When she had not returned by the next morning, Paul called the police and reported his car and girlfriend as missing, July 4, 2011. Paul was worried, he didn’t know what to think anymore. He was unable to sleep. He was worried about her. Something must have gone wrong on their latest robbery. The second night without Jackie, he cried.

A week turned into two, then a month. There was no call or news from Jackie. She had disappeared completely. Paul slowly accepted she was gone, but it hurt. He cut off his ponytail, and shaved his head. He rented a car, not a Ferrari. Once in a while, the police contacted him about possible leads. Paul would drive to where she was reported to have been seen, and spend a day or more looking for Jackie. He drove to Ottawa, Kingston, North Bay, Windsor, or wherever the trail led. He realized during these long drives that he did not really know Jackie very well. Her last name was Shambeau and she told him she was from Quebec. Other than that he knew nothing, not even her birthdate. It was possible that Jackie Shambeau was not her real name, and that she was not from Quebec. She never gave him any clues. He knew nothing about her family, or her home life. He couldn’t contact anyone about her. There was no one to contact. All he knew was the features of her body, her smell, the touch of her skin, her ticklish spots. He thought Jackie was basically decent inside, even though she did steal things, as a hobby. Jackie made him feel alive. She was the sexiest person he had ever met.

By June 2011, Steve and Judy had split up. Judy had started seeing other men and Steve found out. They argued about it and Steve had given Judy a black eye. Judy was furious and threatened to report Steve to the police. Judy moved out of their apartment, but Steve followed her and would arrive at Judy’s new apartment unannounced. More arguments ensued and Steve would become more angry with Judy. Finally Steve had an encounter with one of Judy’s dates and broke the date’s hand. Steve would have turned his anger on Judy, but she pulled out a gun and forced Steve to leave.

As Steve was returning to his apartment that night, he spotted police sitting in an unmarked car, and another policeman near his apartment entrance. He cursed Judy for turning him in, and got out of there without being noticed. He phoned Jackie,

“Hello.” Jackie said.

“Jackie, Judy has turned me in to the police.” Steve told her.

“You didn’t hit her again, did you?” Jackie asked.

“Not her, her date. What are we going to do Jackie? I think Judy may have squealed on us, you and me.” Steve said.

“You idiot, see what love does to things. Why couldn’t you just walk away from her?” Jackie said. “We’ll have to find an abandoned farm in the country, like we did before, and stay there for a couple days.”

“What about the jewels our last robbery?”

“I’ll bring them along. We’ll get a safe and store them somewhere.” said Jackie.

“How long do you think we need to hide out?”

“I don’t think we can go back to Toronto for a couple of years.” she said.

“Where are you going to go?” Steve asked.

“I have a sister in Owen Sound. I think I’ll live there for awhile. You could probably get lost there too. We have to stop robbing shops for awhile. I’ll find a wealthy older dude to live with probably.”

“What about Paul?” Steve asked.

“I have to let him go. For good this time. If I see him again, we’ll be caught by the police. I can’t let him find me.” Jackie said. Steve thought Jackie was sadder than she let on. Jackie was thinking about all of the work they had put in on disguises and techniques; how they watched stores for days or weeks before attempting a robbery. What type of security did the store have, how many people worked there, did they have quality merchandise, what volume of business did they normally have, what were the escape routes. They became experts. Sometimes they would test the security systems to see how long before the police responded.

“He sure fell for you. You could get him to do anything for you.” said Steve.

“I can get any man to do what I want him to do.” she replied.

“Yes, you are beautiful and very sexy.”

“Shut up. I’ll pick you up in an hour, meet you at the Yorkdale Centre.”

Jackie hung up the phone, she was angry with Steve for letting love get the better of him. She went to the bedroom and started packing her things into a suitcase. She knew all along that one day she would have to leave and probably at a moment’s notice. She was not in love with Paul, but she was attached to him and the security he provided. Now she was not safe and there was nothing Paul could do to help her. But she had survived on her own before Paul, and she could do so again. She could not bring herself to write a note to say good bye, or to explain. Just cut and run was her way of dealing with relationships. No fuss, no emotion, just business, her business. She looked around the apartment, grabbed the jewels and the keys to Paul’s Ferrari, and left for the last time.

October 11

Thursday, October 11, 2012, Owen Sound

Jackie sat on a stool in an all white kitchen eating an English muffin with strawberry jam and reading the morning newspaper wrapped in a royal red bathrobe. After wading through the boring political news on the first few pages, there was a story about a Ferrari found in the remains of a fire at an abandoned farmhouse near Mount Dunham. Her eyes widened and she read the story intently. The fire occurred two nights ago. The house burned down completely. It was a deliberately set fire, no suspects, no fatalities.

“Oh no” she moaned to herself, “that beautiful car ruined. I guess we left it there too long.” She re-read the story again more slowly. There was no mention about the barns or the discovery of anything else. That meant that the safe of jewels that she and Steve hid must still be there in a barn. Then she read about the part where the house and property were to be auctioned off in two weeks.

“We’re going to have to get the safe.” she said to herself again.

“What safe, sweety?” said Daniel. Jackie nearly jumped from her seat because she hadn’t heard her current boyfriend come from the bedroom. He was naked and groggy looking.

“Just talking to myself honey.” she replied. “Awww, look at that, the little guy is all petered out.” she said as if talking to a baby, looking at Daniel’s genitals.

“Don’t say that, you’ll hurt his feelings.” He came up behind her and reached into her bathrobe. She wore nothing underneath.

“Oooooo, your hands are cold.” she shuddered slightly.

“I’m trying to warm them up. What’s in the news?”

“Same old stuff, politics, riots here and there, Mayor Ford, Raptors lost again, hockey lockout still on.”

She had not told Daniel about her past, but he knew that she was a thief. Like boyfriends before him, her other attributes always outweighed her one vice. Although with Daniel she was less inclined to steal because he was very rich. She only continued stealing in order to keep in practice and to get the thrill of the possibility of getting caught, the adrenalin rush.

Daniel kept groping under the bathrobe. Jackie turned on the stool and threw her arms around his neck, lips meeting lips. She wrapped her legs around his waist and he picked her up and carried her back to the bedroom. “He’s waking up again.” she whispered in Daniel’s ear.

An hour and a half later, Daniel was asleep. Jackie put on her robe and went to the bathroom. She showered and washed her hair, then returned to the bedroom to put on some clothes. She was living with Daniel and had her own closet in the bedroom. She picked out a light blue blouse and gray slacks. She went back to the bathroom to dry her hair and put on some makeup. By the time she came out of the bathroom she looked like a movie star. She went to the kitchen, got the keys to one of Daniel’s cars, grabbed the newspaper and went to the garage. She backed the Lexus out of the garage, then called Steve on her cell phone while sitting in the driveway.

“Steve, this is Jackie. Have you seen the newspapers?”

“About Mayor Ford or the Raptors?” Steve asked.

“Page 4, front section.”

“Okay, page 4, what am I looking for?”

“See the story about the Ferrari? Read it.” she demanded. She could hear Steve mumbling his way through the story, and she waited.

“Is that your Ferrari? What about the safe?” Steve asked, suddenly aware of what the story means.

“It doesn’t say anything about the safe. We don’t know if they have it or not. We have to go and get it.” Jackie said.

“Now? I’ve got a job these days Jackie. I have to go to work this afternoon, and tomorrow. Can’t it hold until Saturday? If the police haven’t found it, then it should still be there on Saturday, and if they have found it, then it doesn’t matter if we look today or Saturday.”

“Okay, Saturday. Pick me up around 10 am. Bring some flashlights, we may have to go in after dark. Wear dark clothes.” They hung up. Jackie was impatient and wanted to know now if the safe was still there. It was hidden, but not that well hidden.

She backed the car out onto the street and drove off. She wanted to think. It was possible the police had found the safe, and were now waiting in the barn, maybe with security cameras, for someone to come and reclaim it. Then they would capture them. She found herself driving all the way to Mount Dunham and out to the abandoned farmhouse. As she came upon the farm, she noticed that there was nowhere to hide. There were few trees and the land was flat. There were clear views to the houses on either side. There were no vehicles at the burnt out house, but she could make out the police tape and danger signs around the house. It would be difficult to go in there without being noticed. Someone would surely see a car or flashlights. What else was there to do?

Jackie continued driving east. At the next house she noticed a woman getting out of a car. She drove to the next crossroad and pulled off highway 12 and parked on the side of the gravel road. She turned around and pulled into the laneway where it said Palmer on the mailbox. She got out of her car and went to the door of the house and knocked. Mrs Palmer came to the door and opened and spoke through the screen 84 CHAPTER 7. OCTOBER 11

door, “Yes, how may I help you?” she said. “Hi, I’m trying to find the Gunther’s place, Sara and John? I think

I must be lost.” Jackie said.

“I’m sorry, I don’t know that name. Do you have an address?” Mrs Palmer asked.

“No I don’t. I don’t have a phone number either. They have their farm for sale, and I thought I would stop and have a look at it.” Jackie said.

“If you’re looking for properties you could look at the one next door. It’s going to be auctioned soon.”

“You mean the one where the house has burned down, over there?” “Yes, that one.” “Do you know how much land comes with it?” Jackie asked. “Not for sure. You could go into Mount Dunham and ask at the

bank.” Mrs Palmer said, then continued, “Did you know that they found a Ferrari in the garage after the fire?”

“You’re kidding. Who would burn up a Ferrari? Did you see the people that started the fire?”

“To me it looked like one person, in a pickup truck.”

“Did they find anything else in the fire, or in those barns?” Jackie asked. She couldn’t put her finger on it, but Jackie noticed that Mrs Palmer became momentarily quiet with that last question and was looking at her carefully.

“No, I don’t think they found anything else. I don’t think they really looked. Now if you don’t mind I have to put away my groceries.” Mrs Palmer said.

“Yes, of course, please don’t let me hold you up.” Jackie turned and went back to her car, and sat inside.

“Well Mrs Palmer, you became sensitive when I asked about the barns, and you were looking at me kind of oddly, like you recognized me. You must know about the safe.” Jackie thought to herself, “You definitely know something.”

Jackie started the car and drove back towards Mount Dunham, but decided to stop at the Lawrence farm and see if she could gather any other information, but Grace Lawrence proved to know very little about the fire or the barns. However, Jackie did meet Mickey, and made a mental note to bring some meat next time. She returned to Owen Sound in time to go to supper with Daniel.

86 CHAPTER 7. OCTOBER 11

Nate Haslett

October 11, 2012, Police Station, Mount Dunham

Linda made an early morning call to start her day. “Good morning, sir. May I speak to Nate Haslett?” “Speaking,” he said. “This is Linda Logel, I’m a fire scene investigator in Mount Dun-

ham.” Linda imagined his moustache and goatee, and the faint aroma of beer, and seemed to hold her breath as though she were speaking to him in person.

“You’re not asking for money are ya, cuz I can’t help you out.” I guess he’s not one for charity thought Linda.

“No sir, I’d like to ask you a couple questions concerning a recent fire. Do you have time now to answer them over the phone?”

“What fire?” Nate was immediately on his guard. “Sir, a few weeks ago you dropped by the police station in Mount

Dunham and asked about some abandoned property east of town.”

“I might have. I don’t recall exactly. Do you know the date I was there?”

“According to our records, it was September 17. We told you the farm was abandoned and to see the bank if you were interested in purchasing it.”

“Oh yeah, I remember.”

“Sir, did you go to the bank to see someone about the property?” Linda asked.

“Yes, I went to the bank, but no, I didn’t see anyone about it. They had a sign on the bulletin board inside that gave me all the information about the auction.”

“Sir, do you know that the farmhouse on that property burned to the ground the other night?”

“No, really? Guess I can scratched that place off my list.” Nate said.

“What list is that sir?”

“My list of potential new homes. I’m getting married and need some land with a house on it.” Nate said.

“Where were you on the evening of October 9 this week? That is, Tuesday night.”

“I don’t remember, let me see.” He was thinking. “I think I was at the country music bar in Graydon. I usually go there. My girlfriend works there.”

“What do you do for a living, sir?” Linda wondered.

“I do custom service farming. I mainly help farmers during planting and harvesting seasons. I travel all over Ontario. Wherever there is work.”

“Yes, I understand. Would you please stop by the police station in Mount Dunham the next time you are traveling this way. I’d like to get your fingerprints and maybe a DNA sample, if you’re willing to cooperate, in order to eliminate you from our inquiries.”

“Do I have to? I’d rather not.”

“That’s fine Mr Haslett. If we absolutely need them, we’ll get a warrant for them. I had hoped you would prefer to cooperate freely.” Linda said cheerily, but was disappointed inside.

“Well, thank you sir for your time. I hope you find the right property.” Linda hung up. She sat in her chair mulling over the conversation. Nate Haslett gave nothing away, and seemed to be a cool customer, if he was the arsonist. Nate’s work, however, allowed him to travel through Ontario, and so he would be familiar with abandoned farmhouses. She had a strong feeling that Nate was involved in this fire. Linda could not imagine Nate having a girlfriend, and wondered what she would look like. She wrote the new information into her notebook on the case.

Linda started searching the computer for Nate Haslett. There was one arrest for DUI over 5 years ago, but no fingerprints or DNA in the system. Otherwise his record was clean. Linda thought that if Nate Haslett was an arsonist, then maybe he could be involved in other fires. Linda began a search of fires that occurred over the last five years in southwestern Ontario. There were almost a thousand.

“That’s too many to go through.” she thought to herself. She filtered out those that were urban properties or accidents, leaving less than fifty rural fires.

“Still too many.” she whispered. Next she screened out those that occurred during daylight hours. That left nine. She looked for those that were abandoned houses. Five left.

“Now we’re talking.” She started going through them one at a time. She wrote down the names of the last owners, the locations, and the name of the person who phoned in about the fire.

     Where         Owner           Phone in
     ------------------------------------------
     Belmont       S. Jackson      Neal Howard
     Newmarket     D. Wiggans      Norm Haskel
     Paris         A. Fletcher     Norm Howard
     Goderich      W. Grayson      Bob Johnson
     Mt Dunham     M. Blanchard    Betty Palmer

Linda found it odd that the initials of the first three phone ins were all N and H. Was this a coincidence that they were the same initials as in Nate Haslett? She could verify this by trying to phone each of the first three. She went back through the first three files and got the phone numbers of the phone ins. The numbers were all different. The area codes agreed with the locations of the fires. It was beginning to look more like a coincidence. She called the first on her list, Neal Howard.

“Neal Howard speaking.” came the response on the call.

Linda identified herself and the purpose of her call. Neal Howard confirmed that he phoned in the fire. Linda thanked him, and disconnected the call. There was no apparent link of this fire to Nate Haslett.

Linda punched in the next number on the list for Norm Haskel.

“Hello, Norm’s Garage, Bill speaking.” came the answer.

“Is Norm Haskel there please?” Linda asked.

“Sorry ma’am, there’s no Norm that works here. It’s just called Norm’s Garage. Besides, the previous owner was Norm Staley, not Norm Hassle, or whoever you asked for.”

“Thank you Bill.” Linda disconnected, she did not explain about the fire. This could have been a bogus name, there was no Norm Haskel at this phone number. Nate could have called from this garage and used the ’Norm’ part to make up a phony name. This fire could possibly be linked to Nate Haslett.

She called Norm Howard in Paris.

“Hello?”

“I’m calling for Norm Howard please?” Linda said.

“There’s no one here by that name, you must have the wrong number.” and the person on the other end hung up. Linda thought that this could be another possible Nate Haslett call. She did not call the last two names because they did not have the correct initials, and because she had already spoken with Mrs Palmer.

All in all, there was nothing substantial to link Nate Haslett to the other four fires of abandoned properties, except the initials N and H on three of the four fires. This approach was going nowhere.

Linda went back to the files and looked at the evidence that was logged in for each scene, and witness statements. There was nothing that stood out as common among the fires, except that they were abandoned farmhouses in isolated locales and the fires were set at night. No vehicles or people were seen at any of the fires except the last. Then looking at the fourth fire on her list, one of the witnesses was Nate Haslett! The fire was in Goderich, over 100 km from where Nate Haslett lives. Maybe he was working in that area, as he said. However, Nate was now directly connected to two of the five fires.

“I have to get Nate Haslett’s finger prints and maybe a DNA sample too.” Linda told herself. “I better get a warrant for those.”

Just then her name was called on the loudspeaker system. She was wanted at the front desk. Tracy saw her and nodded towards a tall shaved headed man. Linda went up to introduce herself.

“Hello, I’m Linda Logel, fire scene investigator.”

“Good. I saw the newspapers this morning and the story about the fire you had up here with the Ferrari in the garage. I believe that is my car. My name is Paul Selinger.”

“I was going to phone you today, to ask some questions. How good of you to come here instead. Please follow me this way.” Linda led Mr Selinger to the interrogation room. “Please be seated and wait here, I’ll just be a second.” She left him there and went to her office to get the file and her notes, then she returned. “Would you like any refreshments before we begin, Mr Selinger?” she asked.

“No thank you. Please, just call me Paul.”

Linda sat down and took out her pen and notebook.

“Okay Paul. Could I see your driver’s licence please?”

Linda took it and made notes about the address and birthdate. Then she handed it back to him.

92 CHAPTER 8. NATE HASLETT

“Thank you. How long have you lived in Toronto?” Linda asked. “Almost my whole life. I spent two years in Europe back in 2002. I’m

a chef and I was apprenticing in France and Italy. Now I’m head chef at the Terrozi Bread House in Toronto.” He handed Linda a business card, which she took and stapled to the file.

“Do you have any idea how your car ended up in a garage of an abandoned farmhouse 200 km from Toronto?” she asked.

“None whatsoever. I have never been here in my life. You must know I reported the car stolen on July 4, 2011.”

“Do you know who stole your car?” “I think so. My girlfriend, at the time, was Jackie Shambeau, and I

believe she disappeared with it.”

“Have you heard from your girlfriend since that day?” “No, I’ve been looking for her ever since. I love her, you understand.

I think she may have been forced to leave Toronto, by this other man.” “Do you know who this other man is?” Linda asked. “I think his name is Steve. I have nothing else.” “Do you have a picture of Jackie?” “Yes, I brought a couple.” Paul took them from his coat pocket and

gave them to Linda.

“Wow, she is a beautiful woman.” Linda acknowledged. “She looks like a movie star. Could I keep one of these? For the file?”

“You can have them both, I have more of them. I’ve nearly gone mad not knowing where she is or what has happened to her.” Paul said.

“Paul, you know your car was completely destroyed in the fire, right?”

“Yes, I know, but I was hoping to find clues about Jackie. She must have been here.”

“Well, I didn’t know anything about Jackie, until now. I’m sorry I can’t help you today. But I will definitely be looking for her as I go forward in this investigation, and this Steve character too. As you know there were no bodies found in the fire.”

Linda suddenly thought about the sighting of the car at the Black Forest Restaurant, which she had not yet pursued.

“Paul, a Ferrari was spotted at our Black Forest Restaurant back in July 2011. Bring your pictures of Jackie. I need to ask people there about that sighting. Would you like to help? Do you happen to have any pictures of Steve?”

“That would be great. Sorry, no, I do not have any pictures of Steve.” Paul said.

“I’ll get my coat and we’ll go.” Linda took her file and left the interrogation room. She got her coat and hat, then collected Paul as she walked up to the front.

“We’re going to the Black Forest to ask a few questions. I’ll be back in 40 minutes.” Linda told Tracy. “Oh yes, I need a warrant to take fingerprints and a DNA sample from Nate Haslett, could you get that for me please?”

“No problem.” Tracy said.

“Thanks Trace.” Linda and Paul were out the door and headed south down Main Street to the Black Forest. As she walked along next to Paul, she realized how tall he really was. He seemed to be physically fit and wondered if he was a runner. His bald head was a little lumpy at the back, but he carried himself in a dignified manner, with some confidence.

* * * * *

A sign on the front door of the restaurant said “CLOSED”. The restaurant was only open evenings. Linda could see lights inside and so she knocked loudly on the door. After much knocking and banging a woman finally came to the door drying her hands in a towel and pointing to the ‘CLOSED’ sign. Linda waved back showing her official ID. The woman nodded yes, and motioned for them to go around to the back.

The woman met Linda and Paul and held the door open for them at the rear of the restaurant. Linda introduced herself and Paul, and the woman said her name was Mary Stewart. She was somewhat unkempt apparently from washing dishes and cleaning the dining area from the previous night’s business. There were other people helping to clean the kitchen area.

“Paul Selinger, that name sounds familiar.” said Mary.

“I’m the head chef at the Terrozi Bread House in Toronto, maybe you’ve been there.” Paul said.

“Indeed, I have, that must be it.” Mary said happily.

“I’m investigating the fire at the Blanchard farmhouse.” Linda broke in.

“I’ve heard about it.” said Mary.

“Well, I don’t know if you heard there was a Ferrari in the garage that was completely burned out.”

“That rumour has reached me too. So it’s true?”

“The car belonged to Paul here, which he reported stolen in July 2011.”

“And?” Mary said.

“The tow truck driver told me that he saw a grey Ferrari parked outside this restaurant also in July 2011. I was wondering if you or any of your employees might have seen it, and the people who were in it.” Linda asked. “Paul has a picture of the woman, that may help you.”

“You’ve got to be kidding. I might have seen the car, but I sure wouldn’t remember who was in it. My memory isn’t that good.” Mary wailed. She took a glance at the picture and shook her head negatively.

“She’s a pretty one, isn’t she?” Mary said.

“What about any of your employees at the time? Maybe the waiters and waitresses?” Linda asked.

“Waiters and waitresses turn over pretty quickly in the restaurant trade.”

“Can you get me a list of the names of all your employees during July 2011, please?”

“Wait here, I’ll check the records on my computer in the office.”

Linda and Paul took a seat at the bar, in the dark, and looked around the dining area. The front of the restaurant had glass windows all across the front so that patrons, waitresses, and waiters could have seen all of the cars parked in front of the restaurant. There was some parking at the left side of the building that was not visible from the dining area. So it would depend where the Ferrari had been parked. Linda started thinking out loud again,

“Given that the car was so expensive, the driver would likely have parked in the front so that they could keep an eye on it while they ate. So the chances of finding someone who saw the occupants might be fairly high.”

Paul nodded in agreement. Mary returned and handed a piece of paper containing the names of 12 employees during July 2011 to Linda.

“Mind you, not all of them would be working at the same time, nor everyday. All of them, except this one, Sarah Jelenik, no longer work here. Sarah has stuck with us. She’s not here right now, but she’s scheduled to work tonight. Our customers like her because she’s friendly and rarely makes a mistake with the orders. She’s not bad looking either, but she is married.” said Mary. “Is there anything else I can help you with?”

“No, this list is great. Thank you very much.” Linda and Paul departed via the rear door from which they entered. She folded the paper and stuck it inside a pocket of her pants.

As they were leaving Mary said to Paul,

“If you are ever looking for a place to work, please give me a shout.”

Paul laughed and said,

“I’ll keep that in mind.”

Outside the restaurant, Linda said to Paul,

“It’s going to take some time to go through this list.”

“Do you want me to help?” Paul said.

“No, its best to keep the investigation on an official basis. But I’ll let you know if I learn anything about Jackie’s whereabouts.”

“I’ve got to do something while I’m here.” Paul looked frustrated.

“I can’t stop you from looking around and asking people questions, but I wish that you just go back to Toronto. I’ll do my best to find her.”

“Consider this, if Jackie and Steve left the car at that farmhouse, then they must have found another means of transportation. They either bought another vehicle from around here, or they stole a vehicle.” Paul said.

“Or they could have brought another vehicle with them? Maybe there was another person involved.” Linda said. “Anyway, the nearest car dealers are in Orangeville, or south to Arthur. I can check the police files for stolen vehicles around that time in this area. You can check the dealerships, if you wish. Just let me know if you discover anything. A witness told me that there was a white 4-door sedan at the Blanchard house in 2011. The witness could not tell the make of the car.”

Paul’s face brightened up, now he had something he could do.

“I’ll be back later this afternoon and let you know how I make out.” He hurried away towards his car, and Linda continued walking back to the station. Linda thought,

“Paul’s correct. They would have needed alternative wheels to leave Mount Dunham.” Linda stopped and reminded herself that finding the person who started the fire was her priority, and not the persons who left a very valuable car in the garage. She doubted if Jackie or Steve were connected to the fire.

* * * * *

“Where’s your friend, Linda?” Tracy asked as Linda came into the station.

“He was interested in finding the woman that stole his Ferrari, and there was a chance that someone at the Black Forest might have seen her. All I got was a list of people that might have seen her. Now I have to check them out.” Linda said.

“Sooo, where is he?” Tracy asked again.

Linda rolled her eyes finally getting Tracy’s meaning, “He’s gone to check out car dealers in Orangeville and Arthur. He figured that they probably bought or stole another vehicle to get around in.”

“So, he’ll be back sometime?” Tracy asked.

“Maybe. I told him to let me know if he found anything. Say did you get that warrant for me?” Linda asked

“The request has gone in.”

Linda escaped to her office. She took out the list of employees from the Black Forest restaurant. She got out the phone book, and tried to match names to phone numbers and addresses. She found 6 of the 12 names on the list. She began calling the six, one by one, saving Sarah Jelenik for last. Three of those she phoned did not work on the day in question, but they heard about the Ferrari the next work day. None could say anything about the occupants of the car.

Bill Hayes was a waiter that day, and he remembered the Ferrari. There were two occupants in the car, a man and a woman. The woman was very beautiful, he recalled. The man was medium height, barely taller than the woman. The woman’s eyes were light gray and she had long black hair. Everyone in the restaurant gazed at her as she was led to their table. He did not remember anything else about them. A number of customers walked around and admired the Ferrari when they left the restaurant. The man became agitated at the amount of attention the car was attracting. They ate and left very quickly.”

“You’re sure that her hair was black?” Linda said.

“Definitely.” Bill replied. Linda thanked him for his help and said she may contact him again to show him a picture.

Lastly, Linda called Sarah Jelenik, the current employee at the Black Forest Restaurant.

“Yes, Mary told me that you would likely call about the Ferrari. I saw the car. There was a man and a woman. The woman was very neat with long black hair, but it wasn’t really her hair. A woman can tell these things. When I served them, up close I could see a few other light coloured hairs mixed in with the black. The man had bad breath. After I got my first whiff of it, I kept a good distance from him. I don’t recall what they ordered, but once the food was served the man was very much in a hurry. I’m not sure what that was about. I would say that the two people were not lovers. Even though the man tried to give that impression, the woman gave a distinct air of indifference towards him. The woman was polite and thanked me for everything. The tip they left was smaller than I expected, and smaller than average, given that they were driving a Ferrari.”

“Did you hear them talk about where they were going or where they were from?” Linda asked.

“No, they might have said something, but it didn’t stick in my head. When they left they turned right heading north through Mount Dunham.” Sarah added.

Linda thanked Sarah, hung up, and made more notes for the file.

* * * * *

The warrant to get fingerprints and DNA from Nate Haslett did not come in until 4:45 pm. It was too late to try to find Nate today, it would have to wait until the next day. Linda started putting away her notes and files on the fire. She had checked for stolen vehicles around July of 2011 in the Mount Dunham area and found none that were white, 4-door sedans.

Just then her name came across the intercom calling her to the front desk.

“Paul, how did your search go?” Linda asked when she saw who was waiting for her. “Come on back to the interrogation room.”

“Success, in Arthur. The salesman remembered Jackie, although he claimed her hair was black. He said there was a man with her, and that they came in a Ferrari. The man bought the sedan and paid cash. It was a Chevy. He wouldn’t give me any other information because I wasn’t the police.”

“That’s okay, I can phone him and ask for it. Well, great. I guess that doesn’t help us to know where they are now, but it confirms that they were here. Do you think Jackie or Steve would have deliberately set the house on fire to burn the Ferrari?”

“No, Jackie was very attached to the Ferrari, she wouldn’t have destroyed it. And I don’t think she would have let Steve do it either. Neither of them were destructive people.” Paul said.

“Are you going back to Toronto now?” Linda asked.

“No way. I’m closer to finding Jackie now, than I’ve ever been. I’d like to stay and help.” Paul said.

“My priority is to find out who started the fire, not to find missing people. Tomorrow I’m going to try to find a suspect and serve a warrant. I don’t see anything you can do.”

“If you find out the name on the registration of the white sedan, then you might find his name and address among the driver’s license. Then I could go to wherever it is and try to find him.”

They went to Linda’s office and she powered up her computer. Paul gave her the phone number of the dealer who sold the car to Steve. Linda phoned it. Linda explained the situation and gave him her badge ID. After some minutes the dealer came up with the licence plate numbers that were put on that car, and the name of the purchaser. Linda wrote down both of these, and thanked the dealer for his assistance.

Linda made a copy of the name and licence plates on another sheet 100 CHAPTER 8. NATE HASLETT

of paper and gave it to Paul. Then she keyed in the name into the driver licence database and hit search. The name and picture of Steve Swider came up. Linda showed it to Paul and asked,

“Is this the Steve that you knew?”

“Yes, that’s him. Swider is his last name? What’s his address?” “671 Olive Street in Owen Sound.” Linda had two copies of the page printed. “Do you think Jackie will be in Owen Sound too?” Linda asked. “I have to check it out. Maybe Steve knows where she went after

they were here. They only had one car, so chances are they both went to Owen Sound. I don’t really know.” Paul said. “Thank you very much for your help.”

“What will happen when you find Jackie?” Linda asked. “I don’t know that either, but I want to find out.” Paul said. Linda walked Paul back to the front desk and let him out, and wished

him good luck as he left the station.

“Do you believe it, he actually found what he wanted to know?” Linda said to Tracy.

“I’m going home. See you tomorrow.” Tracy said.

“I’ll be going out first thing to find Nate Haslett.” Linda said, “So I won’t be in until later. Justin’s driving me.”

Linda stopped by Captain Muller’s office and updated him on the fire investigation. Then she made arrangements with Officer Justin for the next morning’s excursion to find Nate Haslett. Then she went home to her cat.

Looking for Nate

Friday, October 12, 2012

During the night Linda was awakened by the howling of the wind and the sound of rain hitting her bedroom window. She briefly opened one eye and saw drops of water flowing down the window in random rivelets. Her eye closed and she was asleep again.

Rain was still falling when she awoke and got out of bed. She decided to take her morning run wearing her water resistant sweat clothes with the hood to keep her head dry. Due to the cold and wet, she only ran three kilometers. Despite the rain gear, she was still nearly soaked to the bone. She took a long shower to warm up afterwards.

Officer Justin Jankowisz showed up at 8:00 am. He had found directions to Nate Haslett’s residence, and had punched them into the GPS system in the police car.

“It will likely take us 40 minutes to get there.” Justin said.

“Sure is wet and cold this morning, isn’t it?” Linda said.

“You have to expect it at this time of year.”

“You don’t have to like it though.”

“It’s better if you just accept it. Nothing you can do about it.”

Although Justin carried a pistol every day on the job, he had never had reason to draw it out or to use it. The only time he used it was at a practice range for his annual qualification checkup. Linda learned that Justin was a martial arts master, and that these talents had been useful while on duty. Linda thought that maybe those skills would be tested today with Nate Haslett.

Linda rode along, quietly, looking out the window at the passing country side. With the wet and cloudy day, the farmsteads looked more isolated and lonely than usual. Puddles were collecting in some fields. More rain had fallen overnight than she thought, as she saw little newly created streams of running water along the road.

“You think this Nate Haslett started your fire?” Justin asked.

“At the moment, he’s the best suspect I have.” Linda said. “I checked out other fires of abandoned farm houses, of which there were five, and Nate was a witness to one of them. I think he might be linked to another as well, but there is no direct evidence, just his initials, “N” and “H”.

“So these fingerprints should confirm it or eliminate him.” Justin said.

“If he’s not the one, then I don’t have any other suspects, unless the lab tests identify someone else. Lab report should be on my desk when we return.”

“You know, many of these types of cases are never solved.”

“I want to solve this one. It’s my first case.” Linda said.

They talked about Paul Selinger, the Ferrari, and Paul’s search for Jackie. Linda showed Justin the picture of Jackie from the case file Linda had on her lap.

“You weren’t kidding, she is really good looking.” Justin said. “I guess I’d go looking for her too, if I knew her. Don’t tell my wife I said that.”

“I wonder how much effort I should put into trying to find Jackie myself.” Linda said. “I don’t consider her a suspect in starting the fire.”

“You should probably find either this Steve or Jackie to figure out why they left the car in that garage.” Justin offered. “The presence of the Ferrari needs explanation. It seems like they were hiding it, so they may be wanted by the law for something.”

The GPS lady announced that they should turn right in 1 kilometer. Their destination was 2 kilometers on the left side of the road after the turn. Linda pulled out the warrant and the kit for taking fingerprints and DNA samples in preparation. As they pulled up to the house Linda couldn’t see a pickup truck, only a silver Pontiac Vibe parked in front of a two car garage. Inside the garage were stacks of boxes of empty beer bottles, waiting to be returned to the Beer Store.

Linda was glad the rain had let up to a light mist. Before they went to the house, Justin entered the licence plate number of the Pontiac Vibe into his computer. The owner was Irene Haslett.

“Just checking.” he said. “No unpaid tickets or anything.”

Both her and Justin went up to the entrance of the house. They could hear a big dog barking madly in the house, warning everyone inside of their presence. They looked at the ‘Beware of Dog’ sign before they pressed the doorbell. Justin placed his hand on his weapon in preparation. A skinny woman in her late 50’s, wearing dark slacks, pink blouse, and a dark blue sweater opened the door and squeezed through, closing the door on the agitated, barking German Shepherd inside. A cigarette hung from her lips. Everyone was standing on the stoop, in the mist. She took the cigarette between two fingers and folded her arms as she blew smoke away.

“Good morning, ma’am. I’m Linda Logel and this is Officer Justin Jankowisz of the Mount Dunham Police. We’re here to serve and carry out a warrant on Nate Haslett. Is he home, please?”

“You’re too late. He left early this morning. Won’t be back until this evening, I expect.” she said quietly.

“Do you know where he went?” Linda asked.

“He never tells me where he’s going.” the woman said.

104 CHAPTER 9. LOOKING FOR NATE

“Who are you and what is your relationship to Nate?” Linda asked. “I’m his mother, Irene Haslett.” “Does his father live here too?” Linda asked. “No, Neal died about seven years ago in the feed mill fire in Hills-

burgh.” Linda could see that it saddened the old woman to recall this event.

“I’m sorry ma’am.” Justin said. “Nate told me he was getting married soon. Can you tell us his

fiancee’s name and where she lives?” Linda said.

The woman gave a little chuckle and a twinkle came to her eyes. “He’s been giving out that old story again? He doesn’t have any

fiancee.”

“Are you sure?” “He doesn’t have enough money to get married. He can barely keep

his truck gassed up and buy his own beer.”

“He told us that he does custom farmwork for other people.” “That’s true, but only if he has to. To get money, you see. And he

usually has to get money.”

“Does Nate have a cellphone number?” “Yes, but I don’t keep track of it.” “Could you phone us when he returns, and tell him that we would

like to serve him this warrant?”

“A warrant? What’s this all about?” Irene asked. “It’s in relation to a house fire I’m investigating. I’d like to get

his fingerprints and a DNA sample. Do you know where Nate was on Tuesday night this week?” Linda asked.

“He got home very late, around 2:00 am.”

“How do you know the time?” Linda asked.

“The dog barks whenever anyone comes to the house. Brutus woke me up and I looked at the clock on my night stand.” Irene said.

“Thank you Mrs Haslett. You have been very helpful. Remember to call us or have Nate call us when he gets home today.”

Linda and Justin returned to the police car. As Linda was getting into the car, she thought she noticed a movement in a curtain on the second story of the house. The dog was still barking at the door. She stared closely at the window and thought she could make out the figure of someone through the curtains.

“Mrs Haslett!” she cried. “Who else lives in this house?” Linda pointed to the upstairs window to the right.

“That’s my daughter, Nancy. It’s just her and Nate, and Brutus living with me.”

“Could I speak to her please?”

“She’s probably not dressed yet.”

“Please ask her to come out and talk to us Mrs Haslett.” Linda pleaded.

Irene squeezed back through the door and shut it. Linda and Justin waited by the door. They could hear Irene yell upstairs to her daughter. Two minutes passed before the door edged open again. This time a sleepy looking young woman about 25 yrs old slipped through the door wearing a white robe tied around her waist, and her dark brown hair going in all directions, not brushed or combed. She was a younger version of Mrs Haslett, fairly thin.

“Sorry, to get you out of bed.” Linda said, and she gave her usual introductions.

“Do you know where Nate was going today?” Linda asked.

“Why would I know? I hardly say anything to him, and he pretty well ignores me.”

“Do you know if he has a girlfriend?” Linda asked.

“If he does, I’ve never met her.” Nancy replied.

“Do you know when Nate got home on Tuesday night?”

“I couldn’t tell you. I sleep with these special ear muffs that block out most noise. Otherwise the damn dog wakes you up whenever it decides to bark.” Brutus was still making it known that he did not appreciate having visitors today.

“Okay Nancy, thank you for helping us.” Linda said, and turned to leave. Nancy went back into the house.

Justin and Linda left the Haslett place and got back on the road.

“Looks like Nate is becoming a stronger candidate.” said Justin.

“Yes, he’s been lying. He doesn’t have a fiancee and he was out during the time of the fire, so he could have been there. Nate gave me some details about where his girlfriend worked.” Linda was going through the file to find that note. “Here it is. Graydon, a music bar. Can we go there now and check it out?” Linda keyed it into the GPS and Justin had to turn the car around.

Graydon was a one-stop-sign village in the middle of nowhere. On one corner was a snowmobile repair shop. On another was a corner convenience store that looked closed permanently. Cross corner from that was a hotel with bar, and a sign outside advertising the entertainment schedule for that week. The remaining corner had a white sideboard house, that was clean with well manicured gardens around the house. They parked in front of the house and crossed the street to the bar. There were no customers inside, but the smells of liquor and cigarettes were strong. A slightly balding man with black hair slicked back over his head, sat behind the counter watching a TV program. When he got up he had on a multi-coloured shirt, unbuttoned down the front, with a Tshirt underneath both hanging out over black jeans. He was clean shaven and smelled heavily of cologne. His very muscular arms had tattoos on both forearms. He seemed to be the epitome of a biker.

“Good morning sir.” said Officer Justin.

“Good morning officers. How may I help you?” he said putting both hands on the bar, displaying his tattoos prominently.

Linda pulled out a picture of Nate Haslett, and asked, “Do you recognize this man?” “That’s Nate without his moustache and goatee.” he said. “Correct. Was he here on Tuesday night?” Linda asked next. “I don’t recall seeing him this week.” “What time do you close up?” “Midnight. Business is usually all gone home by then.” he said. “Do you employ any female waitresses sir?” “Yeah, she roams the tables and checks that everyone is drinking or

eating something.”

“Does this employee have any relationship with Nate?” Linda asked. “You’re joking right? My waitress is 44 years old and married. To

me.” he laughed.

“Nate told us his girlfriend worked here.” Linda told him. “Not true, I’m sorry.” and he held his hands out palms up. “What

can I say?”

“Have you ever seen him with a girl in here?” Justin asked. “Nate is usually by himself. His sister came in with him once, a long

time ago. I think it was her birthday, or his birthday.”

Linda asked the owner for his name and phone number, then thanked him for his help and left the bar. The man returned to his chair and turned the volume back up on the television.

“Well, that corroborates the mother’s story that her son does not have a fiancee.” said Justin.

“And that Nate lied about his whereabouts on Tuesday night. Yep, he’s becoming a very strong suspect.” Linda said.

They drove back to Mount Dunham in silence, while Linda made notes of their morning visits in her file. She wrote

• No girlfriend or fiance.

• Not home or at bar during the fire.

• Got home after 2:00 am.

• No money.

• No intention of buying Blanchard property.

“That dog must be a pain, barking every time someone comes to the house.” Justin said, breaking the silence.

“It kept us from going in the house too.” Linda said. “For all we know Nate could have been in the house, except there was no pickup truck there. If I come here again, I’ll bring a warrant to search the house, and Nate’s bedroom.”

“Be prepared to deal with that dog.” Justin warned.

“You bet.”

Owen Sound

Owen Sound, Friday

Paul Selinger pulled into the Owen Sound Tourist Information Centre on the outskirts of the city about 11 am on Friday. He went to the telephone booth and started looking through the phone book for Steve Swider or Jackie Shambeau. There were no listings for anyone named Shambeau, but there was one for Swider. He phoned the number, but there was no response. He wrote down the address by the name, then looked for a complimentary map of the city. There were none obvious, so he asked the old gentleman behind the desk. The attendant pointed to the location of the maps and Paul thanked him. It took a while to find the street on the map. It was a long street that was separated by a river that snaked its way through the city. Olive Street was on both sides of the river, but did not go over the river. Paul asked the attendant if he knew which side of the river would have 671 Olive Street. The attendant went to his computer and put in the address in Google maps, and found the house on the west side. Paul thanked him and left the centre.

Paul studied the map in his car and oriented himself. He memorized the street names where he would have to make turns. He regretted not getting a GPS for his car. He started out on his trip. When he came to his first turn, it was a one-way street going in the wrong direction, naturally. So he continued to the next street and turned right on it. He then parked the car and studied the map again. Under his breath he was cursing one-way streets. He noticed that Olive Street was also one-way, so that he would have to go one more street west then make two rights to come back to the address he wanted. The rest of his journey went as planned and within ten minutes he was parked in front of 671 Olive Street. The houses on this street were somewhat small and rundown, and it was obvious that most of the houses were likely rentals. The house at 671 Olive was two-storey with green siding, and a white porch on the front that was slanted at an angle downward from the door of the house. The steps up to the porch were not level. There was a large green, plastic garbage can to the right of the door. There seemed to be several old newspapers in blue plastic bags under the porch. A cat was curled up on the porch to the left of the door, watching Paul’s vehicle. The driveway on the left side of the house was empty and the garage at back was closed. Most of the neighborhood seemed to not be home. Paul went to the door and knocked, but there was no answer, as he expected. He would have to wait until Steve Swider came home, and he had no idea when that would be.

It was 11:30 am and Paul was hungry. He had noticed a convenience store a few blocks away before he had turned. He locked his car, then walked to the store. The store did not have very much in it, but he bought some wheat buns, a package of cold cuts, a package of cheese slices and a tomato, plus a bottle of ginger ale and a copy of the Toronto Globe and Mail newspaper. Back in his car he got out a spare set of cutlery from his glove compartment and made a sandwich. Being a chef he always had spare cutlery in his car, as well as some condiments. He detested using plastic cutlery. He devoured a couple of sandwiches and quenched his thirst with the ginger ale. He then settled in for a long wait and started reading the newspaper.

After an hour, Paul sensed that he needed a restroom. He recalled seeing a gas station across the street from the convenience store. He locked up the car again, and walked to the gas station, and asked to use their restroom. Once relieved, he returned to his car to wait.

At 4:37 pm, a white Chevy sedan pulled into the driveway at 671 Olive Street. Paul got out and walked up to the driver as he got out of the car. Steve was a little surprised to see someone, but it was obvious to Paul that Steve recognized him.

“Steve, remember me, I’m Paul Selinger.”

“Right, I remember you. What are you doing here?” Steve looked around quickly for police, and saw none.

“I want to find Jackie.”

Steve stood there looking at Paul trying to figure out what he really wanted.

“She went to Vancouver. I haven’t seen her in ages.”

“Do you have an address or a phone number?” Paul asked.

“No man. Look she doesn’t really want to see you again. Don’t you understand man. You were used and tossed away. That’s the way she is with guys.”

Paul was hurt, being told something that he often contemplated but did not want to believe.

“How do you know?” Paul asked.

“She laughed about it as we drove out of Toronto with your Ferrari.” Steve answered. “The only thing she liked about you was your cooking.”

Another wound inflicted on his heart. Could she have been that uninterested in him?

“She uses men. She is addicted to men and sex. I was one of her victims too, once, before you. She’s probably gone through another two or three men by now.” Steve said. “You should have figured that out by now.”

Paul turned and went back to his car and got in. He was defeated. He realized he was an idiot for thinking Jackie would want him back. He started the engine and drove off. He drove without any plan. He ended up near the sound, the northern edge of the city. The houses here were very expensive looking. There was a little park that had a dock going out into the water. He parked the car and walked out on the dock. He sat on the dock and looked out at the water and boats. Sea gulls hovered back and forth over the boats. Tears were in his eyes. Even though it was cold, he sat there a long time looking at the picture of her from his pocket. Steve must be lying to him, he thought. But why had she not ever contacted him? Steve must be right. And so he wavered back and forth. After an hour he concluded that Jackie was gone forever from his life. He decided to go back to Mount Dunham, then return to Toronto.

As he was leaving the parking lot of the little park, waiting to turn onto the street, a brown-grey Lexus drove by right to left, and there in the driver’s seat was Jackie. Their eyes met and both recognized the other immediately. Paul’s face brightened up, and Jackie’s took on a panic appearance.

Jackie sped up and turned right at the first street. Paul had trouble pulling out of the parking lot due to heavy traffic, but he finally poked the front of the car out into the street to pursue Jackie. He saw where she had turned and he also turned, but he could no longer see the Lexus. She had vanished again. Paul was determined to find her now. He had to confront her and get her to admit she felt nothing towards him. Steve had lied about Jackie being in Vancouver. Unfortunately he did not get a licence plate number on the car Jackie was driving. He decided to go back to Steve’s house and try to force an address or phone number from Steve. Steve must know where Jackie is living in Owen Sound, or at least be in contact with her.

* * * * *

After making several more turns Jackie was back on her way to the apartment. There was no sign of Paul in her rear view mirror. She didn’t know how Paul had tracked her to Owen Sound, but if he could do it, then so could the police. Jackie parked the car in the garage at Daniel’s apartment. Her cellphone chimed before she got out of the car. “What now?” she thought.

“Hello, Jackie? Steve here. Your old friend, Paul Selinger, is in town. He stopped by my place an hour ago.”

“So you told him where I’m living?”

“No, I told him you were in Vancouver.” “Well, I just saw him two minutes ago, just five blocks from here. I

managed to lose him. But he knows I’m not in Vancouver now.”

“Jeez, I bet he’ll be coming back to see me soon. He won’t be happy that I lied to him.”

“You better hide out somewhere tonight. Don’t forget you’re picking me up in the morning. We have to go get the safe.”

“Don’t worry, I’m out of here. See you at 8.” Jackie got out of the car and went into the house. Daniel was in the

kitchen, eager to see Jackie. He put his arms around her, but she pushed him off after a quick kiss, she was not in a mood for sex.

“I’m going away for a couple of days. I have to pack.” “Where are you going?” Daniel asked, feeling a little rejected. “Business trip.” she said. “But you’re not employed.” “It’s better if you don’t know.” “Are you in trouble Jackie?” Daniel asked worriedly. “Not any more than usual. It’ll just be a short trip.” “We were supposed to go to a concert tomorrow night.” “You’ll have to go alone or take someone else. How about my sister?”

she answered.

“I could, she looks like you, but she is definitely not you. You know what I mean?” he said.

“You mean her views about sex?” Jackie said. “What views? She has no views. She’s opposed to it. She might as

well be a nun.”

“That’s not true, she’s just very selective. She doesn’t appreciate you like I do.” Jackie was yelling from the bedroom as she packed her suitcase. She had packed her largest suitcase, and as many of her favourite outfits as possible. She had a smaller case for her cosmetics and toiletries. She was planning for the possibility that she would not be returning to Owen Sound. In a smaller second suitcase she packed several of her disguises that she had used in past robberies. After she had packed she moved the suitcases to the front door, and Daniel noticed.

“Looks like you’re going for more than a couple days, if you ask me.” he said.

“I couldn’t decide what to wear.” she said. Jackie put her arms around Daniel’s neck, and he hugged her close to him.

“How about nothing?” he whispered. They smiled at each other, then walked off together to the bedroom. Jackie knew she had to satisfy him to avoid any more questions.

* * * * *

When Paul Selinger parked in front of Steve’s house, Steve’s car was not in the driveway, and Paul knew that Steve had gone into hiding. Thus, Steve and Jackie must have talked with each other. He got out of his car and walked up the crooked front steps onto the porch. The cat scampered off the porch quickly. Paul knocked on the door. He looked around, then tried the front door. It was locked. He went around the house to the back door. It was also locked, but neighbors could not see back there. He forced the door open by kicking it. The lock ripped out of the door frame, but he was in.

He went through the kitchen, looking at everything that was stuck to the refrigerator. There was no address or phone number for Jackie on the door. Paul looked on the bulletin board above the kitchen table. Still nothing about Jackie. There was a phone in the living room on an end table. There was an old rolodex for phone numbers. Paul flipped through them one by one. He didn’t find any for Jackie, but in the ‘S’ section one card had been pulled out, leaving part of it behind. This must have been Jackie’s card.

Just then the phone rang and scared Paul half to death. He sat there as it rang once, then twice. He couldn’t decide whether to answer it or not. So he didn’t. It wouldn’t be Jackie, and it wouldn’t be Steve because they have already talked to each other. After five rings the phone went silent and Paul was relieved.

Paul started looking through the house for waste baskets. There was one in the bathroom. Nothing. There was one under the sink in the kitchen. Then he saw it, the top part of the card from the rolodex. It had the initials J and S, and a number. Paul put the card in his pocket. He went out the back door and tried to keep it closed so that rain or animals couldn’t get in. He got in his car and drove to a motel and booked a room for the night. He tried the number he had taken from Steve’s house.

“Hello?” a man’s voice answered. Paul didn’t recognize the voice.

“Is Jackie there, please?” Paul asked.

“No, I think you have the wrong number.” and the man hung up.

Paul was perplexed and annoyed. Paul called the number again. This time there was no answer. That must have been Jackie’s current male friend, Paul thought, or a wrong number. Either way that number is not going to be useful any longer. He threw the card with the phone number into the waste basket.

“I’ll have to wait for Steve to return to his apartment. Then I can force him to take me to Jackie.” Paul thought to himself, and so he returned to Steve’s apartment.

116 CHAPTER 10. OWEN SOUND

The Money Envelope

October 13, Saturday

Ken Lawrence stared out he window of his bedroom. His parents were preparing to go into Mount Dunham to shop for groceries. He was dressed and ready to go over to the Blanchard place to look in the barn for his friend’s, Jeff’s envelope of cash. He guessed that it would take him 5 minutes to get over there, maybe another five minutes to get the envelope, and then 5 minutes back. A total of 15 minutes.

“Ken!” called his mother.

“Yes Mom.” he answered and went to the top of the stairs leading to the first floor.

“Could you clean the morning dishes for me while we’re gone, please?”

“Certainly Mom.” Ken wished they would leave soon before either parent thought of more things for him to do. Too late,

“Don’t forget about cutting Mrs Palmer’s lawn.” said his father.

Ken winced, he had forgotten about that. He’ll just continue on over there from the barn at Blanchard’s place. That meant he would have to do the dishes before he left. He quickly went downstairs to the kitchen and started in on the dishes, scraping off plates and rinsing them. Then he cleaned out the sink and piled in the dishes. He was starting to put in the water and soap when his parents and Roger said goodbye and left the house. He washed the dishes as quickly as he could. His quality of work was less than his usual standard, but he was in a hurry. When the last dish was in the strainer, he emptied the sink and dried his hands. The dishes were usually left to air dry. He went to his room upstairs and collected his cell phone, in case he had difficulty finding the envelope. Then he could call Jeff. Mickey followed Ken out of the house and Ken locked the door. Ken and the dog started off towards Blanchard’s place, going across the field.

Ken was at the back of the barn, out of sight from the passing cars on the road. When the road was clear he ran to the front and slid open the door. He heard a voice and looked up terrified. Mrs Palmer was waving from the back of her house. Ken looked up and waved back. Now he would have to make up a story about what he was doing, for Mrs Palmer. He only hoped that she wouldn’t tell his parents.

“I’ll be there in a couple minutes Mrs Palmer.” he yelled. Ken could hear an “Okay” come back in answer. Ken ducked into the barn and Mickey followed. Ken went straight to the corner of the barn and the trap door. On top of the door were bales of straw. Ken moved them aside, then lifted the heavy trap door. The envelope was supposed to be attached to the underside of the door. Ken looked, but there was no envelope anywhere.

“Aah, it probably came loose and fell inside.” he thought. Unfortunately Ken had not thought to bring a flashlight with him. He lowered himself into the hole and started feeling around with his hands. Finally after checking every corner and the entire floor, he was convinced there was no envelope. Mickey was looking into the hole too, but offered no help or explanations.

“Now what do I do?” Ken said to himself.

As Ken climbed from the hole Mickey tried to lick his face, but Ken just pushed him back. Ken sat with his legs dangling in the hole, and pulled out his cell phone and called Jeff’s number.

“Hello Jeff? I’m here in the barn.”

“Did you find it?” Jeff said as he answered the phone ring. “The envelope’s not here, Jeff.” Ken said. “Its gotta be.” Jeff exclaimed. “You’re not bullshitting me are you

Ken?”

“No way, Jeff.” “Look again, please.” “Just a minute then.” Ken eased himself back in the hole and felt

the sides as well as the bottom of the box shaped area. He stood up, and said,

“Not here Jeff.” “Man, someone’s taken it.” Jeff said. “Maybe your dad found it before you guys left?” “Naw, he would have never looked there, and he didn’t know any-

thing about it.”

“Well, who would have taken it?” Ken asked. “Maybe whoever left the Ferrari in the garage.” “Well, we don’t know who that was.” “Thanks Ken, for looking. Let me know if you found out anything

about it, will ya?”

“You bet. Sorry man.” Ken hung up, and then climbed out of the hole again.

Ken closed the trap door and re-piled the straw bales in the corner. Then he left the barn and walked on to Mrs Palmer’s house. He went straight to her garage to get the ride-on mower ready. He filled it with gas from a big red plastic gas container, and he checked the oil level. Before he could get on and start it up, Mrs Palmer came into the garage.

“Were you looking for something at the Blanchard’s Ken?” Mrs Palmer asked.

“Just remembering what it was like to play there with Jeff. We kept some old treasures in the barn. You know, neat rocks and hawk feathers that we would find. But there was nothing there anymore.” Ken replied sort of stiffly.

“Ah, boy things.” she nodded.

“Yeah.” Ken tried to laugh it off.

Mrs Palmer noticed Ken’s manner and let the subject drop. Ken got on the mower and started it up. As soon as the engine stopped coughing and sputtering, Ken shifted it into gear and backed out of the garage.

The grass cutting took two hours. Mickey gave up waiting and went home when he saw the family car return down the laneway. Mrs Palmer made some lemonade for Ken when he finished, and some cookies. She paid him $20 for his work, and thanked him. She did not ask him any more about what he was doing in the barn at Blanchards. He prayed that she would not mention it to his parents.

He walked along the road to go home. It was easier than walking across the field. He did not take any notice of the car or the people in the car that passed him by as he walked home. He was thinking of the homework he had to do when he got home. By the time he walked into his home he had forgotten about Jeff and the money envelope, and about Mrs Palmer. He was back to being himself. He went to his room and put away the $20 he got from Mrs Palmer. He counted the money he kept in his old cigar box and it came to $87. He knew Christmas was only a couple months away. He returned the cigar box to the bottom drawer of his dresser and covered it with socks. Then he went to find Roger and Mickey. On his way through the kitchen his mother said,

“Thank you Ken, for doing the dishes.”

“You’re welcome, Mom.” said Ken as he went through the door.

Jackie and Steve

October 13, Saturday

Steve picked up Jackie at 8 the next morning, as arranged. They made good time to Mount Dunham on mainly empty roads. They stopped at the hardware store in Mount Dunham and bought a flashlight. Then they drove out to the Blanchard farm. On the way they passed a young boy walking on the road a short distance from the Blanchard laneway. They continued on past the farm to the next sideroad, and pulled off.

“I wonder what that kid was doing.” Steve asked. “He wasn’t carrying a safe.” said Jackie. “No, but he might have found it and let the police know about it.” “We can’t worry about a kid.” said Jackie. She continued, “Too many people could notice if we tried to look in

the barn now. We’ll have to come back after dark.”

“What about the other houses on either side of the Blanchard’s?” “They shouldn’t be looking out the windows at night.” “What if the safe is gone?” Steve asked.

“Then we really have to get out of here, fast.”

“What do we do until nightfall?”

“We can get a room at a motel back in Mount Dunham.” Jackie said.

“Do you think anyone will remember seeing us from before?”

“I stopped at this house they other day. A retired university professor lives here by herself. I talked with her about the fire. When I asked if they found anything in the barns, she got quiet and looked at me funny. As if she recognized me. It was strange. She knows something, but I don’t know what.” Jackie said. “You’re right, someone might remember us. We’ll go to Arthur instead and get a room there.” With that Steve put the car in drive and sped off.

* * * * *

In Arthur, Steve checked them into the Sunny Crest Motel. For the afternoon they watched TV on the small screen. Around 5:30 pm they became hungry. Steve went out on foot and found a pizzeria, and bought a few pieces for the two of them, and some Coca-Colas.

Around 7:00 pm the sky was starting to get dark. Jackie was getting anxious.

“Once we have the safe we’ll go back to Owen Sound.” Jackie said. Steve acknowledged and started packing up his things. He had dressed into his black jeans, black sweatshirt, black coat, and navy blue touque. When they walked to the car the air was brisk, and you could tell there might be a frost tonight. Jackie drove the car back to Mount Dunham, and continued through town and turned onto Highway 12 east towards the Blanchard farm. Steve gazed at the sky, then said

“The stars are really bright tonight. There are millions of them.”

“I think it is more like trillions.” Jackie remarked.

“Whatever.”

“There is only a sliver of a moon, so no one should see you lurking about. Just keep your flashlight aimed to the ground.” Jackie cautioned.

The car was about 2 kilometers from the Blanchard farm when Jackie turned off the car lights and slowed down. There were no other cars on the road. Jackie could barely make out the edges of the road and rolled along quietly. When they reached Blanchard’s laneway, Jackie stopped and Steve slinked out of the car, but the indoor light of the car came on when the car door was opened. Steve got out quickly and shut the door behind him.

“Damn it. Fucking light.” he said. Hopefully no one was looking at that particular moment. Jackie drove on as planned to the next sideroad and parked the car. She fiddled with the indoor light so that it would not come on again when she picked Steve up. She turned off the car and waited in the dark.

“It’s sure dark out here and so quiet.” she thought. Just then a car zoomed past going west towards Mount Dunham on Highway 12. Jackie looked at her watch, 8:43 pm.

“This should only take about 20 minutes, at the most.” she thought. She stiffened when she thought she heard something outside the car. She peered around out the windows, but couldn’t see anything. Just then a gust of wind caused the bushes on the side of the road to whip against the wire fence. Jackie made a little sigh of relief to see the cause of the noise. Jackie did not believe in ghosts, but a night like this, in the country, could certainly create the feeling of ghosts. She looked at her watch again, 8:50 pm.

* * * * *

Steve groped his way to the barn. He saw the yard light at the Lawrence farm, and could see lights on in their house. He opened the barn door as quietly as possible and went in. He moved the flashlight slowly from one corner to the other. He noticed two tiny red dots aimed at him along one wall, then the red dots turned, and he saw that it was a rat, which scurried along the wall, then disappeared.

“Rats! Just what I need.” he said to himself. He moved his flashlight to the corner where the straw and trap door should be. The bales of straw were piled neatly in the corner. He went to the corner and scattered the bales of straw to get at the trap door. When he looked in the hole there was nothing. Empty.

“Shit, shit, shit.” he swore. He closed the trap door and turned to leave the barn. He could hear something enter the barn and he raised the flashlight quickly.

“Woof! Woof! Woof!” over and over, loud and menacing. Steve nearly had a heart attack. He thought it was a wolf. His flashlight revealed a black dog barking at him. Steve hated dogs, they always seemed to sense that in him. However, Jackie had warned him about a dog, and Steve had brought some meat, which he pulled out and dropped to the floor. The dog grabbed it and ran off out the door.

Steve went outside and pulled out his phone and called Jackie. Steve could see someone at the Lawrence farm in the yard, and lights came on outside at the house to the east too. This was not going very well.

“Jackie, come now, quickly. There’s a dog, and the neighbors are aroused.” Steve said and hung up. He ran as fast as he dared in the dark toward the road. He saw a car driving down the lane at the Lawrence farm.

Jackie started the car, backed up, then put it in forward and stepped on the gas. She was at the Blanchard laneway about the same time as Steve and the Lawrence vehicle. She didn’t bother to dim the lights, but put on her bright lights. This was to blind the driver in the Lawrence car coming at her from the west. She stopped and Steve opened the door and hopped in. The other car pulled in front of her to block the other car from leaving, also with its bright lights on. Jackie drove forward blindly, trying to turn left to go around the oncoming car. There was not enough room for Jackie to turn. The front right lights of both cars crashed together jolting Jackie and Steve. Broken plastic and glass from the collision scattered everywhere. Jackie stepped on the gas and her car pushed the other car backwards into the ditch. Jackie could hear the crunch of metal and plastic as her car pushed forward. The left rear of the Lawrence’s car fell into the ditch and the remaining light now aimed to the sky. Jackie’s car was free and she took off heading west. Steve was holding on, eyes wide open.

“My car! Oh shit.” Steve shouted.

Jackie turned left on the next sideroad and headed south. With only one head lamp, visibility was not great. “We’re sitting ducks in this car. We have to switch vehicles. What went wrong?”

“It was a stupid dog. Scared the shit out of me.” Steve said.

“Was the safe there?” Jackie asked.

“No, it was gone.”

“Who could’ve takenn it?” she asked.

“How should I know? Dozens of possibilities. The safe isn’t there anymore Jackie.”

Jackie came to a small village and spotted a house with two vehicles in the front. She nodded to Steve, and he knew what he was supposed to do. He got out of the car, and shut the door. Jackie drove on a little ways and parked on the side of the road.

Steve tried the doors on the first vehicle, and they were locked, but not the second. Even better, the keys were above in the sun visor. He climbed into the black Ford Expedition and put the car in neutral. He slid out and pushed it backwards out onto the road, quietly, down a gentle slope. Then he jumped in and started it up, then drove on. He passed Jackie and she turned onto the road and followed him. They came to another sideroad and parked both vehicles. Together they transferred all of their belongings to the Ford Expedition. They abandoned the old car where it was.

“You can phone the police tomorrow morning and report your car as being stolen.” Jackie told Steve.

As they drove back to Owen Sound, Steve said,

“Most likely the police have the safe and are keeping it quiet.” I think we have to give it up as a loss.”

“That’s about a million dollar loss Steve.” Jackie said angrily.

“We can’t ask the police about it, without telling them we were the robbers.” Steve laughed.

“If the police don’t have it, then it must be one of the neighbors.” Jackie said. Steve looked questioningly at her. “It’s got to be one of them, who else would even bother to stop there?”

“The arsonist. I bet he has it.” Steve countered.

Steve dropped Jackie off at her apartment. Then he drove the Ford Expedition to the other end of the city and parked it in a parking lot near the bus station. He wiped his fingerprints from the car and left it there with the keys in it. He took a bus back to his house. It was nearly midnight when he sat down in his kitchen to think about the missing safe.

Suddenly his backdoor opened and a man was on top of him, knocking him off the chair and onto the floor. The man had him pinned with his knees on his arms and his hands around Steve’s throat.

“Where is she?” the man yelled. Steve struggled, then got a look at the man’s face.

“Oh, it’s you again.” Steve said exasperated, as he recognized the face of Paul Selinger.

“You lied to me. She’s living here in Owen Sound. I saw her. Where is she?” Paul yelled again.

“Look man, she doesn’t want to see you. She’s living with another guy now.” Steve said.

“I don’t believe you.” Paul said.

“Here, I’ll give you her phone number and you can see for yourself.” Steve said. Paul got up and let Steve sit up on the floor. Steve got out his phone and entered Jackie’s home phone number. He handed the phone to Paul. Paul listened to it ring a few times, then a man sounding half-asleep answered.

“Hello?” the voice said. “I’d like to talk to Jackie please.” Paul said. “She’s not home now. Try again tomorrow pal.” and the phone

hung up. Paul looked at Steve.

“That’s Daniel. He’s very rich.” Steve said. “You’re going to show me where she lives tomorrow.” Paul said.

“I’m going to sleep here tonight.”

“Be my guest.” Steve replied, although he wasn’t pleased with his visitor.

“Where’s your girlfriend Judy?” Paul asked. “She finked us out to the police in Toronto, which is why Jackie and

I had to leave.”

“Finked on you about what?” Paul asked. “You know, one of our robberies.” Steve said. “Are you and Jackie still robbing stores here?” “Sometimes, not so often. Jackie’s got this rich dude, and I’ve got a

steady job, so we are almost retired from stealing, you might say.” Steve said.

“Where were you today?” Paul asked. “We hid some of our loot at the Blanchard farm, the place where we

hid your Ferrari. We went to see if it was still there.” Steve explained. “Was it?” “No, it was gone.” “Was it worth a lot?” Paul asked. “Yes, a lot. Jackie was pissed.” Steve said. “What are you going to do now?” Paul said.

“We have to figure out who has it, if we can.” Steve said. “Look, I’m tired, can we quit for tonight already?”

“Sure, go to bed, but don’t try to run off on me.” Paul warned.

After Steve went to his room, Paul placed some pots on the floor against the front and rear doors, so that there would be a noise if Steve tried to leave while he was asleep. Afterwards, Paul settled in on the couch, but it was only slightly more comfortable than sleeping on the floor. Paul fell asleep and all was quiet through the night.

The next morning Paul awoke, somewhat stiff and not well rested. The clock said 7:30. Paul sat up and rubbed his eyes, then recalled where he was. Paul went to the bathroom and relieved himself, then splashed water on his face. He walked from the bathroom to Steve’s bedroom and knocked on the door. There was no response, so Paul knocked louder. Silence. Paul opened the door and looked in the room. The window on the other side of the room was open and Steve was gone.

“Shit, fuck, damn it!” Paul cursed at himself. Paul sat down at the kitchen table, and thought about what he was going to do now. For sure Steve and Jackie would leave Owen Sound now, and Paul would have little chance of following them. He knew he had been close this time.

Sunday, October 14

October 14, Sunday

The phone rang and awoke Linda from a dream in her warm cozy bed.

“Hello?” she said groggily.

“Linda, this is Justin. We need your help in the office today. We’ve got missing cars, abandoned cars, the Lawrence car pushed into a ditch, and intruders at the Blanchard place.”

“Busy night, eh?” Linda said.

“Just get in here, please.” Justin begged.

“See you in half an hour.” she set down the phone. She usually slept in on Sundays with Ruffles.

She got out of bed and looked out the window. There was frost everywhere. It almost looked like snow. She shivered slightly and tiptoed to the bathroom on the cold floor. Twenty minutes later she was in uniform and ready to drive to the office. She fed Ruffles and cleaned out his litter box before putting on her coat.

On her drive to the station she wondered who would have been at the Blanchard place and why. How could so much happen in one night?

Were all of these events related or separate? She’d have to wait until she was told more. She arrived and parked her car around the back of the station in her usual spot. She could see Captain Muller’s car parked in his spot. Seemed to her like he was always at work. She had not spent a lot of time working with Captain Muller. She found that she was assigned to most of the fire related happenings, which was okay with her. Captain Muller seemed to be busy going to meetings and doing reports.

“Hello, Justin.” she called as she entered the building from the rear door. Justin looked down the hall towards her and said “Good morning.”

“What do you need me to do?” Linda asked. She noticed that Justin was by himself in the office.

“Captain wants you to go to Blanchard’s farm and talk with Fred Lawrence and Mrs Palmer. Whatever happened there last night might be related to the fire.”

“How did the Lawrence’s car end up in the ditch?” Linda asked.

“Fred tried to block the intruders from leaving, but they ran into his car and rammed it into the ditch, then fled.”

“Did he get the licence plate number?”

“No, but a car with one of its front lights out and fender off was reported abandoned this morning about 7 kilometers away, and another car was reported missing. The Captain is looking into those two events as we speak.”

“Okay, I’m off, I’ll take SUV 2. See you later.” Linda said, and stopped in her office to get her notebook on the fire investigation, before picking up the keys to the SUV, and heading out the back door.

As she drove east on highway 12 Linda spotted a red-tailed hawk sitting atop a pole along the road. It was fluffed up to keep warm, but was looking over the fields for rodents. As she approached, the hawk took off and soared, slowly winding to the ground, then it dropped with its talons outstretched. Breakfast was served. Linda thought it disconcerting how the end of life for the rodent was the continuance of life for the hawk. She thought how the abundance of rodents made it necessary to have hawks and other creatures to keep the rodents in check. Nature’s way of balancing the books. If there were too many hawks then there may not be enough rodents, so some hawks may starve. This balance of life and death, she thought, is the same for all living creatures. This was Linda’s religious experience of the day, as Linda refused to attend church of any kind. Linda believed that religion was a human invention, and man-made group belief systems were not necessarily religious. While Linda believed in God, she did not think any one religion had more claim to God than another. To Linda, God was the universe and nature, and everything that exists is part of God. God is just a word that means everything, including eternal. There is no heaven or hell, just being or not being.

Linda suddenly came out of her thoughts and realized she had just passed the Lawrence’s driveway.

“Was I asleep?” she said to herself. She remembered the hawk, then slowed the car as it came to the Blanchard laneway. She pulled in and parked immediately. She grabbed her notebook, and rolled out of the SUV and made her way back to the highway edge. Linda could see pieces of broken headlights and part of the fender of a car twisted, scratched, and broken lying on the ground and in the ditch. She could make out where the back end of the Lawrence vehicle had been pushed into the ditch, by the places where weeds had been trampled and torn from the ground. Linda noticed drips of oil and thought an oil pan might have a hole in it. She also saw the rear tire marks of the car that did the pushing. Everything seemed to back up what Justin had told her. She took pictures and bagged the pieces of tail lights and fenders.

She decided to have a look in the barns again. As she walked to the barns, she saw Mickey, the black lab, racing across the field towards her. Someone was calling Mickey from the Lawrence farm, but Mickey was ignoring those calls. Mickey greeted Linda as she was opening the barn door.

“Good morning Mickey.” she said, and petted him on his head while his tail was wagging up a storm. “Yes, you haven’t seen me for a couple days. No, I don’t have any treats for you.” Once Mickey was satisfied that everything was okay, he decided to return to the Lawrence farm and the person that was calling. Linda watched him run and waved to 132 CHAPTER 13. SUNDAY, OCTOBER 14

whoever was waving at the Lawrence farm.

Linda went into the barn and immediately noticed that the straw bales were no longer piled neatly in the corner, and she could see the trap door. She went over, lifted up and trap door and looked in the hole. There was nothing in it except a few pieces of straw. Linda wondered what could have been hidden in there. Did last night’s intruders find what they wanted and taken it? Linda left the barn and closed the door, then walked back to her SUV. She noticed Mrs Palmer was standing outside and looking over at her. She waved to Mrs Palmer and she waved back.

Linda got in the SUV and backed onto the highway and went further east to Mrs Palmers house, and pulled in. Mrs Palmer was in her garage when Linda pulled in.

“Good morning Mrs Palmer.” Linda called. “Oh, it’s you. I’m sorry, I forgot your name.” she said. “Linda. I’m the fire inspector.” Linda said. “That’s right, now I remember.” Mrs Palmer said. “Did you see what happened over there last night?” Linda asked

directly.

“Not a lot. I heard the dog barking first.” Mrs Palmer said. “You mean Mickey? He was over there too?” Linda asked surprised. “Oh yes.” “Then what?” Linda urged. “Well, Fred drove his car down the lane and towards Blanchard’s.”

she continued.

“Where was the intruder’s car?” Linda asked. “It drove up past my house just before Fred got there.” “So someone was waiting to pick up the person that went into the

barn?” Linda said.

“Yes, there were at least two people, one driving and one in the barn.” Mrs Palmer said. “Then Fred drove his car right in front of the other car. Both cars had their bright lights on, so it was difficult to actually make out what was going on. Next thing, I heard crunching and gravel being thrown. Then Fred’s car was pushed into the ditch and the other car drove off.”

“Where were you when this was happening?” Linda asked.

“I was standing just outside my front door. I walked over after the other car left to see if Fred was okay. But first I went back in and phoned the police station.” she said.

“That’s good Mrs Palmer. Did you see any other activity around the barn earlier in the day?” Linda asked.

“Yes, I did actually.” she replied. “Fred’s boy, Ken, was looking in the barn yesterday morning, and Mickey was with him. After that Ken came over and mowed my grass for me.”

“Did he say why he was in the barn?” Linda wondered.

“He said it was boy-stuff. I didn’t press him any more about it. None of my business.” Mrs Palmer said.

“Did you notice if he took anything from the barn?”

“He said he couldn’t find what he was looking for.”

“Thanks Mrs Palmer for your time. Please let us know if you see any further activity over there, ok? Call me about it right away.”

“I will, I will.” Mrs Palmer said as she waved and started back into her garage. Linda returned to her SUV and made notes in her notebook. Then backed the car out and headed further east until she came to the next sideroad. She could see where a car had pulled off the road, again from the trampled weeds along the side.

“There must have been two people. One dropped the other off at the Blanchard laneway under darkness, then came and parked here until the other one phoned. The idea was to pick something up from the barn. Mickey, however, broke up the party. Since the hole under the trap door was empty, that means they either found what they came for, OR they found the hole empty. With Mickey barking, the Lawrences and Mrs Palmer became aware that something was going on. Fred Lawrence drove his car over to get Mickey and to confront the intruder. That was a stupid thing to do. What if the intruder had a gun? Anyway, the driver gets a call, and comes in a hurry to pick up the partner. Fred tries to slow them down by blocking their path, but gets his car pushed into the ditch. The intruders get away.” Linda pieced the events together in her head. “What were the intruders after? What was Ken looking for? I need to speak with Ken.”

Linda turned around and drove back, passed the Blanchard farm and turned into the Lawrence’s driveway. As she drove around to the back of the house she noticed the Lawrence’s car. The rear end was crunched and there was mud all along the back. The front right portion of the car was also crunched in and touching the front wheel. Linda got out of the SUV and walked around the Lawrence’s car, taking pictures of the damage. Mickey was all over her again, tail and butt wagging like he had not seen her in months. Fred Lawrence came out of the house towards her.

“Good morning, Mr Lawrence.” Linda said. “My name is Linda Logel, and I’ve come to ask you about what happened last night.”

“I shouldn’t have tried to stop them. Look what they did to my car.”

“I agree, you shouldn’t have done that. They might have had a gun.”

“I never thought of that.” said Fred.

“Did you get a good look at the intruders?” Linda asked.

“No, they shined their bright lights at me and I couldn’t see anything. I just blasted my bright lights right back at them.”

“Could you tell how many people there were?” Linda continued.

“Not really, but at least two.” Fred said.

“Do you have any idea what they were after at Blanchard’s?”

“I thought they might be setting fire to the barns. Finish the place off.” said Fred.

“Mrs Palmer said that she saw Ken in the barn over there yesterday morning, before he cut her grass.”

“What? I told him to stay away from there.” Fred said bewildered.

“Could I speak with Ken for a little bit? It could be important.” Linda asked.

Fred invited Linda into the house which led to the kitchen. They were met by Grace Lawrence.

“Hello Linda, busy again I see. Would you care for some coffee or tea?”

“Tea would be nice, thank you.” Linda replied.

“Please, sit here at the table. It’ll be ready shortly.” Grace said.

Fred had gone on in the house, and Linda could hear him calling for Ken. When Ken came into the kitchen he showed fear and guilt at the same time. Ken knew what this was going to be about. Grace served Linda her tea and provided cream and sugar, both of which Linda ignored.

“Hi Ken, I’m Linda Logel. I’m the fire inspector, but today I’m asking questions about last night’s intruders at the Blanchard farm. Mrs Palmer tells me she saw you go into the Blanchard’s barn yesterday morning, with Mickey. Is this correct?”

Ken looked at his parents, then at Linda. He didn’t seem to know how to answer the question. So he kept quiet.

“Answer her Kenneth.” said Grace. His father thwacked him in the back of Ken’s head with a finger. You could hear the contact and knew that it smarted. Ken ducked his head and looked up at his father.

“Yes, I went in there.” he blurted out.

“I told you not to go over there, didn’t I.” Fred said gruffly.

136 CHAPTER 13. SUNDAY, OCTOBER 14

“Please Mr Lawrence, could I speak with Ken alone.” Linda broke in.

Fred and Grace nodded and left the room. Ken said “Could we go outside to talk?” Linda picked up her tea and went out the door with Ken following

her.

“I need to know what you saw in the barn.” Linda said. “I went over to look for something for a friend.” Ken said. “What friend? Why?” Linda asked. “If I tell you, do you promise not to tell anyone else?” Ken said

worriedly.

“I can’t promise anything Ken.” Linda answered. “My friend left an envelope with fifty dollars in it, on the underside

of the trap door in that barn. We used to play in there and use the hole as a hiding place. As a secret hideaway.” Ken explained.

“Why would he hide money there?” Linda asked. “He didn’t want his parents to know about it.” Ken said. “Was the money stolen?” Linda asked again. “Oh no, nothing like that. He earned it from doing odd jobs for

other people.” Ken said.

“Why did he ask you to get it? Why didn’t he get it himself?” Linda said.

“He doesn’t live around here any more. Too far away.” “So where is this envelope?” Linda asked. “It wasn’t there.” Ken said. “Was there anything in the hole?” Linda inquired.

“I looked in the hole for the envelope. There was nothing.” Ken said. “I phoned my friend and told him it was gone.”

“Then what?”

“Then I closed the trap door and re-stacked the bales of straw, and left the barn. I went straight to Mrs Palmer’s to mow her grass.”

“And there was absolutely nothing in the hole?” Linda asked again.

“Nothing, I swear.” Ken said.

“Thanks Ken. If I need to know your friend’s name, I’ll come back to ask you, but we’ll leave him as unknown for now.” Linda said.

Ken took Linda’s cup and went back into the house. Linda packed up and drove down the driveway and out onto highway 12 headed back to Mount Dunham.

On the way back she spotted the red-tailed hawk sitting on a fence post. This time it did not move. Linda wondered to herself how many mice the hawk caught per day, and how many it needed to survive. She guessed that soon the hawks would be headed south for the winter.

* * * * *

Captain Muller had returned to the station by the time Linda got back from Blanchards. Muller was eating in the lunch room with Justin, and Linda joined them.

“Morning Linda, what did you learn?” Don said.

“Well, it was pretty much as Justin told me before I went out there. Seems like there were at least two people. The driver dropped off the other, who went into the barn. While he or she was in the barn the Lawrence’s dog, Mickey, came in and started barking. All of the barking got Fred Lawrence worried about what was going on. Fred thought someone might be trying to set fire to the barns. Mrs Palmer phoned in here, and went over after the intruders sped away. The driver had been waiting at the next sideroad, then came to pick up the other one. The driver arrived first, then Fred tried to block their path, and got rammed into the ditch. Both cars had their bright lights on, so no one could see anything.” Linda stopped to take a bite of her sandwich.

“Do we know what the intruders were after?” Don said.

“That is an interesting point. Fred’s son, Ken, had been in the barn earlier that day. He claimed to be looking for an envelop with $50 in it for a friend, on the underside of a trap door in the barn floor. He said that it wasn’t there, and that there was absolutely nothing in the hole beneath the trap door.” Linda explained.

“We don’t really know if the intruders found what they wanted or not.” Justin said. “Maybe young Ken took whatever was in the barn?”

“I don’t think so.” Linda replied, “Mrs Palmer said he came directly to her place after leaving the barn. She did not mention that he was carrying anything with him. Whatever was supposed to be in the barn, I don’t think the intruder found it last night. That means someone else has taken it, whatever it is.”

“It may not have been in the hole.” Don said. “Your fire investigation report mentioned a small office, maybe it was in there.”

“The office was pretty barren as I recall.” Linda said. “But you’re right we don’t really know what they wanted, if anything. Ken Lawrence, however, said he re-stacked the bales of straw in the corner, on top of the trap door. This morning the bales were scattered all over the barn. So the intruder had to have looked under the trap door last night.”

Don said. “The abandoned car was owned by Steve Swider of Owen Sound. A white Chevy sedan.”

“That name was associated with the Ferrari that was destroyed in the house fire. One of the people who might have assisted in its theft.” said Linda.

“But get this, Steve Swider reported his car stolen this morning.” Justin said.

“That must be a cover-up.” Don said. “It’s not likely a coincidence that his name has now appeared twice in connection with the Blanchard place. I’ve asked my friend Murray in Owen Sound to detain Mr Swider for questioning. I’ll have to go up there when he’s in custody.”

Don chewed and swallowed, followed by a drink of soda, then

“The vehicle that was stolen near where the other car was abandoned has shown up in a parking lot in Owen Sound. Another coincidence? I don’t think so. Mr Swider is somehow involved in this himself.” Don said.

“There was a woman with Mr Swider and the Ferrari. Her name was Jackie Shambeau, sir.” Linda said. “Maybe she was the other person last night?”

“Possibly, although no one got a good look at either of them.” Don said. “Justin, could you look into the whereabouts of Jackie Shambeau. If she is also in Owen Sound, then I’ll ask for her to be detained as well.”

“Okay, captain.” Justin agreed.

“Right now we can only arrest them for leaving the scene of an accident, and attempting to cause bodily harm.” Don said. “We don’t know if they took anything from the barn, or were involved in setting the house on fire. Have you been able to find Nate Haslett and question him about the fire Linda?”

“Not yet sir.” Linda was not pleased at the slow progress in her first case, and now with all of this complicating mess she was being slowed down again.

140 CHAPTER 13. SUNDAY, OCTOBER 14

Nate in Hiding

Nate Haslett was in the locker room of the YMCA in London, Ontario standing in front of a mirror and shaving off his moustache and goatee. There was a kit to change hair colour to blonde sitting on the ledge below the mirror. Nate had become nervous when he learned that the police had been to his home and had also visited his favourite drinking hole. As he shaved he was thinking about other things he had to do to avoid detection. He decided to stay clear of Mount Dunham and his own home, for as long as it took for the police to stop their search for him, forever, if necessary.

Nate was a pyromaniac, and the urge to set something on fire grew stronger each day. He was constantly on the search for any type of abandoned buildings. The urge to set fires had become stronger in the last year. Instead of two or three fires a year, he was up to five in the current year. However, he had spaced these out all over Ontario so that it was unlikely any of them would be linked together. A couple of them had been categorized as accidents, and different investigation crews that did not communicate with each other guaranteed that Nate would not be connected to the group of them. However, he had been careless with the farmhouse. He should have never gone to the police station to ask about the availability of the farm land. He had exchanged licence plates with another pickup truck that was inoperable, and by painting his truck to red. These cosmetic changes to his outward appearance and to his truck should hide him for some time.

He had to stay somewhere, and if he used his credit cards he might as well just call the police and turn himself in. His sister had gotten $1500 in cash for him as well as a suitcase of clothes. Nancy Haslett did not know her brother was an arsonist, but she did not like the police and she did not want her brother to be treated unfairly by them. Neither Nancy or Nate’s mother knew that Nate was responsible for the fire that killed his father.

Nate was thinking about places to hide out. He remembered an empty motel he had seen near Clinton. He drove there from London up highway 4. He stopped at a Canadian Tire store on the way out of London. He bought a number of survival items including a sleeping bag, air mattress, water container, cooler, Coleman stove, and flashlights. He also bought a good rain coat. Then he stopped at a No-Frills grocery store and bought food supplies for the next few days. Then he went on towards Clinton.

On his drive-by of the motel he noticed that he could park his truck behind the motel and it would not be seen by passing cars on the main road. The motel was at a crossroads on the northwest corner of the intersection. Behind the motel was a hill sloping upwards with many towering fir trees which shaded the motel in summer and protected it from wind in the winter. The highway in front of the motel was two laned and the motel was about two kilometers from the little village of Lucan. Traffic on the highway was generally light. The motel itself was set back from the highway about 100 feet, so that at night, in the dark, passing cars would hardly notice the motel was there. The outdoor signs for the motel had been removed some time ago although the posts for the sign remained standing. The gravel drive in front of the motel was overgrown with weeds. The paint on the motel buildings was peeling everywhere, and the exposed wood was beginning to rot underneath.

Nate waited until he could see no cars coming from either direction, then drove his truck behind the main office and along the bank of 6 motel rooms. He got out of the truck and walked through the break between main office and motel rooms towards the front. He waited again until he could hear no vehicles approaching, then he went to the door of the first motel room with a numeral 1 on the door. He had to break in the door to gain access, and it gave easily to his kick. The windows were boarded up on every room.

Nate found the mirror in room 1 was broken, and the bathroom fixtures had been smashed to pieces. The metal bed frame and springs were still in the room, but all other furniture was gone. The carpeted floor was stained everywhere from water that had leaked in from the boarded up window. There were no windows or exits at the back of the room, which made Nate feel trapped.

Nate ran back to his truck when the way was clear, and grabbed his ax from the back of his truck. He went a little ways along the back wall and started hacking a hole in the wall. He kept working at the hole until he could easily pass through without getting on his knees. Now he could enter or leave without exposing himself at the front of the building, and he wouldn’t feel trapped on the inside.

Nate took in his cooler and Coleman stove, and set them up near the front door on a counter. Then he brought in his air mattress and sleeping bag and set up a place to sleep on the floor. The old springs and bed frame he leaned against the wall out of the way.

Nate found an old tarp behind the motel office that had been used to cover a pile of wood logs. He nailed the tarp over the boarded up window to his room, on the inside, to block light from escaping his room, and to help keep out some of the cold at night. He cleared out the center of the room. Here he set up his stove for cooking. He went back to the truck and covered it with camouflage netting so that the truck would not be obvious to planes or helicopters that passed over.

The first night he wondered how long he would have to hide from the police before they forgot about him. He also thought about setting fire to something. He was kicking himself about not checking out the Blanchard farmhouse more carefully. He could have a Ferrari, or the reward from reporting the Ferrari to police, if he had only checked the inside of the house more thoroughly. In his haste to make a fire, to satisfy his lust, he gave up a valuable prize. Who hides a Ferrari in the garage of an abandoned farmhouse?

Sometime past midnight, Nate heard gun shots outside and the sound of a vehicle on the gravel drive in front of the motel. More shots seemed to be coming closer to his room. There seemed to be several guns firing. He couldn’t look out to see what was going on, but he could hear laughter and loud bantering outside. The vehicle was getting closer to his room. Suddenly a bullet came through the door to his room and lodged into the back wall of the bathroom. Nate made himself as flat as he could on the floor. More bullets came whizzing through the door almost at the same time. Another round of bullets followed the first. Then they moved on and he could hear shots going into the office area.

“Got it!” came a yell from outside. More laughter and guffaws came from the car. Then the car spun its wheels on the gravel and the car zoomed back onto the main road and disappeared. Nate was shaken and no longer secure in his choice of hideout.

“Drunken idiots with guns.” he thought. “How often do they come by?” Nate went to the front door and opened it slightly. The number one that was nailed to the door now had a bullet hole through it, and several other holes around it. Nate closed the door. He shined a flashlight onto the counter. His cooler and stove had been hit with bullets. Luckily the fuel tank for the stove had not been hit. “Stupid bastards!” he said to himself. He got pieces of the wall that he had taken out of the back wall and nailed them over the holes in the door. “This is not a safe neighbourhood.” he said. Nate did not own any pistols or rifles. He prided himself in specializing in abandoned buildings. Buildings that he thought were better off being demolished and rebuilt. He was giving someone else something to do, as he looked at it. That was how he justified his compulsion to burn things. He wasn’t into shooting guns or killing people.

The next morning his stove still functioned, even with an extra hole in it. After he ate, he packed all his things back into his truck including the tarp he had used over the window. He looked over the place before he left. “Tonight I’ll be back for my goodbye.” he whispered to himself. “Now I have to find a new hideout.” He had some possibilities to the southwest near Exeter, in particular, an old stone house set way back from the road. The elderly couple that had lived there had been killed in a car accident only about a month ago. The farm was not yet for sale, as the surviving children had not decided what to do with the property.

It took Nate 45 minutes to drive there and he drove down the kilometer long laneway and parked at the rear of the house. He just sat in the car for most of the day. There were tall pine trees around three sides of the house, and a pear orchard on the fourth side. This allowed Nate to walk around the property in the afternoon. He simply froze beside a tree when a car passed by on the road. The house was still full of furniture. The wooden garage attached to the back of the house was cluttered. Nate explored the garage and he did not find a Ferrari, or anything of real value. There was an entrance to the house inside the garage. Nate tried the door, but it was locked. He took a pick ax that was already in the garage and bashed the door with it around the lock. The lock on the inside of the house was torn out of the wood to which it had been attached. Nate could enter the house, but he decided to wait until later. He closed the door and it stayed closed. He leaned the pick ax against the door frame. Someone would have to move it to get into the house.

There were two barns behind the house. One was a very large barn for storing hay and straw, and may have at one time been used for dairy cattle, but now it looked like any substantial wind would be able to blow it over. Nate did not feel safe walking through this barn as there were missing boards here and there in the floor. There was a pile of loose hay and straw in one corner, and Nate knew there had to be rats in the barn. There were a few old machines for farming the land and separating grain from the stalks of plants. Nate did not see anything of any value, to him at least, in the barn.

The second and smaller barn had been used for chickens. Although there were none present now, the smell of poultry feces was undeniable. Nesting boxes were arranged along one wall, two rows, one above the other. There was also a roosting area for the chickens to sleep at night. There was a small room for storing feed, Nate figured no one would hide anything of value in this barn. He did notice that he could park his truck behind this barn and not be seen from the house. He would cover it with the camouflage netting too.

Nate went back into the garage and this time decided to explore the house. The entrance from the garage led directly into the kitchen which was about 25 feet by 25 feet. There were four big windows in the kitchen, the kind you could sit in. Along one wall was an old chesterfield. A place for me to sleep, he thought. On the other wall was another door to the outside between the two windows on that side. There were two other exits from the kitchen into the front part of the house. One to a living room area and the other to a hallway to the two rooms on the left side of the house. These were a dining room and a smaller music room. Between the two exits was a small pantry for storing canned goods. Unlike the motel he had stayed in last night, the water and electricity were still functional in this house. But he would not be able to use the lights tonight because they could be seen from the road.

The end wall of the kitchen held the cupboards, stove, and sink. The refrigerator was next to the chesterfield between the two windows and opposite the exit doorway to the outside. About six feet in front of the sink was a large chopping block on four solid legs, and above it hung pots and pans. The ceiling was almost 11 feet high. In the middle of the room was a round table suitable for seating six people at one time. The flooring was old and well worn. The kitchen was obviously the hub of the activity in this house. The entrance to the kitchen from the garage was on the left side of the cupboards. On the right of the cupboards there were stairs that led to the bedroom above the kitchen. Nate climbed the stairs carefully which were piled with magazines on nearly every step. At the top he looked into the bedroom. It reminded him of a mouse nestbox. There was little clear floor space in which to walk. Chairs and dressers were piled high with papers, clothes, and who knew what else. The only place to sit was on the bed, which was as it was when the old couple arose the morning they drove to their death. Cleaning this place was going to take a lot of work, but a fire would clear it out quickly and easily, he thought.

Nate made his way through the bedroom to the other doorway. Outside the doorway to the left was a bathroom. There was a tub, but no shower. The ceiling in the bathroom slanted with the roof of the house, such that tall persons would have difficulty and would likely bump their heads often. Past the bathroom was a bedroom on the right, and another on the left. There was a smaller room at the very front of the house, which just had a couple of chairs in it. The front bedrooms were obviously not in use, as they were nearly empty without clothes or dressers. There was one large window in each bedroom. On one side the window looked out at the pear orchard, and on the other side the view was of the barns and trees.

The entire house was situated on the top of a hill, and commanded a good view of the countryside. From the front sitting room Nate could see where the laneway meets the road. He could not see the houses of any neighbors, which meant they could likely not see this house. From the sitting room there were stairs leading back down to the first floor. The bottom of the stairs were directly in front of the front doorway. Nate tried to open the door, but it was screwed shut all the way around the door frame. This door was only for looks. It was also clear that the doorway had been winterized to the extreme. Facing north it attracted much of the wind and bad weather, so the owners had decided not to use it. There was no sidewalk or pathway leading to or from the front door. The only exits were the two doorways in the kitchen.

The living room was grand and had a fireplace along the outside wall. The fireplace was a later addition to the house and not part of the original design, but it completed the ambience of the living room and made it cozy. Nate was beginning to like this house very much. He was just unsure of when he might expect unwanted visitors, relatives of the deceased. This was his only concern. The thermostat was on the wall and he thumbed the dial to warm the house for the night.

He looked at his watch, and decided to go back to the motel. First, he stopped at a diner and had meatloaf and mashed potatos for supper. It was good and filled him up. He stopped at a gas station and filled up, and also filled a Jerry can with 5 liters of gas. He continued on to the motel and parked at the back. He took the Jerry can into the room and splashed some of the gas around the room on the walls. Then he went out the back hole in the wall and emptied the remaining gas on the walls of the other rooms along the back. At the end he threw the can up on the roof of the motel. Then he went back to his truck. At the hole in the wall he lit a match and threw it onto a puddle of gas on the floor inside the room. Flames burst out with a “fluff” sound, and quickly spread inside the room. Nates eyes glazed over as he admired the flames dancing around. When the flames popped out of the room and onto the back wall, then Nate scurried around to get into his truck. By the time he started the truck and started backing out, the flames had spread halfway down the back wall of the motel rooms. Nate checked the roads carefully before pulling out from the motel. He did not want to be seen. He pulled out slowly and quietly in the dark and watched the flames grow in his backview mirror as he drove away. He smiled. He headed back to the old stone house for the night.

Paul Returns to Mount Dunham

Paul Selinger sat in the kitchen of Steve Swider’s rented house in Owen Sound and wondered what he could do now. Steve had disappeared during the night while Paul slept. His chances of finding Jackie now were almost zero. He had no idea where she lived. He didn’t know the last name of the guy she was living with, and he knew that Steve and Jackie would now be fleeing Owen Sound. He needed help. He showered and left Owen Sound. On the edge of town he grabbed a breakfast at a deli. He looked in his pockets and found the phone number for Linda Logel. He decided to see if Linda could help him with his search. Paul paid the waitress and got back in his car and headed for Mount Dunham.

An hour later Paul pulled into Mount Dunham and turned right at the corner where the police station was located, and parked along Buffalo street. Because it was Sunday the small town seemed deserted. It was half past noon. Paul entered the police station. No one was manning the front desk, but by opening the door an alert signal had been made in the station. Linda Logel appeared from the back area and she smiled when she saw that it was Paul.

“Paul, what brings you back to Mount Dunham?” Linda said.

“I need help.” he said.

“Why don’t you come back to our lunch room, and you can update all of us at the same time.” Linda suggested. Linda buzzed him through and led him to the lunch room.

“This is Captain Muller, and this is Justin Jankowisz. Sir, this is Paul Selinger.” They both stood up and shook hands with Paul.

“Please, have a seat. Help yourself to food from the machines if you like.” Don said.

“No thank you. Remember, I’m a chef. This stuff isn’t food, it’s plastic.” Paul said disgustedly. Justin looked carefully at the sandwich he was holding.

“Suit yourself.” Don replied.

“Sir, Paul said he needs our help.” Linda broke in.

“Help with what?” Justin said.

“I’ve been looking for my girlfriend, Jackie Shambeau. Last time I was here I discovered that a white car in the name of Steve Swider was purchased in Arthur in July 2011. The time when my Ferrari went missing. Linda helped to find Steve’s address in Owen Sound. So I went there. I found Steve, but he told me that Jackie had moved to Vancouver, and that she didn’t want to see me again. So I drove to a park on the bay in Owen Sound and sat there for some time trying to digest this news.” He paused and noticed that Linda was staring at him in an odd way. When their eyes met, Linda looked away quickly. Paul continued, “When I was leaving the park, there in front of me in a car was Jackie. I know she saw me too. Due to heavy traffic I was unable to pull out and go after her. By the time the way was clear for me, her car had disappeared. Anyway, I realized that Steve had lied to me. I drove back to Steve’s rental house. He wasn’t there. So I broke in and looked around trying to find Jackie’s phone number or address, but no luck.” Paul rested for a second. He glanced at Linda again, and caught her watching him again.

“Then what?” said Linda, munching on some fruit.

“I waited for Steve to come back.”

“Did he?” Linda asked.

“Last night, sometime past 11 o’clock. I jumped him and pinned him to the ground. He gave up fighting when he saw it was me. He said Jackie had this other guy, Daniel, who was rich, and that Jackie didn’t really want to see me again. Steve phoned Jackie’s number and let me take the call. Daniel answered and told me she wasn’t there and to fuck off.” Paul said. “This morning when I woke up, Steve was gone. He climbed out through his bedroom window. I was going to force him to take me to Jackie this morning. But now they are both probably gone for good from Owen Sound.”

“What do you want us to do?” Linda asked.

“I wasn’t forthcoming with you before.” Paul said.

“What do you mean?” Linda replied.

“Well, I didn’t tell you that Steve and Jackie are robbers. They plan and execute robberies of all kinds of stores using disguises and elaborate ruses. When they left Toronto with my Ferrari they were being sought by the police. Another woman, Judy, who was part of their gang and an ex-girlfriend of Steve’s, ratted on them to the police. So they took their loot and the Ferrari and left Toronto in a hurry.”

“Wait a minute. Were you part of their gang too?” Justin interrupted.

“Jackie tricked me into being part of two robberies. At one she acted like I was her boyfriend who had just asked her to marry him, and she jumped on me and made me fall into a pressure sensitive counter glass, which set off the security alarms in the jewelry store. While the clerk was focussed on us and the noise, Steve and Judy made off with a bag of jewelry. I had no idea what was happening until I saw Steve and Judy take off. That was my first day with Jackie.” Paul answered.

Paul continued, “As far as I know, none of them were employed, but seemed to live well from their occupation. Jackie actually bought my Ferrari with cash, so she had lots of money. She enjoyed the thrill and adrenalin rush of each robbery.”

“Did they ever use weapons in their robberies?” Captain Muller wondered.

“No guns.” Paul said.

“So you knew that the three of them were robbers, but you never told the police.” Captain Muller said.

“Yes, that is correct. Have you ever seen Jackie? I was totally under her spell. I didn’t care what she did, as long as I was not involved. I made her promise me. Living with her was heavenly. Before you ask, I do not know how many robberies or which robberies were committed by them after that. I just know that they continued. Jackie made sure to keep me out of the loop. Occassionally I would read about a robbery in the newspaper and wonder if it was them, but I never asked.” Paul said.

“Who was the leader of the gang? Was it Steve?” Linda asked.

“No, definitely Jackie.” Paul answered.

“You say when Steve and Jackie left Toronto with your Ferrari they also had some loot from their latest robbery?” Linda said.

“That’s right. They apparently hid it at the Blanchard farm, and they went there yesterday to retrieve it, but it wasn’t there. That’s what Steve told me last night.” Paul replied.

“Do you know what kind of loot it was or what it was worth?”

“No. Sorry.”

Linda looked at Captain Muller and said,

“That agrees pretty well with what happened last night sir.”

“Yes, now we know that there was something hidden in the barn, but that it wasn’t there last night. We know that Steve and Jackie do not have it, unless Steve was lying to Paul again.” Captain Muller said.

“I don’t think he was lying, Captain. He seemed pretty upset when he came in last night, before I jumped him. And he didn’t have any of it with him.” Paul said.

“Justin, could you look into robberies in Toronto back in June 2011, and see if we can determine what was taken. Then we’ll know what we are looking for.” Captain Muller said.

“Will do, sir.”

“Paul, how much do you know about Jackie? You know. Where was she born? Did she have any brothers or sisters? How old was she? Did she use other names? That sort of thing?” Captain Muller asked.

“Sorry, Captain. Even though I lived with her for over a year I never met or talked with any of her relatives. She claimed to be from Quebec. Her age, I’d only be guessing. She never told me much about herself. I’m ashamed to say that I don’t really know much about her. But I can’t believe she didn’t love or care about me.” Paul said.

Everyone sat around the table in silence for a minute. There seemed to be no further questions for Paul. Justin finally got up and cleared his spot at the table, then went out to the front. Linda finally spoke up,

“We have another problem now, Captain. We have the person who started the fire that burned down the farmhouse, and we have stolen property that used to be hidden in the barn at the farmhouse, but which is again stolen by persons unknown. My priority is to find the arsonist. Paul’s priority is to find Jackie, and now someone has to find the stolen loot. How do we proceed?”

“I think I might have a plan.” said Paul.

“Go ahead.” said Captain Muller, “What have you cooked up?”

“This might work, I think.” Paul said, “Since the loot was not where Steve and Jackie left it, then they must think that the police have found it. If Jackie knew for sure that it was here, she might try to come up with a plan to steal it back. Then she would come to us, and we wouldn’t have to chase her and Steve.”

“No one would be stupid enough to try and steal it from a police station, even ours.” Linda said, shaking her head negatively.

“Jackie would do it for the challenge and thrill. She’s like that.” Paul answered, “At the very least, she would come to Mount Dunham and look at the possibility of stealing them back.”

“Are you certain about that?” Linda asked.

“No, but do you have a better idea?” Paul said.

“I agree with Linda, Jackie’s not going to take that kind of risk.” Captain Muller said.

“Our best bet is to figure out where the loot has gone before Jackie does. From what you’re telling us, she seems smart enough to do that. If the police had found the loot, that would have been included in the reports about the fire. But since we never mentioned it, it sort of implies that we didn’t find it.” Linda reasoned. She continued, “Because Steve and Jackie did not find the loot where they expected, that means someone else has found it, any time since July 2011 when the Ferrari and loot were left there. The possibilities are first, someone in the Lawrence family. Secondly, the neighbor Mrs Palmer. Thirdly, the arsonist. And lastly, someone else unknown.”

“Have the Lawrence’s and Mrs Palmer been questioned about it?” Paul asked.

“No, we didn’t know for sure until now that there was any loot at all.” Linda said, “Once Justin determines what was taken back in Toronto, then we can ask them. The arsonist is in hiding, but it is unlikely he took it because he never noticed the Ferrari in the garage. If it was persons unknown, then we are out of luck I think. However, I think persons unknown would have run into Mrs Palmer or Mickey, the black lab at the Lawrence’s, and they haven’t reported any such visitors.”

“If it is not the arsonist or persons unknown, then that only leaves the Lawrence’s or Mrs Palmer.” Captain Muller summarized, “Do you think Jackie will have come to the same conclusion?”

“Most likely.” said Paul, he was impressed with Linda’s reasoning.

“If that’s true, then Jackie and Steve will be visiting Mrs Palmer and the Lawrence’s soon. We should get out to them soon.” Linda said.

“I’ll get my officers out there right away.” Captain Muller said as he rose and left the lunch room.

“Maybe we should warn the Lawrence’s and Mrs Palmer about possible visitors.” Linda yelled down the hall behind him.

“Not if they are suspects. They might hide the loot again and just lie to us and to Jackie and Steve.” Don yelled back.

Linda went back into the lunch room and told Paul, “We better help Justin figure out what they stole.” Suddenly there was work to be done, an urgency to move to action.

* * * * *

“Thank you for talking with me.” said Justin Jankowisz. “You say a robbery on June 28, 2011 involved a man and woman team. Do you have any names associated with those people?”

“The names left by an anonymous caller were Steve and Jackie, but no last names. The caller also left the Toronto address where Steve was living.” said Bruce Darrow of the Toronto police.

“I can give you the names. I have Steve Swider and Jackie Shambeau. They currently live in Owen Sound. They left Toronto in a Ferrari owned by Paul Selinger and hid out in an abandoned farmhouse here in Mount Dunham in July 2011. Could you tell me what was stolen and roughly its value?” Justin said.

“We guessed that the missing Ferrari and the jewel heist were connected because the name Jackie was associated with both. The jewels were valued at $843,000. They picked the most valuable pieces during the heist. The victims claimed the two used magic tricks, replacing the real pieces with imitations. This meant they knew exactly what they were after.” said Inspector Darrow. Then he added, “Thanks, we didn’t know they were in Owen Sound now, we’ll see about working with their police force and share information.”

“Ah, they are not likely there any longer. Do you consider Steve and Jackie to be armed and dangerous?” Justin asked.

“There were never any reports of weapons in their robberies, but that doesn’t mean they didn’t have any. So, I’d say yes, treat them as armed and dangerous.”

“Okay, thanks, we’ll let you know if we apprehend them or recover the jewels.” Justin said.

“Roger that.” replied Inspector Darrow. Justin hung up his phone and sat back in his chair. He was about to get up to tell the others when Linda and Paul came up to the front.

“We’ve come to help you out.” Linda said.

“No need. Just finished talking with a guy in Toronto. Steve and Jackie made off with $843,000 in jewels. He considers them to be armed and dangerous criminals.” Justin said.

“That’s a lot of money. I’m sure they are going to be back to see the Lawrence’s and Mrs Palmer. We better go see them right away. Steve and Jackie may be there already.” Linda said worriedly.

“I’ll see if Captain Muller can man the front desk, then I’ll go with you.” Justin said.

“Can I go with you?” Paul asked Linda.

“Sorry, that’s not officially allowed, but you can follow us in your car. Just don’t get in the way or say anything.” Linda answered. With that Paul left the station and went to his car and started it up.

A minute or two later Captain Muller and Linda drove out from behind the station in an SUV. Paul concluded that Justin was left behind to man the front desk. Paul followed the SUV. Once they were on highway 12, the SUV picked up speed. Linda was in a hurry. Captain Muller, in the passenger seat started calling his other two on duty officers to get them to come to the area too as soon as possible.

The afternoon had become overcast and dull with the possibility of rain. Linda was trying to determine how to make someone confess to taking the jewels from the barn. Or more likely, how to tell if someone was lying about it. One of her courses during her early training had been on questioning suspects of crimes and looking for outward clues or mannerisms that indicated lying. She had not applied the information from that course so far. Linda inherently assumed everyone told the truth, until it was proven otherwise. Linda was taken suddenly away from her thoughts when Captain Muller said,

“We’ll stop at Lawrence’s first. I’m betting one of their kids found the jewels and stashed them somewhere in their rooms. And that’s without the parents’ knowledge.”

“Possibly, but it seems to be a close knit family and the boys are not thieves.” Linda said.

“I know that. But they probably don’t know what they have or how valuable it is. They’re not hiding it from their parents, they just haven’t told anyone about what they found.” Captain Muller explained.

“We’re almost there.” Linda said, as she signaled to turn left.

158 CHAPTER 15. PAUL RETURNS TO MOUNT DUNHAM

Pursuit of Thieves

Mickey, the Lawrence’s black lab, greeted Linda as she got out of the SUV. Linda had to acknowledge the dog’s existence and exuberance. Captain Muller ignored the dog as much as possible. Fred Lawrence came from the garage and everyone said “Good afternoon.” Paul Selinger drove in, and introductions were made all round. Mickey seemed to like Paul and tried to push a ball into Paul’s leg. Linda noticed the interaction of Mickey and Paul, and she was glad that Paul took to Mickey and vice versa. That meant Paul was a pet type of person.

“Mickey, settle down. You’ll have to excuse Mickey. Too many visitors for one day.” Fred explained.

“Visitors?” Captain Muller asked.

“Yes, there were police officers from Toronto here earlier.” Fred said.

“What? They should have informed us they were coming. Excuse me while I phone the office.” Captain Muller said, and he went back to the SUV.

“Hello, Justin, Don here. Can you phone the Toronto police and see if they sent officers to the Lawrence place? If they did, ask them why they did not inform us before they came. Get back as soon as you can Justin. Thanks.”

Captain Muller turned back to Linda, Paul, and Fred.

“Has Linda explained why we are here today?” Don asked.

“No, we’ve been talking about my boys.” Fred said. “Please come inside and we’ll have some tea or coffee.” When they got to the door of the house, Don’s phone rang and he turned away to take the call.

“That was fast.” Don said, “What did they say?”

“Sorry Captain, they do not know of any officers that were going to Mount Dunham today.” Justin said.

“Thanks Justin.” and he closed his phone, then went into the house.

In the kitchen Linda and Paul were sitting at the table along with Fred. Grace was getting a kettle going on the stove, and fussing about cups and saucers. Don took a seat at the table and unzipped his jacket.

“Tell me about the officers that were here earlier. How many were there and what did they ask you folks? Did they show you any identification?” Captain Muller asked.

“There were two of them. The woman was in a pant suit, and the man had on a uniform, but no pistol or other items on his belt. Yes, they had identification, but we did not look at it closely. It seemed legitimate.” Fred said.

“That woman was much too beautiful to be a police woman, sort of like Angie Dickinson was, if you ask me.” Grace interrupted.

“What colour were her eyes?” Paul asked.

“They were light gray. She was indeed a very handsome woman.” said Fred. Paul looked at Linda, and she nodded back. Linda pulled out a picture of Jackie from her files, and showed it to Fred and Grace.

“That’s her, definitely.” they both agreed. “They wanted to know if we had ever seen a safe in the barns over at the Blanchard’s place.” Fred continued.

“A safe?” Linda asked surprised.

“Yeah, well I never go over there, so I’ve never seen a safe.” Fred said.

“I never saw one there either.” Grace added. “I don’t think the Blanchard’s ever had one.”

“What about your two boys?” Linda asked.

“Ken, my older son, claimed to have seen a safe, but not at the Blanchard place.” Fred said.

“Where did he see it?” Linda said.

“In Mrs Palmer’s garage. He noticed it when he went over to cut her grass. It’s been there a long time he said.” said Fred.

“Once the two police officers heard that, they left immediately.” said Grace.

“How long ago was that?” Linda asked worriedly.

“They came while we were eating lunch. About an hour ago.” Grace replied.

Linda, Don, and Paul got up and excused themselves, with urgency.

“Thank you very much. You have been very helpful. By the way, they were not real police officers. They were robbers and now Mrs Palmer may be in danger. We must get over there immediately. Please excuse us.”

“Oh dear!” said Grace with a shocked look on her face.

* * * * *

Captain Muller pulled into the driveway at Mrs Palmer’s house. Only Mrs Palmer’s car was in the driveway. The garage door was partially open. Don jumped out of the SUV and ran to the front door of the house and started knocking. There was no response.

Meanwhile Linda went to the garage and peaked in. There on the workbench was a safe with its door open, and empty on the inside. She left the garage and went to join the Captain.

Paul Selinger arrived in his car.

“Paul, stay in the car please.” said Linda.

“There’s no answer.” Don said.

“Try the door.” Linda said. Don tried the door and it was not locked. He opened it and called in,

“Mrs Palmer! Are you home?” he shouted. There was still no response. Linda came in and went past him. Both of them noticed that the house was in disarray. Steve and Jackie had been looking through the house.

“Hello, Mrs Palmer?” Linda shouted. The upstairs toilet was flushed. Linda ran to the stairs and took them two at a time to the top. Captain Muller was behind her. They both shouted her name again while looking for the bathroom. They both arrived at the door at the same time. Inside, Mrs Palmer was tied to the toilet and gagged. She had managed to flush the toilet with her elbow. Linda worked to untie her. She had a bruise on her forehead and some drops of blood had dripped from the side of her mouth.

“Oh thank you, thank you.” she nearly cried.

“Mrs Palmer, what happened? Are you alright?” Captain Muller asked.

Mrs Palmer pushed past both of them, out the bathroom and turned left. Linda and Don followed her to her bedroom. Inside it looked like a tornado had struck the room. Everything was torn inside out. Mrs Palmer went straight to her closet and looked in a shoe box on the floor.

“Oh no, they’re gone.” she cried.

“What’s gone Mrs Palmer?” Linda asked.

“My jewels.” she wailed. Suddenly, she realized Don and Linda were the police.

“Mrs Palmer, did you get these jewels from the Blanchard’s barn?” Linda asked. Mrs Palmer did not answer. Paul Selinger suddenly arrived in the bedroom.

“Please Mrs Palmer. We know there were jewels hidden over there.” Don said.

Mrs Palmer looked at Linda guiltily.

“Do you remember that I told you I saw a man and woman with a white car over at the Blanchard’s, shortly after I moved in here.”

“Yes, the information about the car helped Paul to find the man, in Owen Sound.” Linda said.

“Well, I didn’t tell you that I saw the man carry something heavy to the barn. The woman went with him. They were there for ten minutes or so. When they left, the man wasn’t carrying anything. So I concluded they left something in the barn.” she told. “Well, I was curious, so the next day after they had gone, I went over to explore. That’s when I found the safe, under a trap door in the floor. Later I brought my wheelbarrow over to the barn, while the Lawrence’s were out, and I brought the safe to my garage.”

“Why didn’t you tell us earlier?” Linda asked.

“Once I had the safe in my garage it took me about a week before I was able to break it open. I didn’t know the code to open the door. When I saw all those lovely jewels inside, I liked them. I knew they were probably stolen, so I couldn’t sell them anywhere. So I decided to make necklaces and bracelets out of them and to wear them myself. I took one of the jewels to a friend of mine who made necklaces and things as a craft hobby. But once she saw the jewel she knew it had to have been stolen, so she refused to deal with it. She hasn’t spoken to me since then. I’ve been trying to learn how to make necklaces and things on my own.” Mrs Palmer said.

“Do you realized the jewels were worth more than $800,000?” Don asked.

“I thought they were nice and I wanted to wear them to special do’s, but I had no idea what they were worth.”

“What about the people who were here?” Linda asked.

“A police officer was at the door when I answered it. He wasn’t very nice. A woman came in behind him, and they took my arms, one on each side of me, and they forced me upstairs, then tied me to the toilet as you saw using a rope they took from my garage. They wanted to know what I did with the jewels, but I didn’t say a word. That’s when the man slugged me in the face. I think I bit my tongue.”

“Looks like they found them without your help Mrs Palmer.” said Don.

“Not all of them, I have one or two in the basement in my craft shop area. After they tied me up they looked through my bedroom, then they quit after they found the shoebox. They left after that. You missed them by about twenty minutes.”

“We’re going to have to take those jewels in the basement, Mrs Palmer. You may be facing charges for not reporting your discovery of stolen property. You’ve certainly prevented us from capturing the robbers sooner.” Captain Muller was angry with Mrs Palmer.

“Do you know if these were the same people you saw put the safe in the barn?” Linda asked.

“I never got a good look when they were over at Blanchard’s, but it must have been them. Last week a woman stopped in, said she was looking for a particular person. While she was here we were talking about the fire next door. When she asked if anything had been found in the barns, then I knew she was the one that had been here back then. I didn’t say any more to her about the fire after that.” she said.

“You did have a good look at them today, however. Did they mention their names while they were here, either to you or to each other?” Linda asked.

“It was the same woman today. I hadn’t seen the man before. They only asked me where the jewels were. After that they didn’t talk to each other, until the man said ‘Found them’.” she said.

“You’re going to have to come to the station with us Mrs Palmer. I’m going to have to charge you with keeping stolen property, and obstructing our investigation. Linda could you take Mrs Palmer and get the jewels from the basement. I’ll put the safe into the SUV, we’ll need that if we ever catch them.” Captain Muller ordered.

“Yes sir.” Linda replied, then escorted Mrs Palmer to the basement. Paul turned to Captain Muller and said,

“Another cold trail it seems. We have no idea where they have gone, plus they have the jewels.” Paul was very frustrated, seeming only inches away from catching up with Jackie, to be eluded again.

“I don’t think we will be catching the robbers today Paul. You might as well go home. If you hear from Jackie or Steve, be sure to let us know.” Don patted Paul on the shoulder.

“I guess I will. Thanks for letting me tag along with you. And you let me know if you catch her. I would like to face her at least one more time, to talk with her directly.” The two men shook hands and Paul left the house. Captain Muller waited for Linda and Mrs Palmer to come back from the basement.

Linda, Don and Mrs Palmer locked up the house, and got into the SUV. Don put the safe from the garage into the back of the SUV. Linda had the jewels from the basement in her pocket. Mrs Palmer was a little disheveled from her ordeal and had not yet cleaned the blood from her mouth. The two other officers arrived and Captain Muller told them to go back to their usual routine. He informed them that the thieves had made their escape.

Rain was falling lightly but evenly now, as they drove back to Mount Dunham. Captain Muller phoned the Lawrences and informed them about Mrs Palmer’s safety so they would not worry. He knew they had acted too slowly, and he was angry with himself.

* * * * *

Steve and Jackie were feeling much better now that they had recovered most of the jewels, as they drove eastward away from Mrs Palmer’s house.

“Where are we going now?” Steve asked.

166 CHAPTER 16. PURSUIT OF THIEVES

“I thought we could go back to Toronto, exchange these jewels for some cash, then go our separate ways.” Jackie answered.

“Where are we going to stay?” “We could pay a visit to Chef Paul, and use his apartment for a few

days at least.” Jackie said.

“What if he’s there?” Steve asked again. “He’s not going to like seeing me again, and he’ll certainly bug you.”

“We’ll have to tie him up and gag him, catch him by surprise. It’s only going to be one or two days.”

“I want to give my respects to Judy.” Steve said. “You’ve got to stay away from her Steve. She’s set on turning you

in. If you see her, we could be in trouble again. I don’t like being forced to run.”

“Well I don’t like snitches.” “Look, you’ll have a good chunk of money in two days, and you can

leave Toronto and never see her again. Please, stay away from her, you’ll only get into more trouble if you see her.” Jackie pleaded. “Promise me you won’t go to her.”

“Promise!” Jackie said firmly. “Promise.” said Steve sullenly.

Home to Toronto

The rain poured heavily at times as Paul drove back to Toronto from Mrs Palmer’s house. The rain matched his mood. He had been very close to catching up with Jackie once again. He had missed her by twenty minutes. He also thought about Linda and how he had caught her staring at him at various times, but then she averted her eyes when he had made contact with them. To be sure, he thought she was very clever, and it was her logic that brought them close to catching Jackie and Steve.

Jackie flashed across his brain again. He could almost smell her body aroma and it reminded him of their bedroom escapades. Then came the pain of knowing that she left him and never contacted him for over a year. This was re-enforced by Steve’s comments and the fact that she was living with another man. These were like ice picks driven into his heart. He caught his eyes tearing up, and tears running down his cheeks. Not a lot, but a few drops. He started to get angry with Jackie and with himself. He had let his desire for Jackie take over his life. She obviously did not want him to be part of her life any longer. Why did he still want to see her again? These thoughts kept racing through his head all the way back to Toronto. As he got closer to Toronto the traffic became heavier, and he nearly pulled into another car that he had not noticed. Such near catastrophe’s have a way of waking one up. Watch what you’re doing, idiot.

Once he was within the city limits Paul decided to go to the Terrozi Bread House and check on the state of the kitchen. The weekend was nearly over. He would begin work again tomorrow, but he may need to order some food. He had been gone for a couple of days, and there was a co-worker, Gaetan, he could rely on to keep things stocked. Everyone was surprised to see him back in the kitchen, and it seemed all of them started to become more busy now that he was there. Paul looked out in the dining area and saw that the crowd was starting to dwindle out. He knew there would be a few more late nighters to show up. He checked the stores of food and made notes for ordering. Monday would be a light day for dining, so many of the items were not essential to order tonight. He made a list and gave it to Gaetan to order. Then he said good night and told them he would be back tomorrow, as usual.

Paul drove home and pulled into the parking garage of his apartment building. He sat in the car for a few moments. All the thoughts of Jackie went through his head again. At the end he found himself thinking about Linda again. Was his brain trying to convince him that he should move on, past Jackie, towards a new future? He suddenly felt tired, and got out of the car and headed for the elevator to his apartment.

He unlocked the door to his apartment and went inside, then hung his coat on a hook in the hallway, and took off his shoes. He walked a little ways and turned left into his small kitchen. There in front of him standing at the other end of his kitchen table was Jackie. She smiled at him and said sweetly,

“Hello Paul.”

Before Paul could react or say anything he was whacked on the head from behind, and he fell to the floor unconscious. Behind him was Steve holding a rolling pin. Steve and Jackie carried Paul to his bedroom and tied him up to the bed and gagged him.

An hour later, Paul came too with a terrible pain in his head. He tried to rub his head, but he couldn’t move his hands. He looked and saw his wrists were bound to the bed. He blinked and tried to determine where he was. After a few seconds he realized he was in his own bedroom. Then he remembered seeing Jackie in his kitchen. He looked around his bedroom, it was dark and he was alone. He remained still and listened. He could hear the television set in his living room. He started to make noise. He waited, then yelled some more through the gag. Finally he could hear steps and his bedroom door opened and the light was switched on blinding him immediately.

“Sorry about that Paul.” said Steve. “We need to hide out in your apartment for a few days. Then we will be gone again, and probably out of your life forever.”

Steve checked over the restraints to make sure Paul had not started to come loose. His hands were cuffed to the upper bed posts, and his feet were tied to the lower bed posts.

“If you promise not to yell, I’ll take the gag out of your mouth. Okay? Just shake your head yes. Great.” Steve undid the gag.

“Why did you have to bash me in the head?” Paul asked. “We needed your immediate cooperation.” Steve answered. “Where’s Jackie?” “She’s out. She didn’t really want to see you, but we had no choice,

no place to hide. We needed time to plan things out. In the meantime, you have to stay out of the way.”

“I have to go to the bathroom.” Paul said. “Sorry Paul, you’ll have to hold it until Jackie returns, she’s getting

you a catheter.”

“What!!!” Paul exclaimed. “You heard me.” “What about food and number 2’s?” “No food, no pooh.” “How long do you plan on imposing on my hospitality?” “Shouldn’t be more than a week, two tops, I would guess.” “Two weeks!!” Paul yelled.

Steve held up the gag and shook his head at Paul.

“Am I going to have to put this on you again Paul?”

Paul was furious, but realized he could do nothing. Steve left the room, turned out the light, and closed the door behind him.

* * * * *

A few hours later, the light was switched on in his bedroom again and woke him up. Jackie was there holding a catheter and bag. Steve had not been kidding. Jackie undid the belt and zipper on his pants then pulled his penis out the opening in his underwear.

“Please don’t do this Jackie.” Paul begged.

“Sorry, we can’t watch you all the time, and it would be unwise to untie you from the bed. This is best for both of us.”

She hooked up one end of the catheter to the bag and then carefully inserted the catheter into his penis. Paul winced at the sensation. He could feel it moving inside him. When it reached his bladder he could feel the relief of the escaping urine. Paul had never had a catheter before.

Jackie brought in a glass of water and helped him to sip from the glass.

“There, that should do you for now.” she said.

“Jackie, what about us? Where have you been?” he asked.

“There is no us anymore, Paul. You need to move on.” Jackie rose and went to the door.

“I don’t understand, what did I do?” Paul begged.

Jackie did not look at him, but turned off the light and shut the door. Paul was in darkness again. Later he heard the apartment door open and close. Someone had left. After that all he could hear was the television, faintly. His head was aching and he couldn’t rub it with his hands. Eventually he fell asleep again.

* * * * *

Jackie had hoped to stay in Paul Selinger’s apartment for one or two weeks before moving on. Her plan was to rest and relax from the sight of the police. Jackie converted some of the jewelry into cash while she was out on the first night. When Jackie returned to Paul’s apartment she split the money with Steve. After which Steve left the apartment to celebrate on his own. Jackie slept on Paul’s couch with the TV on, but muted. The next morning she awoke and there was a news headline about a murder of a woman. Police had captured the killer, Steve Swider. The victim was Judy Dixon, described on the news as a past girlfriend of Steve’s. Jackie lifted her head slightly and panic engulfed her face. Her mind started racing very fast, and adrenalin started pumping through her veins. How long would it take the police to figure out that Steve was involved in the jewelry robberies, and how long before he would break and tell the police about her? She realized she could not stay at Paul’s apartment. She could not count on Steve to keep quiet to the police. Jackie showered and dressed, then packed up her belongings. She realized that she did not have a car, but remembered that Paul did. So she searched for his keys and found them very quickly in a pocket of his coat hanging in the hallway.

Jackie went into Paul’s room,

“Sorry, I can’t stay a moment longer. Steve has gone and murdered his ex-girlfriend, Judy.”

“Why did he do that?” Paul asked.

“I’m sure he didn’t plan to. Judy has a way of getting under his skin, and Steve has a quick temper.” Jackie said. She took a key from her bra and left the key on the bed, just out of reach.

“That doesn’t mean you have to leave, does it?” Paul asked.

“Let’s just say that Steve can sing like a canary. Police will be all over this place today. I have to get out of town. We probably won’t see each other again Paul.” Jackie smiled.

Paul heard Jackie leave the apartment. The key was out of reach, but Jackie was right. Police showed up later that afternoon and broke into his apartment. Paul was embarrassed to be found tied up, with a catheter in him. The police released him from the bed, and Paul picked up the urine bag and hobbled to the bathroom where he withdrew the catheter slowly from his penis. He emptied the bag into the toilet. He undid his pants and tucked in his shirt, then did everything up. He washed his hands and face in the sink. He now felt the lump on the back of his head where Steve had hit him. He wondered if he needed to see a doctor about a possible concussion. He went to the living room to meet with the police.

“Paul Selinger?” asked Officer Bruce Darrow.

“Yes” Paul answered as he rubbed his head.

“How did you know Jackie Shambeau and Steve Swider?”

“Jackie used to live with me, until July 2011 when she stole my car and left Toronto.”

“Were you a member of their gang?” asked Officer Darrow.

“No I was not, but yes, I knew that they were committing robberies.”

“Why didn’t you tell the police earlier?”

“Because I was in love with Jackie, and I thought she was in love with me.”

“What about Steve Swider?”

“I met him only once. Jackie told me that he and his girlfriend, Judy, had gone to Vancouver.” Paul said.

“Steve killed Judy last night. What do you know about that?” asked Officer Darrow.

“I’ve been tied up on my bed for at least two nights. I’ve got this knot on my head, and I’ve not eaten for two days. What day is this anyway? Jackie told me about it this morning before she left.”

“Why were they staying here?”

“They recovered their lost jewelry in Mount Dunham a couple of days ago. They knew they could not go back to Owen Sound. So I guess they came to Toronto to change the jewels into cash.” Paul said. “Jackie had a key to my apartment when she left here in 2011.”

“So you agreed to let them stay?” “They were already in my apartment when I got home that night.

Steve crept up on me from behind and whacked me on the head with something. When I woke I was tied up in my bed.”

“Doesn’t sound like Jackie loved you, if you ask me.” said Officer Darrow. Paul just sat there and nodded in agreement.

“Are you going arrest me?” Paul asked. “Not now. Maybe later. Did Jackie mention where she was going?” “No.” The police were finished with him, and they packed up all of their

evidence collected from Paul’s apartment and left him sitting on the chesterfield.

After it was quiet, Paul phoned Linda Logel. “Linda, this is Paul. Paul Selinger.” he said. “You made it home safely?” she asked. “Yes, and no. Jackie and Steve were waiting for me in my apartment

when I got home. Steve knocked me out and they tied me up on the bed.” “You’re kidding!” Linda exclaimed. “Jackie left this morning. The police have been here for the last hour

or two. I thought I should let you know where they have been. Also, Steve has been arrested for the murder of his girlfriend, Judy.”

“How did that happen?” Linda asked. “I don’t know, that’s just what Jackie told me.” “You’re okay now?” Linda asked.

“Yes, except for this lump on my head.”

“You should see a doctor.”

“I will do that next. One other thing Linda.” Paul said.

“Yeah, what?”

“I’m sorry I left yesterday without thanking you for all of your help and letting me tag along.” Paul said.

“You’re welcome, I was glad to help.” Linda said. “Well, thanks for letting us know what happened.”

They had no more to say to each other, so they said goodbye and hung up. Next, he phoned his doctor to see if he could be checked out. The doctor could see him if he could get to the office before it closed for the day. Paul shaved, showered, and dressed for his doctor’s appointment. Before leaving he phoned the Terrozi Bread House to explain why he wasn’t in the last two days. He spoke with Gaetan, his friend. Gaetan said,

“The boss isn’t very happy Paul. You’ve missed a lot of work lately.”

“Look, I’m different now. Jackie is forever out of my life now. I’m back to my old self. I’ll be there tomorrow. I have to see a doctor today. If I’m okay, then I’ll be there tomorrow first thing.”

He hung up and headed for the door. He put on his coat and felt the pockets for his keys. The keys were gone. He realized that Jackie had probably taken them. He rushed down to the parking garage, and sure enough his car was gone, again!. Why does she always take my car? He didn’t have time to phone the police, but instead ran to catch a bus, and then a subway train to the doctor’s office. It was going to be tight on time.

Paul managed to enter the doctor’s office before closing time, but then he had to wait in the waiting room for another 20 minutes. He went outside to phone Officer Darrow about his car. He explained he was currrently at the doctor’s office. Officer Darrow thanked him for the call and got the details on his car, colour, make, plates.

The nurse finally called him in, and the doctor checked him out. Some special images of his head were ordered and Paul had to go to the hospital to have them taken. The doctor suspected that Paul had suffered a concussion. Paul was given a prescription to reduce the possibility of complications, and he was told to go home and stay inactive for two more days, after the images were taken.

Paul returned home after having the images taken at the hospital, He sat in his living room in silence, and thought that he would probably never see or hear Jackie again. He figured that was a good thing, given all the misery she had caused him in the past. He had no interest in finding Jackie any longer. He was glad that burden was now gone. Jackie could take care of herself, he thought. As long as she had her looks and brains, she would be able to convince many men to do what she wanted. Paul was ready to get back to work in his kitchen at the Terrozi Bread House.

That night Paul collected all the items in his apartment that reminded him of Jackie, and took them all to the trash bins in the basement of the apartment building. He was not going to waste any more of his time thinking about her. Starting today he was going to let his hair grow back.

176 CHAPTER 17. HOME TO TORONTO

Search for Nate

Linda smiled as she hung up the phone. She got up from her desk and walked to Captain Muller’s office and called in Justin. She repeated what Paul had just told her about Jackie and Steve being in Toronto.

“Linda, I think we have to let the OPP take over the hunt for Jackie. They were not the ones that caused the fire at the Blanchard home, but were only incidental to events there. We need to concentrate on finding Nate Haslett, our only suspect so far in the fire.” said Captain Muller.

“Yes, sir, I agree. I heard of a report of a fire at an old dilapidated motel last night near Lucan on highway 4. It’s close to Clinton which is out of our jurisdiction, but it could be Nate. I was going to drive out and see if I could find out anything.” Linda replied.

“Justin, call Clinton and see who has jurisdiction on that fire and ask them if it would be okay for Linda to have a look.” Captain Muller asked.

“Will do. Should I go with Linda?” said Justin.

“I don’t think that’s necessary, is it, Linda?” Don asked looking at her.

“No, not necessary Captain. I should be fine.” Linda nodded. Justin seemed disappointed and resigned to staying in the office another day.

Linda went back to her office and prepared her file with pictures of Nate Haslett. She had no way to be sure that Nate was involved in the fire at the motel, but she may be able to find a link if he was involved. She found an empty disk for the digital camera and got it ready. She sat still for a moment at her desk and tried to think of anything else she might need. On her desk was a folder for Paul Selinger. She opened it and saw the picture of Jackie. At least Paul had apologized for leaving without saying goodbye. She thought it was unlikely that she would ever see him again. His part in the case seemed to be over. Tracy came into Linda’s office and brought Linda out of her reverie.

“Here’s information on the motel fire from last night and the names of the officers investigating it. Justin said they would be happy to have you look things over.” she said.

“Oh, thank you. You two are so efficient.” Linda said smiling. “I’ll probably be out most of the day Tracy.”

“Gotcha, have fun.” Tracy said as she blew back out the door and back to her desk in the front area.

Linda took her things out the back door of the police station and loaded them into her car. She drove to the gas station and filled up. She saw Al, the tow truck guy, and waved to him. He waved, but did not seem to know who she was, and ignored her instead of coming over to talk to her. She realized she had not made an impact on him obviously.

“Be that way then.” she thought to herself. She had been trying to get to know the people in Mount Dunham, but most of them seemed to pay no attention to her. She made a deliberate visit to see Mrs Schrandt, the old lady who complained constantly to the police about minor infractions of the law. That seemed to upset Mrs Schrandt rather than put her at ease. Linda wasn’t sure whether to give up and clam up or to keep pecking away at the armor around everyone. Her only converts so far seemed to be Mickey, the black lab at the Lawrence’s place, Tracy and Justin.

Mrs Palmer had been let go after being charged with possession of stolen property. She would have to appear before a judge in court and probably pay a fine, but not necessarily go to jail. This was because she only had a few pieces of stolen jewelry remaining after Jackie and Steve were there. She signed a statement about how she came to obtain the jewelry, and how Jackie and Steve had taken it away from her. She gave the remaining jewelry pieces she had to the police and these were returned to the store from which they were stolen.

Linda was on her way by 9:30 am. Rain was coming down again, off and on, and the temperature was around 10 C. Winter was definitely coming. She took highway 17 west until it crossed highway 4, then she turned south. The scenery was similar to that around Mount Dunham, flat and open, desolate looking especially in the rain. When she got into Clinton she stopped at Tim Horton’s for a break before going to the police station. She had a peanut butter cookie and orange juice. She got to the police station and walked in. The officer at the front desk, took her ID and snapped her picture, then led her to Officer Reggie Lushbaugh.

“Just call me Reg.” he said. Linda nodded and said,

“I’m Linda Logel from Mount Dunham. I’d like to look around at the motel where the fire was. We had a fire at an abandoned farmhouse last Wednesday night. We suspect we know who set the fire, but he seems to be in hiding. I think he is a serial arsonist and he might have been involved with your fire.”

Reg shook his head negatively, then said, “I think our fire was set by a bunch of drunk high school boys. We heard they were out shooting at the motel on Saturday night. I think one or all of them went back last night and set the place on fire.”

“Do you have proof of that?” Linda asked.

“Not yet, but one of my officers is at the high school as we speak gathering up one or two of them. You can be part of the questioning if you like.”

“Well, I hope it is that simple.” Linda said, and hesitated. “Would you mind looking at this picture and tell me if you’ve seen this person around here lately.” Linda handed him a picture of Nate Haslett.

“What do you know, that’s Nate Haslett!” said Reg. “I know him, he’s been arrested before for disorderly conduct. Basically for getting drunk and causing trouble. He works for various farmers around here during harvest time. What about him?” Reg handed the picture back to Linda.

“He’s the one who likely started our fire. Have you seen him in Clinton lately, like Saturday or Sunday?” Linda asked.

“Nope, I haven’t seen him lately. I don’t think Nate is an arsonist. I’ll ask the other officers later for you, if you like?” Reg smiled. Linda did not like the way he smiled, sort of in a patronizing fashion.

“Do you have any other abandoned buildings in the area? Houses, farms, other buildings?” Linda asked.

“I can find out for you, but I don’t know of any off the top of my head.” Reg said.

“That would be great. Should I wait or come back?” Linda asked.

Reg picked up his phone and punched a couple buttons.

“Hi, Shirley? We have an officer here from Mount Dunham, and she would like to know about any abandoned buildings we might have in our area. Thank you Shirley. Also, can you put out a notice on Nate Haslett. If anyone sees him, then he is to be detained for questioning.” Reg waited, then said, “That’s right Shirley, thank you.” He then hung up.

“You can see Shirley on your way out, she’ll give you a list of the currently abandoned buildings. I’m sorry you wasted a drive out here.” Reg said as he got up to show Linda the door.

“I’d still like to look at the fire scene while I’m here.” Linda said as she rose.

“Be my guest. Shirley can give you directions to the site.” Reg added. “If we find Nate, we will let you know.”

“Thank you for your help.” Linda replied with as much gratitude as she could muster. She left his office and walked to the front area. There was an older woman, short and pudgy, with glasses. There was no doubt that she was the one who knew everything. She had probably worked there for 25-30 years. Linda was expecting trouble from her.

“Hi, I’m Linda Logel, should I wait somewhere until you get the list of abandoned buildings for me? Also, I need directions to the motel that burned last night.” Linda asked politely.

“You may sit over there, I should have everything for you in a moment. Would you like a coffee or tea?” she replied very softly and sweetly, not at all what Linda expected.

“Oh, nothing to drink for me, thank you.” Linda said, and went to a chair where she was directed to wait. After five minutes, Shirley brought a couple of sheets of paper to Linda, and said,

“Here you are Officer Logel. The first sheet is a list of abandoned buildings in our jurisdiction and the second is a map to the motel. If you need anything else, then please do not hesitate to ask.” Shirley smiled. Linda thanked her, then turned and left the police station.

Linda thought Reg seemed to have made up his mind about those kids setting the fire. Maybe he was correct, from his experiences. Linda just wanted to have a look to see for herself. She followed the map that Shirley had given to her, and soon she could see the smouldering site ahead. She pulled into the gravel drive and crawled slowly along. The rooms farthest from the main office were burned the least. It pointed to the fire beginning in the first room or the main office. She parked close to the main office, then got out and put on some boots, fire coat and helmet. She walked to the first room and looked in. Most of it had burned completely to the ground. There was a bed frame and springs. The room itself was covered in ashes. The smashed bathroom fixtures had not burned. She walked along the front of the motel. As she reached room 6 at the very end, the walls were still up, but the roof at the back had burned. She looked closely at the door to room 6. She could see the bullet holes which agreed with Reg’s statement that the kids had been shooting guns.

She continued around the end of room 6 to the back of the motel. The entire back wall of all rooms had burned. There was not much left. On the floor in room 6 was a melted Jerry can. She took pictures of it. She continued on along the back towards room 1. When she got to room 1, Linda noticed that the grass was bent over like a vehicle had driven over it a few times. There were four well packed places that indicated a vehicle had been parked for some time. Linda could see the path the vehicle took coming in and leaving this location behind the motel.

Linda was now behind the main office. It too had not burned completely, indicating more that the fire had started in room 1. She suddenly had questions for those high school kids. She went back to her car, took off her coat and helmet, and boots, then drove back to the Clinton police station. Upon entering the station Linda saw Shirley and asked,

“Are the high school kids here now?”

“Yes, come with me you can watch the questioning.” Shirley said. Shirley led Linda to a room from which she could watch Officer Reg asking two boys questions about the fire.

“Sir, we have told you that all we did was to shoot rifles at the numbers on the doors to the motel rooms, and that was Saturday night. We had nothing to do with the fire. We were all at home on Sunday night.”

“We’re going to have to put you in jail until you confess. Who else was with you on Saturday night? It could be one of them.”

“I want to speak with my Dad.” said one of the boys. “Me too!” said the other. Reg shook his head and got up to leave. Linda left the room hurriedly to see if she could go in and ask the boys some questions. She encountered Reg as he was leaving the interrogation room.

“Oh, it’s you again.” Reg looked surprised.

“Could I ask the boys a couple questions?” Linda asked.

“Certainly.” Reg opened the door and allowed Linda to enter. Reg followed. Linda introduced herself to the boys.

“Did you boys happen to drive behind the motel on Saturday night during your target practice?” Linda asked.

“No, there are no doors with numbers on them at the back, duh.” said one boy, the other boy smiled.

“Did you boys leave a Jerry can at the motel?”

“No, we didn’t have one. We didn’t set any god damned fire.” the other one retorted.

“Thank you boys.” said Linda. She rose and left the room, Reg followed. Reg took her arm and led her to his office.

“What was that about?” Reg demanded.

“There were tire tracks behind the motel that only went as far as the first room. My walk around the site showed me that the fire must have started in room 1 because that room was the most completely burned out. Lastly, I found a melted Jerry can in room 6. Only the back of rooms 2 to 6 were burned through.”

Reg looked perplexed.

“Those boys say they didn’t start the fire. They also stated that they had not driven behind the motel.” Linda said. Reg stared at her.

She continued, “That means someone else had been behind the motel. I think it was Nate. The melted Jerry can also would have been left by the person that started the fire. The fact that the tire marks in the grass indicated someone had parked for some time, I’m guessing that Nate stayed in the motel on Saturday night. Most likely in room 1, but I’m just guessing. He was probably shocked by the gun shots that night, and decided to move somewhere else.”

“That’s your theory is it?” Reg laughed. Linda nodded yes. Reg stared at her for a while, thinking. Linda continued,

“If we believe the boys, then the evidence indicates that someone else was there and perhaps caused the fire. I’m stretching to say that it was Nate, but Nate is in hiding somewhere. Unfortunately, he is likely nowhere around here any longer, if he was the one. Your men probably won’t see him. Thank you for your help today.” Linda got up and went to the door.

“You’re welcome.” said Reg, as Linda walked away. Linda got in her car and drove back to Mount Dunham. She stopped by the police station to offload her files and to take care of the pictures she took. She wrote some notes in her files. Then she went home for the night. Ruffles, her cat, welcomed her at the door and brushed against her legs.

That night Linda decided to find out where Nate had been working in Ontario during the last year and the dates. She would also look for places where Nate may have been arrested for disorderly contact. Then she would find out all of the abandoned buildings in those areas, and all of the fires and their dates. She had to learn as much as she could about Nate if she was to find him and arrest him. If she could, she would like to pressure him into moving around. The more he moved, then the more likely someone might catch sight of him.

The next day at work Linda set up a map of Ontario in her office on a bulletin board and put blue pins in places where Nate was known to have been working in the last year. Then she started inquiring about abandoned buildings in those areas and fires of abandoned buildings and used yellow and red pins for those. This was going to take a lot of time, but it had to be done. She asked Captain Muller if Tracy could help gather information. By noon time the map was quite full of yellow and red pins, but not many blue pins. Tracy found Nate’s arrest history throughout Ontario, and used green pins for those places on the map.

“It all looks random Tracy. Pins here and there. But the green and blue pins seem to clumped together in pairs. There are some green without any blue pins. Some of the red and yellow pins are near to blue or green pins, but not always.” Linda talked to herself mainly.

“This doesn’t really help me see anything.” Tracy said.

“We need more blue information.” Linda said at last. “I think I’ll visit Nate’s mother again.”

By seeing Nate’s mother, Nate would likely find out about it, and would keep in hiding, and maybe move to new areas of Ontario. Thus, Nate would feel the pressure, and could possibly make a mistake.

Linda took her own car again, on a drive to Mill Valley, on her own. She remembered the German Shepherd that barked at them during her last visit. At least today the sun was out, although clouds were drifting through. Linda remembered the road turns and found the house. The cases of empty beer bottles were still piled in the garage. She saw an old truck parked against the garage. She noticed that it did not have any licence plates. The dog was barking vigorously again as she approached the door. Irene Haslett came out of the house again to meet her, another cigarette dangling from her lips. Linda was glad to be talking to her outside. Linda abhored cigarette smoke. For some reason, if there was anyone smoking around Linda within 50 feet, the smoke somehow found its way to Linda’s nose like a magnet.

“Good morning Mrs Haslett. Have you seen Nate lately?” Linda asked.

“He’s not been here since you were last here.” Irene said.

“I was wondering if you could recall the places that Nate worked during the last year.” Linda said.

“Not exactly, I didn’t go with him, you know.” she replied.

“Did he keep any records of his work, like written on a calendar, or in a work file? He must have kept records for his income taxes. Does he have a desk in the house?” Linda was throwing out ideas, hoping one of them hit the mark.

“There might be something in his bedroom.” she said.

“Could I take a look Mrs Haslett?” Linda asked.

“Brutus might tear you to shreds if I let you in the house. I’ll go up and see what I can find, okay?” Irene said. Linda did not like this option, but she did not like confronting a vicious dog either. Irene went to the door and squeezed through. Just then the door burst open and Brutus jumped out and was on top of Linda in a flash knocking her to the ground. The dog jumped on Linda grabbing her arm that she held up to shielf off the dog from her face. Linda could feel the teeth clamping down and was surprised at the weight of the dog on her. The dog was growling viciously as it bit and tugged on her arm. With her feet Linda managed to throw the dog clear of her and it fell over, but righted itself quickly. Linda was also up quickly and faced the dog as it was ready to lunge at her again. As Brutus came at her in the air, Linda spun and lashed out both feet at the dog’s head. Linda was pleased to feel both feet connect fully with the dog’s head. Linda landed on the ground again, but bounced up again quickly expecting another attack. Linda fixed her eyes on the dog The dog was silent and did not move or make a sound. Linda waited for the dog to get up. Irene Haslett ran to the dog and kneeled beside it, stroking it. She turned her head towards Linda and said,

“You’ve killed my dog.” Then again, “You’ve killed my dog!”

Linda couldn’t believe it.

“You’re dog attacked me Mrs Haslett. I was only defending myself.” The attack had been so sudden that Linda was not sure exactly what happened. Had the dog gotten out of the house on its own, or did Mrs Haslett let the dog out at her? Linda took some pictures of the dog and the back door, of Mrs Haslett, and of her own arm.

“You better leave now, and don’t bother to come back.” Irene said.

Linda knew that Mrs Haslett was not going to help her any further, and maybe she had no intention of helping from the start. If Linda came back she would need backup and a warrant. Linda drove back to Mount Dunham and had Tracy dress the wounds on her arm, and related the story about Mrs Haslett and Brutus. Justin sympathized and was impressed that Linda had dispatched the dog. Tracy advised Linda to see a doctor, just to be safe.

“We need a warrant to search the house for records of Nate’s employment. And we need to get in there before Mrs Haslett or Nate hide or destroy the records, if there are any.”

“I’ll go with you this time.” Justin offered.

“Thanks, I think I might need it.” Linda replied.

Two hours later, the warrant was ready and Justin and Linda were off to Mill Valley again. The German Shepherd was still lying on the ground where it had died when Justin drove up to the house. Justin knocked on the door as Linda stood to the side holding the warrant. There was no answer. Linda heard something coming from behind the garage. She and Justin made their way to the back of the garage. They could see two people and one of them was digging a hole in the ground.

“You again, dog killer.” Irene said. Nancy Haslett, Nate’s sister, was digging the hole. She stopped to look over the visitors.

“Mrs Haslett, we’ve come with a warrant to look in your house for records of Nate’s employment activities.” Linda said handing Mrs Haslett the warrant. There was no resistance from Mrs Haslett.

“Nate’s room is up the stairs and to the right.” she said.

“Stay out of my room, if you please.” Nancy added.

Justin stayed with Mrs Haslett and Nancy while Linda went into the house.

As Linda entered the house she was in the kitchen area. The air was thick with the smell of cigarette smoke. The wall papered walls were stained slightly yellow from the smoke. There was an ashtray on the kitchen table full of cigarette butts and ashes. Other than that the kitchen was clean and tidy. Linda walked through and found the stairs to the upstairs, but first she looked around the ground floor. The living room had a TV, chesterfield, and La-Z-boy recliner. An end table had another mostly full ashtray on it. There was no sign of a desk or place to keep records. The next room was a dining room with large table and several chairs around the table. It had a nice hand made table cloth and did not look like it had been used for a long time. There were a couple of magazines on the table.

Through a door off the dining room were the stairs leading upstairs. Linda climbed them and looked to the right and left. On the left was a bedroom with an unmade bed and female clothes lying all over the floor. Apparently Nancy was not inclined to keep her room neat and tidy. On the right was Nate’s room, she assumed. The bed was made, but no one had been there. There was a desk next to the window with a computer on it. The closet door was open and Linda could see a file cabinet. She pulled the top drawer and found what she thought she wanted to find. She went slowly through the files in the drawer. Nate seemed to be organized like his mother. Linda found the spot with the records of Nate’s employment. She also found copies of Nate’s tax returns from the last two years, just what she wanted. She put these in her binder and made a list of the things she was taking.

The second drawer of the file cabinet contained owner manuals for various electronic gizmos around the house, and veterinary records for Brutus. Linda closed the drawers and turned her attention to the computer. She pushed the Power On button and waited for it to boot up. It was an older model computer and took a long time to boot. Linda did not need a password once the system was running. She clicked on the web browser. She clicked on browser history, and looked at the list of most recent sites visited. There was a real estate site and Linda clicked on it. Up came a property for sale. A beautiful stone house. Linda started an email message to herself and put the link to this house in the message. She left it open, she figured she might find some other things to put in it.

There were another two real estate sites and places that Nate had been looking at. She added those links to the message to herself. There was nothing else that seem relevant to Linda. She went to Nate’s email and looked through his Trash to see who had sent Nate messages lately. There were several thousand messages in Trash. Linda took out a memory stick and plugged it in. Then she copied everything in Trash onto the stick. She could go through it later back at the office. She was starting to feel sick. Nate too appeared to be a smoker as there was an overfull ashtray on his desk. Linda powered down the computer and left Nate’s room.

She glanced into Nancy’s room again, and saw a laptop computer on the night stand. She went into the bedroom and looked at it. It was on. Linda looked into Nancy’s email server and found the Trash. She plugged in her memory stick when she saw two messages from Nate in the list. Linda copied everything to her memory stick. Then she unplugged and left the room. Once she was outside she made a copy of the list of papers that she was taking. She went to the back of the garage where everyone was. Justin had dragged the dog’s body back to the hole, which was still not large enough for the dog. Linda gave Mrs Haslett the list of papers that she was taking with her, and signed and dated it.

Irene and Nancy said nothing. Linda and Justin thanked them for their cooperation and left. In the SUV, Justin started the vehicle and put the car in reverse. After turning around, he said,

“You smell like cigarette butts.”

“The entire house is a cigarette butt.” she sighed. “I was starting to feel sick to my stomach.”

“Did you find what you wanted?” Justin asked.

“I found some tax returns, and I copied all the emails he has received from his computer.” Linda said. “And guess what?”

“You found pictures of naked women?” Justin guessed.

“No, I didn’t. But he had visited three real estate websites and there were pictures of houses for sale. I just have to see if any of these places are abandoned.” Linda said.

“Sounds like you got everything you could have hoped for.” Justin said.

“Yes, it does. It seems like it was too easy, doesn’t it?” Linda asked.

“I expected Mrs Haslett and Nancy to kick up more of a fuss, to be honest. Maybe they were more concerned about their dog than about Nate.” Justin wondered.

“Nate sent Nancy some emails lately. I copied them too. Haven’t looked at them yet.” Linda said. “Maybe some of this stuff was set up to throw me off. I’ll have to verify everything.”

“That must have been some kick you gave that dog. Caught his head just right. He was big and strong.”

“You forgot to mention his powerful jaws.” Linda said, pointing to her arm.

190 CHAPTER 18. SEARCH FOR NATE

Nate Relaxes

Nate had been in the old stone house near Exeter for a couple nights. He slept on the chesterfield in the kitchen in case someone came in, then he could escape quickly, even though he may have to jostle someone violently to get out the door. The kitchen could be kept warm by shutting all doors into or out of the kitchen. The cans of food that were left in the pantry were still before their best before dates. Nate had bought his own cigarettes and beer, and several empty bottles were on the counters and dining table in the kitchen. The smoke haze in the kitchen was getting thick.

Nate was at the library in Clinton as the sun was going down, and just before the library was to close. He was composing an email message to his sister, Nancy. Nate had been upset when he was told about Brutus. The dog belonged to Nate and he took care of the dog, most of the time. Nate wanted to kill the pet of that police officer, if she had any. People like that usually don’t have any pets. Nancy Haslett had also told him about the police taking some of his documents from his room, and he knew they were concentrating on him as the primary suspect in the fires. He was glad that he did not keep records or souvenirs from his previous fires. That way the police would never know which ones he had actually started.

As he typed his thoughts into his message, an incoming message surprised him. The sender was Linda Logel. Nate had no idea who Linda Logel was, but the subject of the message was “Nate, Where Are You?” He read the message. His mind raced, how did this woman get his email address? As he read the message he realized Linda Logel was a Mount Dunham fire scene investigator. Nate felt trapped in a corner. The police were closing in on him. They might even be tapped into his sister’s computer, so he couldn’t send her any further messages either. He was raging angry and he swore inside his head. He deleted the message he was composing and logged off the computer. He tried to remember what he had told his sister in previous messages because now Linda Logel knew what it was. Had he told Nancy where he was staying now? He just said it was an old stone house, but he hadn’t told her where it was. But the IP address might give the police a clue to the general area of Ontario. He would have to move again, and soon.

Suddenly he remembered the real estate websites he had visited lately, and one was for the house in Exeter. The police will be checking that place out soon, if they weren’t already there. He gathered his things and left the library. He had to go back to the house and collect his things and leave.

He was on the country road coming up to the laneway to the old stone house, which sat back about a kilometer from the road. Before he turned he looked at the house and he could see lights on in the front of the house. He braked immediately and pulled to the side of the road. He looked more closely, and he could see more than one person moving in the house. Again, rage filled his head. Who was in the house? No matter who it was, he could not go in now, and whoever it was, they now knew that someone was staying and sleeping in the kitchen. They either were the police or they would likely phone the police. Nate drove past the laneway and to the edge of the property where he could watch the house. Where was he going to sleep tonight, he wondered. What did he leave in the house that might identify him. He checked his truck. He had most of his things, except the beer and cigarettes. He could buy more of those.

Soon a police car arrived and went down the laneway to the house. That’s it, he thought. Time to go somewhere else, this house was spoiled now. He put his truck in drive and drove off. It was getting dark, and he had no idea where to go. He drove south to the 401 highway and got on headed east. After an hour, he came to an exit with a trucker’s gas station, where trucker’s could park and get a few hours sleep, in their trucks. Nate pulled in and parked beside a tractor trailer that was at the edge of the property. His truck was obscured from the gas station. He walked to the station and bought some potato chips and chocolate bars, and another pack of cigarettes. Then he went back to his truck and made himself comfortable inside. After eating some of the chips, and a chocolate bar, Nate covered himself with a sleeping bag and used the unfilled air mattress as a pillow. He was soon asleep.

Around one in the morning Nate was jarred awake by the sound of air brakes and the roar of the truck engine parked beside him, as it shifted into gear and started to pull out. The driver had rested long enough and at this time at night the roads were pretty deserted. At first Nate couldn’t remember where he was, and thought the police might be there. He tried to move quickly, but he slid off the seat of his truck into the well, and then had a hard time getting up out of it. Once he was up he looked around. His truck could still not be seen from the gas station, but it was out in the open now. He moved his truck beside another parked trailer, out of sight from most other trucks again. Within a few minutes he was asleep again.

As the sun was rising Nate was awakened by tapping at his window. It was an OPP officer. Nate sat up and started to get out of the truck.

“Good morning sir.” said the officer. “I have a couple of things to talk to you about. Could I see your driver’s licence and vehicle registration?”

Nate hesitated at the request, but slowly pulled out his wallet.

“Sir, are you awake?” the officer asked.

“I am now, officer.” Nate handed him his driver’s licence, and then he reached into the vehicle to get his registration. That’s when he realized the licence plates he had put on his truck did not match the information on his vehicle registration slip.

“I’m sorry officer, but I don’t seem to have my vehicle registration with me.” Nate said worriedly.

“That’s not good sir. It should be available at all times in case you are involved in an accident.”

“I know officer, but I just can’t find it now.” Nate said frustratedly.

Another patrol car drove up and parked behind Nate’s truck. “First of all, you are not supposed to be parked here and sleeping overnight. This spot is for truckers only.”

Nate looked around, there were only two tractor trailers left in the big lot. All of the others had moved out. His truck was a sitting duck.

“I didn’t know that officer. I was tired and thought I would catch a few winks rather than driving on in a tired condition.”

“That’s why motels were invented, sir.” the officer said wryly.

“Yes, of course.” Nate said sheepishly. Nate was wishing he could escape, but decided to control himself.

“The other matter I would like to raise is your licence plates.”

Nate realized immediately what was coming next.

“Our records show that these plates belong on another vehicle owned by Neal Haslett. Your driver’s license says that you are Nate Haslett.”

“Yes, officer, Neal Haslett was my father, and yes, those plates are from my father’s truck. My plates came off somehow, and so I borrowed these plates.”

Just then, the other patrol car started flashing its lights, and the driver leaned out the window and yelled,

“Big accident west on 401 about 5 kilometers. All cars to respond. Let’s go.” Then the car backed up and peeled out of the lot. The officer questioning Nate said,

“I should arrest you, but this accident is more important. Get new plates for your vehicle. I’ll be checking up on you to see that you do it.” He handed back Nate’s driver’s licence then got in his patrol car and peeled out, lights flashing.

Nate started shaking once he realized how close he had come to being arrested. It would not have taken police long to link him back to the fire in Mount Dunham. He climbed into his truck and moved it to the gas station and parked. He went inside to use the washroom, and to get some coffee.

As he drank the coffee in his truck, he wondered what he should do. The police seemed to have a net over him, and they were pulling it in around him. If that was the case, then he might as well have some fun before he was caught. When he finished his coffee he seemed to have a plan of action. He started the truck,

“Mount Dunham, here I come.” he said, to no one.

196 CHAPTER 19. NATE RELAXES

Paul Sacked

Paul Selinger spent the last week recovering at home from the concussion Steve had given him. During this time the police visited him twice to ask questions about Steve and Jackie. The same questions came up on both occassions, and Paul answered the same way both times. There was not much else to do when there was nothing to hide. His car was found and returned to him, at his expense.

Today, he was feeling well enough to go to work. He showered and dressed, and managed to be at the Terrozi Bread House by 8 am that morning. As he went into his little office, he noticed a big box on the desk. A look inside the box showed a tumbled mess of Paul’s cooking gear. Paul wondered what had happened, a fire, a flood, pests? The was a letter in an envelope beside the box. Paul sat down and opened it. It was from his boss.

Dear Paul This letter is to inform you that your services are no longer required at Terrozi Bread House. Your employment is hereby terminated effective immediately. You are an excellent chef, but your dedication to the job has suffered over that last year. I wish you the best for the future.

Paul couldn’t be;ieve it, so he read it again. His friend Gaetan had been promoted to chef. Paul sat in the chair stunned. He had never been fired before. He felt empty and abandoned. He had no more expected income. He didn’t have a purpose. Surely this was a mistake, or a joke he thought.

Gaeten finally came into the kitchen, and came over to Paul.

“Man, what can I say? I didn’t know they were going to fire you.” Gaetan said to Paul.

“It’s not your fault Gaetan. I didn’t realize they weren’t happy with me.” Paul said quietly.

“What are you going to do?” Gaetan asked.

“I have to look for another job.” Paul said. “Although, in Toronto that may be difficult. Everyone will have heard that I was fired here. And rumours will have spread.”

“Did you ever find that girl?” Gaetan asked.

“Yes, but she made it clear she didn’t want to see me any more. So I’ve accepted that now, and I’m ready to move forward. She is no longer on my mind. I don’t need her. I was ready to start working again.” Paul said. “I suppose the boss isn’t in yet?”

“No, you know him, not until dinner time.” Gaetan said.

Paul stood up and shook hands with Gaetan and wished him well. He picked up the box of his personal belongings and made his way out of the kitchen and to his car. He drove home and picked up a newspaper on the way. When he got home he started looking through the help wanted ads. There were few ads for chefs, and what ones there were, they were not for chefs of his calibre. Mainly jobs for high schools or senior residences.

By noon he had drank 3 bottles of beer. He was feeling pretty sorry for himself. He put on his coat and went out for a walk. He was soon on the subway line and headed to the middle of Toronto. He found a bar that was open. He sat down at a table in front of a TV screen with the news headlines rolling by on the screen. Someone was talking about the stories, but the volume was not loud enough for anyone to hear. Paul drank, hard liquor now. Before he knew it, it was nearly 5 pm, and he was drunk. He hailed a taxi and was taken home. Once he was in his apartment, he thought the room was spinning and he felt as though everything in his stomach was going to come up. He made it to the bedroom and plopped face first on the mattress. The room was still spinning and he had the feeling that if he made another move of any sort, that he would vomit all over the place. So he remained motionless on the bed with his eyes closed. He felt as though he were flying as the room seemed to swirl around him. He closed his eyes and within seconds he was asleep.

The next morning, he awoke in the same position. He had not stirred all night, out cold so to speak. His mind had been a blank. He had been on the bed for over 12 hours. The sunlight seemed very bright. He rose and stumbled to the bathroom. As he sat on the toilet his digestive system emptied completely. He sat there for almost 30 minutes. His

mo�����